A Bibliography of Publications in Information Processing Letters: 1980–1989

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

20 October 2023 Version 2.16

Title word cross-reference

#10015 [Day80a].

(s, t) [Has81]. (s, t) [EMM89]. 0 [Llo81]. 1 [BP89, Llo81, Vid89]. 2 [BP89, BWD84, CL88b, Shi89]. 2.5 [Sle80]. 2n + k [Ber86a]. 2 × 2 [Cha86]. 3 [CM89]. 3NF [JF82]. 4n [Tie84]. 5 [Sze82]. 5 × 5 [Mak86a]. ^{∗} [McG83]. A [vR88]. B [CS83a, DLF87, GS86, Leu84, MS81, Oli80, PK88, Ans87]. modk [Van87b]. C [Güt83]. d [Eas81, EZ82, HY82, MS86, Sil81, Lea88]. δ + 1 [GP87a]. Δ_{2}^{p} [Sch83a, Bal86]. Dspace(log n) ≠ Nspace(log n) [KL87]. GF(p^{m}) [IT89]. K [Zak82, BYV82, DJ89, Eas81, EZ82, Er83c, GL89, HY82, HS85, MS86, PP89, Pet80b, Rei88, Sil81, Szw87, WS80b, YG87, Lea88]. k [WU88]. κ [IM88]. L [Bui77, Bui79, Day80a, Sal89, CH88b]. L_{1} [AN84, GS83]. λ [Mis88]. LL(k) [Nij82b]. log(N) [Ryt86a]. log log n [FMNT87, Sze87, Sze88, Sze89a]. log n [Dym88, Sze87]. LR(k) [AG81, Nij82b, Ukk85]. l × l [Kir88]. m [BY87, ER83b]. N [Rei87a, Roh83, Tam83, Ber86a, Bha80, FR89, Her89, IMR80, Ray87, Urb80, Wel85]. n − 1 [Tau89]. n^{1.618} [Vit85a]. n^{O(log n)} [KPS88]. NC^{0} [Hås87]. NC^{1} [IJR88, MC89b]. NC^{2} [ZG89]. NL [CH88b]. NP − P [Har83]. NPRIORITY(1) [IT88]. n × n [SV86]. O(1) [Che86, Tar83b, Wil84]. O(k^{2} log n) [Pet80b]. O(k^{3} log n) [Pet80b]. O(log n) [GIR88, PR86, Urb80, WT87, WS80b, SV86]. O(N) [TS86, Zem84, DS88]. O(n log^{2} n)

[HS89]. O(n^{0.695}) [EW86]. O(n^{1.5}) [HRY87]. O(n^{1/2}) [Deh86a]. O(n^{2}) [AB87b, Wel85]. O(n^{3}) [Dub88]. O(n log n) [Lip84, Röh82, KS87a, Nos85, Her89]. O(n log n + m log log n) [SH89]. O(n log n log log n) [KC87]. O(n × 3^{d^{2}}) [Cla86]. ω [DK84, Fre83b, LT86, Sao84, Yoc89]. Ω(log l) [Kir88]. Ω(n) [Lev87]. Ω(n log n) [BR89, Avi80, van80b]. P [HS81, MB85, BI85, CL80, GGPS89, Mio84, Mor88d]. PSPACE [Bab87a]. R^{d} [Kal84]. S [Day80b, LBM88]. T [Hat86, vR88]. Θ(N) [Wil83a]. Θ(N^{2} log N) [SR81]. × [GP88]. V [HS81, MB85]. ε [Vit80]. x + y [MA85]. Z [Jos88].

-adic [Mio84, Mor88d]. -ary [BY87, vR88, Zak82]. -automata [Sao84]. -Calculus [Mis88]. -chain [Shi89]. -closure [Szw87]. -CNF [CL88b]. -colorable [YG87]. -Coloring [GP87a]. -complete [GGPS89]. -Connected [IM88]. -cube [Bha80]. -Fold [HS85]. -Free [ER83b]. -gon [DJ89]. -grammars [NS86]. -Group [BI85]. -immunity [Bal86]. -infinity [Sal89]. -Languages [LT86, Fre83b]. -layer [CM89]. -Line [Wel85]. -lower [Tie84]. -matchings [Ans87]. -moves [Vit80]. -Order [vR88]. -Oriented [Güt83]. -path-free [BYV82]. -paths [EMM89]. -pebble-automaton [Sze82]. -polynomial [Hat86]. -process [FR89]. -processor [BWD84]. -Rational [DK84]. -rule [Yoc89]. -sided [Ber86a]. -sorters [PP89]. -stability [Day80b]. -stable [Bui77, Bui79, Day80a]. -tree [CS83a, DLF87]. -trees [GS86, Leu84, MS81, Oli80, Tam83]. -unification [LBM88]. -writer [Vid89].

03020a [Bal87]. 0L [Kob83].

10 [Avi80, Hel81, Llo80a, Mei81, vW80a]. 11 [Mor81]. 12 [Bin82, Bv82, Kri82]. 13 [NI80]. 14 [Kuč83, Tar86, YWL85]. 15 [GPB83, Žák84]. 16 [SY93]. 17 [Wyr84]. 18 [Bes84a, Sar80]. 1981/82 [Žák84].

21 [PG87]. 22 [AG88b, LMW88, TTv86]. 23 [AG88b, MT87b]. 24 [TV87]. 25 [Bal87, CIR88, VT90]. 26 [Sak88]. 27 [Flo91]. 29 [TM89, Ver89a, Ver89b].

32 [Dun90, Gro91]. 35 [Pet91].

56 [Day80a].

6 [Day80a]. 68034a [AG88b, TTv86].

68110a [Sar80].

7 [Fou79, Ov80].

8 [AJV86, APT82]. 80 [Dai86]. 80b [APT82, Fou79, Ov80]. 80e [NI80]. 80f [Sar80]. 81a [CM80b, Llo80a, vW80a]. 81k [Mor81]. 82d [Bv82]. 82i [Bin82]. 82j [Kri82]. 83g [Kuč83, YWL85]. 83h [Žák84]. 84m [SY93]. 85a [Wyr84]. 85h [Bes84a]. 87b [PG87]. 87j [AG88b, TTv86]. ’88 [Bli88]. 88e [CIR88]. 88i [Bal87]. 89d [Flo91]. 89m [Ver89a].

9 [And80b, CM80b]. 90b [Ver89b]. 90k [Gro91].

A. [Ov80]. Abelian [BI85, Ili85]. Absence [Udd86]. absolute [Wil83d]. abstract [GS88b]. abstraction [CDL87, KS87a]. accelerating [NA89]. acceptable [Sze88]. Acceptors [NA81, Nak81, NI80]. Access [MST84, ZBS83, CSO89, Deb80, Kuč82, Kuč83, Lee89, Mye83, San88, SY88, TJP87]. Accesses [PM84, CMT88]. accessibility [Cou81]. Accessing [Gun82, Win81]. Accuracy [Les83]. Achievable [Vit80]. achieving [GY87]. Achromatic [Bod89]. Active [Tho85]. activity [DLF87]. Acyclic [EG89, Pit87, GT89, Ita88]. acyclicity [dM88]. Ada [WS80a]. Adaptive [CH84, PM88, LP89b, WW89a]. addendum [Kuč83]. adder [Joh80]. Addition [Son85, Vol85, Cha88, Gra80, Mar85a, Mar85b]. Addition/Subtraction [Vol85]. Additive [Pan86a, CP82, Pan81a]. Address [KR85, NI83, RSD81, BDS82]. adic [Mio84, Mor88d]. Adjacency [RSR80]. Adjacent [Atk85]. Adjunct [YJ83]. Adjusting [Pag85, Wog89]. advances [FL87]. adversary [AK81]. adversary-based [AK81]. Agreement [TC84]. ahead [FV89b]. AKS [BP85b]. algebra [Hou87, Ida83, VB86, Sza88]. Algebraic [Mio84, Sei88, SV86, BM82, Dem82, FY80, Lau80, Pan81a, Ram87, RH88, SR88a]. Algol [Ten87b]. Algol-like [Ten87b]. Algorithm [AD86, AB87b, Bar81, BH87, DP83, DFv83, DD81, FM87, GS87, Ham86, HS86, IS86, JM86, KNI86, KC87, Las86, LR86, MF84, Mak86a, Mar86b, MA80, MNS84, NP85, Ryt86a, Sat86, Sch82b, Slu87, SV86, Tak86, van80a, AK89, Akl84, AGS89, Akm87, And79, And80b, Ans87, Apo87, AC88, ARG86, AG88b, APT79, APT82, Ata83, AC89, AS88, Awe85, BES81, BH89, BDD^{+}82, Bin82, BP87, Bir85, Blu86, BM89a, BM89b, Bui80, CL80, Car86, CDS89, CS83a, Cha83b, CCL89a, CK89, CLW80b, Cid88, Cla83, Coh83, CY85, Col88, CL88b, Cro81, CR89, DM89, DS88, Der80, Dub88, Dye80, DF84, EM84, Ebe81, EMM89, EY86, Flo91, FZ82, Gab81, GT88, GM88, GT89, Gol89, Gol80, Gos89, Gra72, GZW85, HS88a, HZ87, HS89, HL87a]. algorithm [HRY87, Hua81, II86, ISHY88, IT89, Iwa89, Jaz80, KK89a, KNT87, KST82, Klu84, KS87b, LMT87, Lev82, LP83, Lip84, LD80, Llo89, Loi80, LP82, Mak81, MU84, MM85, MG89, MR84, Meh88, Mét83, Min88, Mis89, Mit88, OIM84, Ola89, OT89, Ov80, ÖA88, Pap85, PR86, PS80, PS81a, Pet80b, RGG89, Ram89b, RM85, RR89a, Ray89a, Ray87, Roh83, Rub87, SSA89, SS89, SM92, SS84a, SV88, Sha88a, Sha82a, Sha89, SIM89, SH89, Shi89, Sho89a, SY88, Sla82, Sle80, SPL^{+}89, Sri87, Ste89, Sys82, Szp86, TTv86, Tar82b, Tar83a, TV87, TM89, TL83, TV81, Tsa81, Urb80, VT90, Ver87, Ver88b, Ver89b, WW89a, WT87, Wat81, WS82, Whi81, WW85, Wil84, WS80b, Yap83, ZG89, Zem84, Zer89, Zhe83, ZC87]. Algorithms [Cha86, CK88a, CH84, Dv82, Fre84, GT83, GLPG82, GK86b, GS84a, Hig86, HV87, IMR81, Ili85, IMM81, KR85, LS86a, Mor85, NM82, NI83, Pan86a, Van87a, AR80b, AB87c, Bar84, Ber82, Ber88b, CYL88, Deh86a, FS88a, Fra86, Fre80, Fri87, GP86, HL88, IM81, JM88, Kal84, Ko82, Man82, Mar86a, Muk89, Mur83, RR89b, Ray89b, RTV89, Roh88, RS89, SR88b, Ver88a, Ver89a, VV87]. alignment [Lee89, WW85]. Allocation [LR84, MF84, Slu87, App87, CP82, Cha85, NM87, RTV89, Rob80, Slu85]. Almost [TS86, Car87b, Wri87]. Alpha [BH87, McG83]. Alpha-Beta [BH87, McG83]. Alphabet [Car84, Car83]. alphabets [CC87]. Alternating [Hro85a, LL88, CIRB87, CIR88, Fre81, ITT82, Sze89b]. Alternative [Bar81, VU86, AGvG89, Ans87, Szp88]. Always [Pet91, Sni86, Dil87b, Kan83, VT90]. Always-true [Pet91, VT90]. Ambiguous [Tar82a, AFG87]. among [Fei89]. amortized [GST89]. Analyses [GPB82, GPB83]. Analysis [AR80a, Ili85, MA80, Mis88, Ram89a, Tam83, Yao80, Bab87b, Blu86, CDS89, FV89a, Fre80, Gun89, Jaz80, Ram89b, RG89a, SS84a, Szp86, Szp88, Tar82c, Tar86]. analytical [Bar84]. AND/OR [PPK88]. Andrei [Kot89]. Anf [Fre83b]. angle [Har89]. angular [RD87]. annealing [FS88a]. Anomalies [LS86a, Vos88]. APL [EPS82, TB84]. Appear [Kat84].

Application [Hug86, Ryt86a, AN86, Del83, Hou87, LLH89, Mai82, RH87, VB86, Vos88, Wat80]. Applications [BM81, Mil83, Sup85, Cul89, Szp88]. Applicative [Bro83, Mei85, Mye83]. approach [Ans87, BSM^{+}88, BG89, BCC83, CMT88, Mio84, RR88b, Rei87b, SS81, Vog89, WM89, Yoo85]. Approximate [GS86, Leu84, SSS86, DM89, SS83]. approximated [Sch81a]. Approximating [EK89, Les83, PM84, Pob86]. Approximation [IMM81, MSR85, BH89, Flo91, GP86, Meh88, Mih89, RG88, TV81]. approximations [Wu89]. April [Avi80, LMW88]. Arbitrary [Sha83, DS88, Hua88b, SY88]. arc [AH87, RR89b, SH89]. Architecture [LR80, BT87, DK88, GS81, ZM87]. arcs [Yap83]. Area [CL86b, ĎSVT85, Lot84, PV80, Bre80, DJ89, FG89, SR81, SV87, Vř’t87]. Area-Time [PV80, CL86b, Lot84, Vř’t87]. argument [FH87]. Arithmetic [Cho82, Ros86, SM83, Jaf81]. Arithmetical [Sza87]. Armstrong [FV83]. Arora [TTv86]. arrangement [CV89, ten80]. arrangements [Mat88, Wee88]. Array [CW86, Nak81, Ros86, DK88, Jaf81, NA81, RT85b, Wan80, pW82, Yod84]. Arrays [CC83, AN86, Deb80, IJ89, LT88b, PW87]. arrivals [Sha82a]. Articulation [Lau85]. artificial [Ram87]. ary [BY87, Zak82, vR88]. Ascending [Kal85]. ascent [Kru88, Rob89a]. Asserting [GS84b]. Assertional [Pet91]. Assignment [AR80a, AR80b, BMRS88, CCL89a, CL88b, SSA89]. Associative [Ore82]. assumption [HLY80]. assure [FL82]. Asymptotic [Dob82, MKK89, Blu86, Szp86, Weg82a]. Asynchronicity [Pan86a]. Asynchronous [Van81, Bod88a, Maz88, MW87, Tsi89]. Atomic [AG88a, Vid88, Vid89]. attack [Hel81]. Attribute [RB82, Alb89, TW89]. attributed [NS86, SIN87]. attrition [Sun89]. August [Dun90]. authentication [IW81, Kaw87]. authorization [PP80]. authors [CIR88]. Automata [BP85a, Cza80b, Miy84, Mon84, Ryt83b, Ryt84, SV84, UMS82, Wät85, Bud87, CR87, Fre81, IT80, Kin80, Mic81, Mos82, NI80, Rob89b, Rys83, Ryt81b, Ryt81c, Ryt82a, Ryt83a, Ryt89, Sao84, Var89a]. automated [VD87]. automatic [AK86, Gin82]. Automating [Koc87]. Automaton [LY86, Sto82, Eme87, Sze82]. Auxiliary [McC89, Mor88a]. Average [CDS89, GPB82, GPB83, Lav89, MPY87, Ram89b, Reg81, Ver88a, Ver89a, BY87, Bop89, Buc89, Lin86]. Average-case [CDS89, Ram89b, Bop89]. AVL [Che86]. Avoidance [FM87, Klu84]. avoided [Kes80]. avoiding [Gol87]. Axiomatization [Sza87].

B [AG88b, Bv82]. back [Wal83]. backout [Kre87]. Backtrack [Wil84, Mit88]. backtracking [BW80, Bru81, MTK89, Wol89]. backup [HK81a]. backward [PPK88]. Bad [WS88, Lin86]. Balanced [KK83, Ric83, Sil81, Tar83b, Sem81]. Balancing [Sni86, VV87]. Band [CRL85]. bandwidth [SM87, WW89b]. Banker [MF84]. barriers [Lin89]. Base [SM80, CM80a, Spy82]. Based [Bro83, NP85, Tak88, Tra84, AK81, BP83, CK89, Gei84, HS88a, IW81, Lei80, Mal87, MTK89, Mat89c, Muk89, NA89, OT89, RBP82a, SM92, TM89, YYS83]. Bases [AD83, Sik81, Won89]. basic [McL85]. Basis [BI85]. Batch [Wil80a]. Batched [Man86, LP89a]. batching [Wil80a]. Be [Hof83, LY86, App87, BCW80, Sch81a]. before [Bar83]. Behavior [Her83, Wei83, WFL83, Dob82, Dwe88, Pan86b, SN89]. Bell [McL85]. Ben [Sei88]. Ben-Or [Sei88]. Bendix [Mét83, Toy89]. Benign [Lam88]. Berman [CIR88]. Best [Weg82a, RG89b]. best-case [RG89b]. Beta [BH87, McG83]. better [Bod88a, McG83]. Between [Lum85, Pan86a, Tha85, Van81, ABM84, AAI87, Ata83, CW84, CH88b, Kem89, LB89, MC80, Mos82, Nij82b, Oli80, Oza83, Ray87, Roh88, Sch81b, Sze87, Yap83]. BFS [GM88]. Bi [Shi83]. Bi-Weighted [Shi83]. biconnected [CFHP89]. Bidirectional [Sam85, Bod88a]. Bifaithful [LT88a]. Bilinear [Pan86a, Ber82, CL86b, Pan81a]. bimodal [Ver88b, Ver89b]. Bin [Leu89, FV89a, Fre80, Lia80]. Binary [GT83, Hik83, Kor81, SS85a, Shi83, Sil81, TS81, TC84, YH87, ABM84, Bou85, Bre80, CYL88, CA86, Gor86, Gre89, GP87b, GH87, Kor82, LS82a, LP89c, Mäk88b, MPY87, Oli80, Pal87, Sch86a, Taz84]. Bini [Tsa81]. Bipartite [GR82, KC87, Ber84b, IM81, LK89]. bisection [Bui80]. Bisimulation [Dar89]. Bit [AD86, GMZ89, Bir85, Jak82]. bit-reversal [Bir85]. Bit-String [AD86]. bit-vector [Jak82]. Bitonic [HS85]. bits [Fus83]. BL [Bro80a]. Block [PM84, ZBS83, CMT88, EY87]. blocking [CM89, GT89]. blocks [NU88]. board [RV87]. Boas [Avi80]. bonds [PS89]. Boolean [APT79, APT82, AS88, BB88, BP87, BCW80, GGPS89, Hro85b, MP87, PS80, PS81a, SN88, Sky83, Tie84, Vys84, Weg83]. border [Lic84]. Bound [GK86b, LS86a, San80, Sei88, AK81, Avi80, Ban80, BR89, Bod88a, Bro80a, BP88, CP81, CCL89a, FL82, GST89, Hua88b, Kla87, Lev87, Lia80, LP89c, Man84, Ó’D88, SE87, Tie84, Tsi85, Vit85a, van80b]. boundary [AH89]. Bounded [McK84, TS86, BU89, BW87, Chr84, Kfo85, LJ86, LPP88, Par87, dV89]. bounded-depth [BW87]. bounded-reversal [Chr84]. Boundedness [Abi89, Mäk83, YJ83, Cha88, HRY87]. Bounding [LS86a]. Bounds [ĎSVT85, Har85, Hro85b, JT85, LM85, Llo80b, ARW89, BMS80, BS86b, CV89, DJ89, FG89, Llo80a, Mar86a, Mor87a, Mot82, Pan81a, RV87, Ram87, SV87, Tom81, Tur89, Ukk85, Weg82a, vW80a]. Boyce [Bis89, KM80]. Boyce-Codd [Bis89, KM80]. BPP [Lau83]. Bracket [Ryt86a]. Branch [LS86a, CCL89a]. Branch-And-Bound [LS86a, CCL89a]. branchings [TL83]. breadth [ZC87]. breadth-first-search [ZC87]. bridges [TC83]. broadcast [Szp86]. Bruijn [Sri88]. Bubble [Dob80, CLW80a, CLW80b]. bucket [Chl88, Chl89, Eas81]. bucketing [OIM84]. Buffer [AB81]. buffering [Wri87]. Bui [Day80a]. Bunch [Heh81]. Bykat [Ov80]. Byzantine [TC84].

C.F.L. [Sen81]. cache [Lop89, Vog89]. Calculation [KR85, NI83, AS86a]. calculator [Jos88]. Calculus [Mis88, Stå89]. can [App87, BCW80, Sch81a, Sze82]. Cancellation [BS86a, BO85]. Cannot [LY86]. canonical [Sch88]. capacitor [PM88]. Capacity [DRT81]. cardinality [Kan80]. Carlo [Ber82]. Carter [Sar80]. Case [JM86, KK85, Wil83a, Apo86, Bop89, CDS89, CS83a, KS87b, Lin86, Ov81a, PG87, Ram89b, Röh82, RG89b, Sun89, WI88]. cases [WS88]. category [MC88, MC89a, WS82]. Catenative [ER83a]. Cause [Cza88, Cza89, HS88b]. Cause-effect [Cza88, Cza89, HS88b]. CCC [LP89a]. cells [SSW87]. Cellular [Sto82, UMS82, Kob87]. Center [HS86]. Centering [CFHP89]. Centers [Car84]. centralized [RBP82a]. centres [Mur83]. Certain [SS84b, APT79, APT82]. certifier [RBP82a]. CFL [Ber86b, Gra87, Mai85]. chain [GS84b, LP82, Shi89]. Chains [Sup85, Vol85, BB87, BR89, Rot89, Yeh85]. Chandy [Ngu85]. change [EM84]. changes [BO86]. Character [Par84]. characteristic [Pan88]. Characterization [Bea89, Cza80a, HF86, Arb88, Bar80, GT84, Tou84, Vos88]. Characterizations [AMA88, Tag86]. Characterizing [Fei89]. chasing [Kie86]. cheap [Boo88]. Checking [Gun82, Man84, Wol89]. Checkpointing [Mag83, Dud83, VRL87]. chemical [Ev89]. Chip [ĎSVT85, SV87]. Chips [Jin84]. Choice [KK85, Sil81, PG87, Zem84]. choosing [WM89]. Chordal [Ede87, HL88, YG87]. chordality [dM88]. chromatic [Kuč89, Zer89]. Church [Boo81a]. Circle [LL84, Sup85, AC89, Ber88b, BM89b, Dam89, SS89]. Circle-Cover [LL84, AC89, Ber88b, BM89b, SS89]. circuit [Dam89, Gol80, Nar89, SN88]. Circuits [Ber83, Hro85b, Kei85, Yas81, BB86, BW87, Dun90, Huỳ87, Koc89, Mor87a, Par87, SR88b, Tur89]. Circular [RSR80, AH87, Boo80, RR89b, SH89, TW89]. circular-arc [AH87, RR89b, SH89]. circumscription [Kri88a]. Class [FS84, JM86, WB86, YG83, GS89, Jȩd83, LD88, RS89, SIN87, Sen81, Wag86, YD86]. Classes [AS86c, Oya84, Tha85, HH88, Kem89, Sar80]. classical [Stå89]. Classification [Lei86]. clause [McC88]. clauses [LS89, MM85]. Cleaning [Kas84]. clique [HL88, HL89, PY81a, SH89, Whi81]. cliques [AH87, HL88, KPS88]. clocks [Ray87]. closest [AH89, Dye80, HNS88]. Closure [JM86, Jȩd88, Mor88c, Cho81, Ebe81, Jȩd83, Jȩd87, Sze89b, Szw87]. closures [IK83, Khu89]. cluster [Mur83]. clustering [Mur83, Tar82b, Tar83a]. CNF [AB87b, CL88b]. co [EH88a, LS82b, FS88b, Ber86b, BHZ87, Gra87, Mai85]. co-CFL [Ber86b, Gra87, Mai85]. co-context-free [EH88a]. co-NLogSpace [LS82b]. CO-NP [FS88b, BHZ87]. Coalesced [Vit81]. Codd [Bis89, KM80]. Code [Han80, Fra83, OP85, Ric86, Won89]. Codefinitions [Las86]. Codes [CH84, Hea89, Kri82, MR81, Sie88]. Codeterministic [BP85a]. Coefficient [CRL85]. Coefficients [BM81]. Coffman [Llo81]. cographs [Bod89]. cogroup [AGvG89]. coherence [Lop89]. Colbourn [CM80b]. collection [App87, Van87a]. collector [Mor82, RM83]. collectors [Jok89]. color [Kuč89]. colorable [YG87]. colored [Pet80a]. Coloring [Fre84, GP87a, HL88, Nao87, Wil84]. colors [Nao87]. colouring [GIR88, Ste89, Wil80a]. Column [Bit85]. Columns [MA85]. combination [YYDH84]. Combinational [Yas81]. Combinations [NI83, AGS89]. combinatorial [Bru81, de 81]. Combinators [de 87, Bur82b, CDL87, Nos85]. Comment [Dun90, PG87, McL85]. Comments [AJV86, Avi80, Bal87, Day80a, DS79, Fou79, Pet91, Sak88, TTv86, TV87, TM89, VT90, Won89, AG88b, Gro91, Ov80]. Common [AD86, Apo86, Apo87, AB87c, Rei88, RT85b]. Communicating [Bro83, Bro87, RT85a, Van85b, YG83, Bro86]. Communication [Fle89, Har81b, MRS84, VU86, Van81, BO86, Hel81, KW82, LM80b, Mar85a, Mar85b, Mei81, RS87, Zöb87]. communications [Szp86]. Commutation [Dub85, NO88]. Commutations [MO87b]. Commutative [Dub85, BL87, CD86]. commutatively [Kri82, MR81]. compact [CS83a]. Compacting [Naw83, Mor82]. compaction [Mal87, RM85]. compare [Gru87]. comparison [Bar84, Won89]. Comparisons [Atk85, Car87b]. Compilation [TB84]. Compile [Gun82]. compiler [Toy89]. compilers [Pag80]. complement [Ban80, Sze89b]. Complete [AS86c, LMW86, LMW88, Ber81, Bod89, Dam89, FPT81, GGPS89, GH87, LS82b, MC88, MC89a, MW87, Rys83, Sch86a, Sho89b]. Completely [Van85b]. Completeness [BT82, MGKL88, Eis89, SV82, Kři86b].

Completion [FS84, Kel86, Mét83, Toy89]. complexities [BU89, CYC89]. Complexity [AS86c, BYV82, BS86a, BMS80, ER83b, ER83a, FY80, GK86b, JM86, LBB81, Pan86a, Ryt83b, Ryt86b, SM83, TS81, Ban80, Ber80, BV88, Bha80, BC87, BKV^{+}81, Bod88b, Bop89, BW87, Che86, CM89, CH88b, Col79, CM80b, Deh86b, ER81, EL81, FL87, Fle89, GMZ89, HH88, Hoc87, IMR80, JS87, Joh80, JF82, Kal84, Kan80, Kel87, Kla87, Kři88b, Kub87, LT80, LP83, Lis81, MC89b, Mur83, Ó’D88, PS89, Pan81a, RDM88, RS87, Röh82, RG89b, Ryt81c, Ryt82b, Sni81, Stå89, Sze89a, Tie84, Tur84, Ver88a, Ver89a, Vř’t87, Wey84, RG88, Rol89]. component [Koc87]. Components [NM82, Van81, Ima82, Pep87, Ram87, Tar82b, Tar83a]. Composition [Sha82b, TB84, Cza89]. compound [KS87b]. compressed [PK88]. Compression [DRT81, Jak82, Ric86]. Comput [Sar80]. computability [Kri88a, Win82]. Computation [BM81, GLPG82, Lum85, BL89b, Blu83, CC86, CCL89b, CH86, GPS87, Hag88, IK83, Kuč82, Kuč83, KM88, Mat89a, Mor88d, SY88, Szw87, WW87]. computational [Kan80, Sch82a, Stå89]. Computations [DFv83, TS81, Top84, ARG86, AG88b, CYC89, CL86b, Coo88, DS80, HS88a, JK85, LL88, Ryt88, TTv86, TM89, Vit80]. Computed [Nja85]. Computer [HS85, Las86, Deh86a, Gei84, Gos89, Gun89, Heh81, Min88, OT89, Yoo85]. Computers [MRS84, GS89]. computes [GM88]. Computing [AR80a, AC87, Ber84a, BI85, CH89, Der85, Ebe88, Er83c, GL80, Jos85, Pan88, PGR87b, PGR87a, Sho86, Val80, AH89, AR80b, BMRS88, Cro81, EH88b, GS83, HL89, IMR80, JM88, KNT87, Khu89, Ola89, Pet80b, Sho89a, SS83, Urb80, WS80b]. concatenated [Won89]. Concavity [DEGS84]. Concept [AS81, KK89b, KW82]. concepts [SN89]. concerning [CIRB87, CIR88, Cou84, Mal87]. Concurrency [Har81a, Kum89, HM88, RBP82b, RBP82a]. Concurrent [Bro80b, Mag84, Wol87, AK86, Bar80, Hol84, Mag87, Och87, Sis89, van89a]. conditional [EH88b]. cones [Mor88c]. conference [LLH89]. configuration [Gei84]. configurations [Cou81, Wil80b]. Conflict [Deb80, Kat84, HRY87, Mag87]. Conflict-free [Deb80, HRY87, Mag87]. conflicts [Kuč82, Kuč83, SY88]. Confluence [Van88]. conic [NA89]. conjecture [Boo81b, De 82]. Connected [HS85, IM88, Mey83, NM82, Van85b, Deh86a, GS89, Ima82, LS86b, OT89, Pep87, VV87]. Connection [SB89]. Connectivity [Gus83, BDD^{+}82, EH88b, Rei87b, Sri88]. Conquer [Wal84, BCC83, KO88, Ver88a, Ver89a]. Consecutive [ten80, Taz84, Tru80]. consensus [Bab87b, FR89, Mis89]. Conservative [Kir88]. consideration [Asa83]. Considerations [Oza83, IT88]. consistency [FL82]. conspiracies [Bes84a, Bes84b]. Constancy [GH81]. Constant [CRL85, GP87a, Hro85a, SS87, ER81]. Constant-Degree [GP87a]. constants [PT85]. constrained [Shi81, Sla82, van89a]. constraint [Mal87, Rom88]. constraint-based [Mal87]. constraints [AC87, Kan80, RD87]. construct [CS83a, Dev85]. constructible [Sze87]. Constructing [PP89, Sto81, Wel85, CYL88, CR89]. construction [CCS89, Fri88, Kut88, Mel87]. constructions [BSZ89]. consumption [Jan89]. Containing [Jȩd88, NO88]. Containment [Oya84]. Context [Har85, Mäk83, Ryt84, Sli82, AGvG89, AFG87, BCC83, Cho81, Ev89, EH88a, IJR88, Mai82, Mäk85b, NA81, Nij82a, Wag86, Yu89]. Context-Free [Har85, Mäk83, Ryt84, Sli82, AFG87, BCC83, Cho81, IJR88, Mai82, Mäk85b, NA81, Nij82a, Wag86, Yu89]. Contiguous [Naw83]. continued [CCL89b]. contour [WW87, Woo84]. contraction [MC89a]. Control [Tak88, AS86a, HM88, KPK89, Kum89, Muk89, PM88, RBP82b, RBP82a, San88, YYS83]. Controlled [CK86a, MG80, Lop89]. Controls [Sch84]. convergence [Asa83, FS88a]. converse [FI87]. conversion [ABM84, Mio84]. Convertibility [ter80]. Converting [Par84, Vid88]. Convex [Byk78, GS87, Kal86, SSS86, AH89, AK89, Akm87, And79, And80b, Ata83, Avi80, Ber86a, CW84, DS79, Dev80, Dub88, Fou79, Goo87, Gra72, Jar81, Kal84, Kla87, Mel87, Ola88a, Ov80, Rei88, Roh86, Roh88, RD87, Sch81b, SS83, van80b]. Convexity [DEGS84]. convolutions [Mat89a]. Cooperating [Fra89]. Coordinator [KK89b]. Cope [Van85b]. copies [Wol87]. Coprocessors [Mül87]. copying [Die89]. Coroutines [All83, Sam85, All89]. correct [Wyr83, Wyr84]. correcting [SS84a]. correction [CM80b, Kra80, Wil81a]. Correctness [Hol86, HL87a, Tir83, Ham87, McC89, Upf82]. Correspondence [Tha85, Kem89, Pan81c]. Corrigenda [Ver89a, Ver89b]. Corrigendum [And80b, GPB83, LMW88, MT87b, PGR87b, Tar86]. cost [CC86, Hua88b, MP88, Weg85]. Costs [MST84, Ber80, CGR87, JS80, Ryt88]. counter [CL86a, Deb80, HY82, Vř’t87]. Counterexamples [Toy87]. Counters [Miy84]. Counting [Hik83, HL89, KNI86, RH87, Wal82]. coupling [Mih89]. Cover [LL84, Sav82, AC89, BS89, Ber88b, BM89b, Cla83, GP86, SS89, Sho86]. covering [FPT81, KM80, RV87, WW89a]. covers [Got87a, Mou80]. Creation [Bud85]. credit [Mat89c]. Criteria [MSR85, Wey84]. criterion [Nij82a, UY88]. critical [Ray89a, Tur84]. Cryptanalysis [SS84b]. cryptanalytic [Hel81]. Cryptographic [San88, Sha83]. cryptography [AS83]. cryptosystem [Hel81, LLH89, LM80b, Mei81, SSA87, Yun85]. CSP [ABC87, Zöb87]. Cube [Mey83, Bha80, LS86b, OT89]. Cube-Connected [Mey83, LS86b, OT89]. Cubic [Sto81, Wil80b]. Cumulative [Har82]. Currying [Ida83]. Curves [SV86, BP83]. Customizable [Jok89]. Cut [FP82, NI89, Ski89, Whi81]. cut-set [Ski89]. cut-sets [Whi81]. Cuts [CC84, GI89, Hoc87, JK85]. Cutting [HS86]. Cycle [Shm83, Loi80]. Cycles [Mey83, Dil87b, Fri87, Kan83, LS86b, Pro89]. Cyclic [Mag85, Sat86, Atk81, FVG87, Wal83, ten80]. cycling [Gol87].

D [Bv82, Day80a, PG87, Žák84]. D0L [CL81, ER81, ER83b, SSA87]. D0L-T0L [SSA87]. D0S [ER80]. dags [Gab81]. Dangling [CK84, NM87]. DASD [Bra88]. Data [AD83, AMM84, And80a, DRT81, Kak85, LO89, LBB81, Mil83, Mor88b, Ric86, RTV89, SM80, VB86, CCS89, CM80a, Fie87, FMNT87, GS88b, HHS87, IW81, Jaz80, Jos88, KV81, Lee89, LW82, MTK89, MPY87, Mor88a, OS88, PT85, SO87, Sak88, Sik81, SY88, Spy82, Sun89, Szp88, Wey84, Won89, WSA84, YYS83]. data-dependency-based [MTK89]. data-driven [YYS83]. Data-Structure [And80a]. data/knowledge [Won89]. Database [HF86, Jaj85, PM84, RB82, Row84, CH88a, Che82, CMT88, HK81a, Leu82, RBP82b, SY88, Var81, Var84, VB86, Zan80]. Databases [FV83, Kel86, Lei86, MST84, Sch84, AGvG89, Kan80, LD80, RBP82a, ter80]. Datalog [Abi89]. Davis [GPB83, GPB82]. deadline [Jan89]. Deadlines [Fre83a, SY93]. Deadlock [Cza80a, FM87, Llo83, OR86, RT85a, JS80, Klu84, SPL^{+}89]. Deallocation [CK84]. deap [CCS89, Car87a]. Decidability [IR81, Gir86, Gra80]. Decidable [TS86, Sab80, Sab81, Sen81]. decide [HRY87]. decided [BCW80]. Deciding [Cha81, Hea89, IT89]. Decipherability [Ryt86b, AG84]. Decision [Cza80a, IMR81, Sei88, Cha83a, IT80, Jar81, Ram87, Var81]. declarative [MT86, MT87b]. decoding [Sie88]. decomposable [MO81, Ov81b, Ov81a, vW80b]. Decomposing [Sup85, BR89, LN88]. Decomposition [Kal85, BS89, Tak87]. Decompositions [DP83, BV88, Hou87, Var84]. Dedekind [CC84]. dedicated [Ber80, Kub87]. deduction [Bal87, Far82, Wil87, YD86]. deductions [YYDH84]. deficiency [Bal87, Wil87]. define [All89]. Defined [BS86a, GL80]. Defining [AS85, RB82]. definition [Bar83]. definitions [Pag80]. Degree [GP87a, McK84, BDQ82, Lei88, LPP88, Shi81, Sla82, WU88]. Degrees [GH81, AS86b]. Dekker [Mar86b]. Delaunay [Dil87a, Dil87b, Kir80]. Delauney [Lin86]. delayed [Erk80]. delays [RS87]. deleting [Ita88]. Deletion [Tsi88, Che86, Ov81a]. Demand [WFL83, HL87b, San89a]. Denning [Bv82]. Denotational [Bro83, Bro86, Gor80]. Density [MG80, LFQ^{+}81]. Departure [Les83]. dependence [Kor87a]. Dependencies [AM84, FV83, BB88, Cas83, CH88a, GT84, GM82, MGKL88, Sho86, Var81, Zan80]. Dependency [HF86, Kat84, GT84, KPK89, MTK89, MGKL88, Tag86]. dependent [SR88a]. Depth [Rei85, Sav82, Yas81, Awe85, BW87, Cid88, Dym88, LMT87, Mor87a, Par87, SV88, Sha88a, SIM89, Weg82a, Weg83, dI89a]. Depth-First [Rei85, Sav82, SV88, Sha88a, dI89a]. depth-first-search [Awe85, Cid88, LMT87, SIM89]. Deques [GT86]. Derivation [DFv83, Las86, Mäk84, Pet80b, Hem80, Leu84]. Derivations [RAK85]. derived [Bar80, MGKL88, Mor87a]. descent [Bar83]. Describing [Bro87, Sli82]. description [HS88b, SR88a]. Design [CC83, MA80, Tho85, CP82, Gei84, Ray89b, VB86]. designing [LLH89]. designs [CC88a]. desynchronized [Kob83]. Details [Nja85]. Detect [Kak85, RT85a, YD86]. Detection [DFv83, GT83, Shm83, Top84, YG83, ARG86, AG88b, DS80, EO87, HS88a, Hua88a, Hua89, JS80, Mat89c, MF81, Rei88, SPL^{+}89, TTv86, TM89, Tse87]. determinacy [WSA84]. Determinancy [Mos82]. Determinant [Ber84a, GP89, Pan88]. determination [Avi80, BES82, Bou85, van80b]. determine [LP82]. Determining [Ber86a, Pep87, Akm87, Gra72, RW87]. Deterministic [IR81, IM81, LY86, Mon84, UMS82, Cho81, Fro81, FV89b, Kin80, Mic81, Par87, Yeh85, Yu89, Zip89]. Deutsch [Ham86]. developing [Rem84]. development [BW80]. deviation [Win82]. Devices [Sch83b, TV81]. diagonal [Lin89]. diagonals [SM87]. diagram [OIM84]. diagrams [Lin89]. Diameter [McK84, BDQ82, EK89, EG89]. difference [Urb80]. Different [MSR85, CDS89]. Differential [WB86, Bui77, Bui79, Day80a]. difficulties [Leu82]. diffusing [DS80]. Digital [Lea88, Reg81]. Digitizing [SSW87]. digraph [Sys82]. Digraphs [IN84]. Dijkstra [Der85]. Dijkstra-Like [Der85]. dimension [Lic86]. dimensional [AN86, Aur86, Cha88, Fre80, GH87, IT80, KO88, Li88a, LT88b, NI80, pW82]. Dimensions [EM85, And79, And80b, Byk78, DS88, DS79, Fou79, Ov80, Pap85, Sle80]. Diophantine [Row84]. Direct [All89, Bou85, Bir85, Rus87, Toy87].

Directed [FP82, Pit87, EG89, GM88, Ita88, JK85, NI89, Nar89]. Direction [AN86, LMW86, LMW88]. Direction-independent [AN86]. disc [Ó’D88]. Discrete [ĎSVT85, MS86, Kaw80, RW87, Ver88b, Ver89b, Wu89]. Discriminators [Sil81]. discs [SY84a]. disjoint [Gus83, Iwa89, Per83, TL83]. disjointness [Man84]. Disk [Bit85, Hof83, CV89, Cha85]. dissection [Gil88]. Distance [Lum85, Ata83, CW84, CR89, Hua88b, LP89c, Mäk88b, Pal87, Sch81b, Sho89a]. Distortions [Nja85]. Distributed [AR80a, AS81, ARG86, AG88b, DFv83, HC85, LR86, LBB81, MF84, Mor85, TTv86, Top84, AR80b, Awe85, Bab87b, Ber87b, Bod88a, CM80a, CP81, CP82, Cid88, Fra89, GMZ89, HS88a, HZ87, Hua88a, Hua89, KPK89, KT88, KW82, Kut88, LY87, LMT87, Lav89, Mat89b, Muk89, Ran83, Ray89a, Ray87, Ray89b, RBP82b, RBP82a, SSA89, SM92, SIM89, SPL^{+}89, TV87, TM89, VT90, Ver87, ZC87]. distributing [Eic83]. Distribution [Har82, Ric83, ER81, KLV87, LLH89, Mat89c, Sha82a, SY88]. distributions [FS88a]. Distributive [van80a, AN80, Han86, HL87b]. Divide [Wal84, BCC83, KO88, Ver88a, Ver89a]. divide-and-conquer [BCC83, KO88, Ver88a, Ver89a]. Division [BP85c, Asa83, BP87, Dem82, SR88b, Vys83]. Divisors [SS85b]. Do [Kat84]. Does [BHZ87]. DOL [ER83a]. domain [BSZ89]. Domains [Bou84, BS86b]. domatic [Ber88a, RR89a]. Dominating [Ber84b, BYV82]. domination [Ber86c, Kei86, RR88a, RR88b]. dominator [Lei88]. dominators [PGR87b, PGR87a]. domino [VRL87]. domino-free [VRL87]. Done [LY86]. door [IW81]. Double [Car87a]. Double-Ended [Car87a]. down [FV89b]. DPDA [Oya84, Vit80]. DPDAs [Cou81, Cou84]. Drawing [Har89, SV86, Sto82, KK89a]. drift [Ray87]. Driven [Ryt86a, Dym88, OP85, YYS83]. dual [FV89a]. duality [LC85, Mak81]. duals [Kan83]. duodirun [Zhe83]. During [NP85, Sha82a]. Dynamic [Güt85, HP85, KL88, LT88b, Reg81, Shm83, Bra88, FB88, FI87, GS89, Gos89, KV81, LW82, Rei87b, Ryt81a, Slu85]. Dynamics [RB82]. Dynamization [vW80b, MO81]. dynamizing [Ov81b].

E0L [RAK85]. ears [Hig82]. east [GS83]. easy [Kal86, Kuč89]. ECDF [Deh86a, DS88]. ECLR [SIN87]. ECLR-attributed [SIN87]. Edge [GT83, BP89, Ber81, EH88b, GIR88, Gus83, Miy88, Sch86a, Shi89, TL83]. edge- [Miy88]. edge-colouring [GIR88]. edge-connectivity [EH88b]. edge-disjoint [Gus83, TL83]. edges [Ita88, LS86b]. EDISON [Dai86]. EDISON-80 [Dai86]. edit [Hua88b]. edit-distance [Hua88b]. Editor [Pet91, Kri82, Žák84]. effect [Cza88, Cza89, HS88b, MT87a]. effectiveness [Dar89]. Effects [KT88, Dud83, Lam88]. Efficiency [MS86, CGPD89, CGPdS89, MPY87]. Efficient [AT87, AB87c, BM81, Bit85, BGLY81, CC86, CA86, Der85, Dob80, Dun90, FM87, GLPG82, HL88, JM86, KF85, Koc89, PPK88, SR88b, Sik81, Vis85, WS80a, vN82, And79, And80b, AFK88, Bur82a, Col88, FP87, Gol89, Gol80, Gra72, GL89, IT89, LMT87, Mat89b, RGG89, Roh88, Ryt88, SIM89, ZM87]. efficiently [GS89, Sho89a]. eigenvalues [BES82]. Eilenberg [BI88]. Election [LMW86, LMW88, Tau89]. elegant [Vid89]. elegantly [HNS88]. element [Urb80]. Elements [Atk85, Deh86a, Ove81]. Eliminating [Gor86]. Elimination [GR82, HL89, Lic86]. embedded [Pro89]. Embedding [HK88, PW87, Sha83, SY84b, CKV87, Gre89, Pla85]. Emde [Avi80]. employing [Har81a]. Emptiness [AK85, ER80]. emulations [Bod88b]. enclosing [Pro89]. Encoded [And80a]. Encoding [AT87]. encryption [GY87]. Ended [Car87a]. endofunctors [MC88, MC89a]. enlargement [Jȩd83]. entries [Ray89a]. Enumerable [CC84]. Enumeration [BM81, Lau85, BW80]. envelope [Her89]. environment [HL87b, LO89]. environments [Kaw87]. Equal [Hik83, Weg82b]. Equality [Ros86, SM80, EH88a]. equation [Sch88, Urb80]. Equations [WB86, AC88, Bui77, Bui79, Day80a, EY87, FY80, Kra84, Lic86, Pan88, PS80, PS81a, Sri87, WS82]. equidistribution [Fus83]. Equivalence [AK85, BFJM87, BCW80, CK86b, IR81, Jaj85, KW84, BMS80, EO82, GM82, Sab80, Sab81, Sen81]. Equivalent [GP86, SS85b, Tra84, Cou81, GT84, Kri82, MR81]. EREW [HR89]. Errata [Llo80a, vW80a]. Erratum [APT82, Bes84a, Bui79, CIR88, Kuč83, Mar85a, PS81a, Wyr84, YWL85]. Error [CV89, Kra80, CC86, MF81, SS84a]. error-correcting [SS84a]. Ershov [Kot89]. Estimate [Bou84]. estimates [BS86b, Dob82]. Estimating [MST84, ZBS83, BV88, Che82, Zer89]. estimation [Ber82]. Estimator [Pit87]. ETOL [YWL82, YWL85]. Euclidean [DF84, EK89, GK86b, Ima82]. Eulerian [Ebe88, PY81b]. evaluated [Sch81a]. Evaluating [Van85a, CMT88]. Evaluation [Jak82, Mag84, Mag85, Alb89, CD89, FH87, GP89, GS89, Han80, Mag87, MKK89, SN88, Sik81, Szp88]. evasiveness [ACK89]. Even [Der85, Tak88, Ebe80]. Evenly [Wal84]. Events [Eva86]. Every [Ber86b, Sze82]. Everywhere [TS86]. Exact [CH86, Sni86]. example [Deb80, HY82]. exceptions [All89]. exchange [AS83, Bir85, LS86b, SR81]. Exclusion [KP84, Mag85, Mar86b, AJV86, DT80, FP87, MB85, Mor79, Pet81, SM92]. execution [BF87a, BF87b, CP81, DL88, Dud83, PPK88, Wol87]. Exercise [BH87, Kes88]. Exit [Wil83c]. Expected [LBB81, Mur83, Dev85, Dob82, GJ82, JS80, KNT87, Pan86b, TV81, Wil84]. Expected-time [Mur83, KNT87]. Experiments [Eva86, MNS84, BS86b]. Explicit [EY87, Zip89]. exploitation [BDS82]. exponential [BA80, BP88, EG89, Sha82a]. exponentially [Mic81, Sky83]. exponentials [Cal87]. Exposure [RT85a]. expression [GS89, Ten87a]. Expressions [DK84, Jȩd88, Van85a, Cas83, GM82, Ryt89, SN88, Sch88]. Extended [BS86a, BO85, MW87, HL87a, Wil81b]. Extendible [BK84, Tam82, Yao80, Kum89, Wee88]. Extending [Pit87, San80, TC84, Cia86]. Extension [MF84, Mar87, FZ82, MZ82, Tom81]. Extensions [TB84]. external [LBM88, Wri87]. extra [GJ82]. Extreme [CS82, EM85].

F [Bv82, Mor81]. Facilities [Har81b]. Facility [CW86]. Factorial [BU89, dV89]. factors [dV88, Ebe80]. failure [SR88a]. failures [Gos89, Tau89]. Fair [FP83, MB85, RAK85, GP88]. Fairness [Bes84a, Bes84b, Cza80a, Zöb87]. families [GGPS89, Tag86]. family [Kri82, MR81]. Fan [Hro85b]. Fan-In [Hro85b]. farthest [AK89]. Fast [BD80, BP85c, CCL89b, CRL85, CL86c, Güt85, LBB81, Lum85, MNS84, Sch82b, Sha82b, Sie88, Toy89, Wu89, van80a, ABM84, AG84, BS87, BP83, Ber82, Car86, Coh83, CR89, Dye80, IM81, II86, Llo89, MM85, MG89, Mit88, Nao87, OIM84, Per83, Wat81, Zhe83]. Faster [Sch83b, App87, Flo91, Meh88, Roh83, Sch81a, TL83]. Fault [DF88, IM88, Var89b, BJ80, Kut88, ÖA88]. Fault-Tolerant [IM88, DF88, Var89b, BJ80, Kut88]. Faulty [Van85b]. FD [Got87a, Vos88]. FD-covers [Got87a]. feasible [HH88]. February [AJV86]. Feedback [Wät85, BDP88, Dwe88]. few [Yap83]. fewer [CYC89]. Fibonacci [CH89, Er83c, GL80, MR84, Pet80b, WS80b, de 81]. Field [Gun82]. fields [Cho82, IT89]. FIFO [Bur82a]. file [CC88a]. Files [Lei82, Man86, Won89]. filling [BP83]. filters [Mar83, PM88]. find [BM89a]. Finding [AH87, Ber83, EM85, Goo87, Her89, Hig86, Ima82, IMM81, Ita88, Kei85, MS86, PS88, RSR80, Sch81b, AK89, BD80, BYV82, Bod88a, Bod88b, CM89, CL88b, Dev80, Dub88, Fri87, GT88, GT89, HL88, JK85, KST82, Lav89, Llo89, Loi80, MP88, Mar86a, TL83, TC83, WW89a, Wat81, Whi81, Wit88]. Finitary [Žák84]. Finite [AB81, BI85, Cza89, LY86, SV84, YG83, AK86, Byk78, Cho82, CLW83, DS79, FG89, Fou79, Fre81, Gra72, Hea89, HS88b, IT80, IT89, Mic81, NI80, Ov80, RH88, Ryt89, Suz88, Sze82]. Finite-State [YG83, AK86, FG89, Suz88]. Finitely [LT86, MO87a, Tag86]. firing [Cul89]. First [Rei85, Sav82, Sza87, Awe85, Cid88, LMT87, Mar86a, Miy88, SV88, Sha88a, SIM89, Sho89b, ZC87, dI89a]. First-Order [Sza87]. fit [Ram89b, RW87]. Fitted [SM87]. Five [Fre84, Nao87]. Five-Coloring [Fre84]. Fixed [CKV87, LNS82, NU88, AC88, Bod88b, Cha83b, EM84, MC88, MC89a, Ram87]. fixed-point [AC88]. fixed-radius [Cha83b]. Fixed-sized [NU88]. flag [Gor86]. flexible [DF88]. floor [Sch82a]. flow [CM89, GT89, Has81, Jan89, Jaz80, NI89, Wey84, WSA84]. flow-shop [Jan89]. flowgraphs [Tse87]. flows [NI89]. Floyd [TST83]. fly [BK88, RM83, Van87a]. Fold [HS85]. Folding [CC83]. folk [LNS82]. Forest [SM86, Akm87, Dub88, Kut88]. forests [LP82]. form [ABC87, Bis89, YWL82, YWL85]. Formal [CK86b, DK84, Hem80, JB84, Upf82, BI88, Kaw80]. Format [Mil83]. forms [Ber82, Bis89, CL86b, KM80, OSW81, Sni81]. formula [CL88b, Weg83, Van87b]. formulae [Min88]. formulas [APT79, APT82, Coo88, JS87]. formulations [FB88]. Forward [Wol89]. founded [Mor89]. Four [San80]. Fourier [ĎSVT85, Nja85]. FP [Ida83]. fractional [Man82]. fractional-on-line [Man82]. fractions [CCL89b]. Frederickson [SY93]. Free [Car84, Dub85, ER83b, Har85, Kat84, Mäk83, Ola88b, Ryt83b, Ryt84, AJV86, AFG87, BYV82, BCW80, BL87, BCC83, Car83, CD86, Cho81, Deb80, EPS82, EH88a, FR89, HRY87, IJR88, Mag87, Mai82, Mai85, Mäk85b, Mor79, NA81, Nij82a, Sli82, VRL87, Wag86, Yu89]. Freeness [Cha81]. frequency [CV89]. Full [Sto81]. Fully [LR86, Hua88a, Sze87]. Function [Cha81, Hug86, Hua88b, IW81, Kes88, MZ82, Sch82a, Tie84, Ver88b, Ver89b, dV89]. Functional [AM84, Bro87, Clo83, FV83, Bur82a, Bur82b, Cas83, CH88a, Dwe88, GS88b, Sho86, Zan80]. Functions [Har82, LS86a, SS85b, SY84b, AMA88, BU89, BW87, Bur89, Dun90, EH88a, GGPS89, GL80, Koc89, Mar87, MP87, Pob86, Sar80, Sze87, Wat88, Weg82a, Weg83, Van87b]. Further [Gro91, Ov80]. Future [Eva86].

G [AG88b, Bal87, SY93]. gallery [Nta86]. game [Lou80]. games [LT80]. Garbage [App87, Jok89, Mor82, RM83, Van87a]. Gauss [RTV89]. Gaussian [Lic86]. General [Ili85, KR85, Mil83, Ove81, San80, BCC83, CLM89, Gol87, KK89a, TC83, WS80b]. generalisations [Leu84]. Generalization [Mar86b, TST83, Fra83]. generalizations [BB86, SV82]. Generalized [Dem82, Mor87a, OR86, ACK89, DK88, HK81b, Taz84]. Generate [McK84, Sat86, Sha82a]. Generated [LT86, RS89]. Generating [vR88, AGS89, CL80, EM84, JYP88]. Generation [RD87, Sem81, Zak82, BK88, Pag80]. generators [OP85]. geometric [Kři86b, LC85, WK88]. GF [CL80]. given [BDQ82, Har89, Jan89, RS87]. Global [Mat89c, Mor85]. Glushkov [Zem82]. Goldberg [Iwa89]. gon [DJ89]. Good [Jin84, Sni86]. GOTO [Lif84]. Graham [GS87, Llo81]. Grammar [CK86a, Gru84, OSW81]. Grammars [Mäk83, Mäk86b, RB82, Tar82a, AB87a, Boo81a, Ev89, Kra80, KR84, Mäk85b, NS86, Nij82b, SIN87, SN87, Sli82, TW89, Wan80, Yod84, Łuk82]. Grammatical [Bud85, MN82, Tak88, Mou80]. Grant [Bis84, PP80]. Graph [Ata84, Lau85, Wel85, Arb88, BB86, Bha80, BKV^{+}81, CFHP89, Ede87, EH88b, GM88, Ima82, Khu89, Kie86, Kor87a, KR84, Mit88, Rei87b, Ryt88, SPL^{+}89, Sri88, Sys82, Szw87, Tsi85, Tur84, WW89b, Wil84, Zer89, dM88, Van88]. Graphs [Ber83, CS83b, Fre84, GR82, GP87a, JM86, Kei85, KC87, Kuč89, Ola88b, Pit87, Sto81, Sup85, AH87, BYV82, BDD^{+}82, BDQ82, Ber84b, Ber86c, Ber88a, BCW80, Bod88b, Bod89, CS82, Dam89, Ebe80, EMM89, Flo91, Fri87, GIR88, Hat86, HL88, HK88, IK83, Ita88, KK89a, KNT87, Kei86, Kři88b, LFQ^{+}81, Li88a, LS86b, LP82, Meh88, Nao87, Nar89, OR89, PY81a, PY81b, Pro89, RR88a, RR88b, RR89a, RR89b, RS89, Sha88a, SE87, SH89, Ski89, Ste89, YG87]. Gray [Ric86]. Gray-code [Ric86]. Greedy [Hof83, Kor81, Cla83, Gol89, Lev87, Lin86]. Grid [TB84, Gre89, Ryt88]. grids [Nta86]. Group [BI85, Bv82, DS81, DMS81]. Groups [Ili85, McK84, MO87a]. Growing [Kor82, Lau87]. growth [BU89, dV89]. guarantee [MC89a]. Guarded [AS81]. guided [Bli88]. Gupta [TTv86].

H [Žák84]. Håstad [Mor87a]. Hadamard [ĎSVT85]. Halfplanar [EW86]. Halin [GIR88]. halving [Mir87]. Hamilton [Fri87]. Hamiltonian [Ber81, Ber83, BB86, Dam89, Dil87a, Kei85, Nar89]. Handling [Tsi88]. Hanoi [Atk81, BL80a, SS84a, Wal82, Wal83]. Hard [RT85a, Huỳ87, Ko82, Rob80, Sli82, Weg85]. harder [Mor81, Rom80]. hardest [Ryt81b]. hardness [SY84a, Wat88]. Hardware [RSD81]. Hash [LL85, Pag85, OT89, Sar80, Wog89]. hash-based [OT89]. Hashing [BK84, CC88b, Reg81, Tam82, Vit81, Yao80, BL89a, Kum89, OS88, Ram89a, Wee88]. Hausdorff [Ata83, Sho89a]. HCB [PK88]. Head [LY86, Mon84, Vit85b, TV81]. Headed [Hof83]. Heap [Car87a, GT86]. heaps [Fri88]. Heapsort [Car87b]. Height [Reg81, Ric83, PK88]. Height-Balanced [Ric83]. Heuristic [AR80b, BP83, WI88]. heuristical [Zer89]. Heuristics [DI89b, CGPD89, CGPdS89, CSO89, FV89a, KV81, Mak88a]. Hexagonal [Kři86b]. hierarchic [Mur83]. Hierarchical [AT87, Jak82, Tar82b, Tar83a]. hierarchies [LD88]. Hierarchy [Jȩd88, Son85, Bab87a, BT89, CH88b, HM88, Lau83, San88, San89b, Sze89b, pW82]. High [LFQ^{+}81, Vit80]. higher [Bur89, Dob82, Pan86b]. higher-order [Bur89]. Hilbert [Lea88]. histories [Bar80]. History [CC87]. Hoare [BT82, Tak87]. homomorphism [Spy82]. Horizontal [DP83, RM85]. Horizontally [LJ86]. Horn [AB87b, JS87, LS89, MM85, Min88, YYDH84, YD86]. Horner [GPS87]. Hsu [Apo87]. Huffman [CH84, Sie88]. Hull [GS87, AH89, Avi80, Byk78, DS79, Fou79, Goo87, Gra72, Jar81, Kal84, Kal86, Kla87, Mel87, Ov80, van80b]. Hulls [SSS86, And79, And80b, Dev80, SS83]. Hunt [Apo86]. Hunt-Szymanski [Apo86]. Hybrid [CK89, MA80, van80a, Röh82, Yap83]. Hybridsort [Hag89]. Hypercube [CKV87, CR88, CL88a, ÖA88, PS88]. hypercubes [EG89]. hypergraph [dM88]. hypersphere [GZW85, WW89a]. Hypothetical [AD83]. Hysterical [MS81].

I/O [Dwe88, SM87]. ibid [AG88b]. idealized [TST83]. idempotent [Bea89]. identical [BF87a, BF87b]. Image [AT87, MPY87, Sho89a]. Images [GT83]. imbalance [Bra88]. immediate [MF81]. immunity [Bal86]. immutability [Hea89]. Implement [Car87a]. Implementation [Bar81, CRL85, FR89, Hoo86, VU86, WS80a, AG84, Asa83, BS87, Bar83, Bur82a, CLM89, HS81, Min88, San88]. implementations [Lau80]. Implementing [GS88b, AG81, Sto86]. Implication [HF86, Got87b, Stå89, Vos88]. implicational [Tag86]. Implicit [AMM84, OS88]. implies [KL87]. important [Jȩd87]. impossibility [MW87]. Improved [HV87, IS86, JT85, NP85, Bro80a, Bui80, Cha83b, CK89, GM88, Lev82, MTK89, Par87, Tar83a]. improvement [CL80, Han80, Iwa89, Jaz80, MF81]. Improving [Apo86]. In-situ [FVG87]. in-the-past [Fie87]. Inaccessible [Lei86]. Includes [AB87b]. including [GS89, YD86]. Inclusion [FV83, Hoo86, CL81, Kin80]. incomparable [Gru87]. incompleteness [Ngu85, Sza88]. Increasing [CGPD89, CGPdS89, Fus83]. Incremental [Mey85, Alb89, ISHY88, Kal84]. Independence [Har85, KOR87b]. Independent [Les83, AN86, BWD84, CL88a, JYP88, JM88, Kub87, KOR87b, Mar86a]. Index [Man86, Oza83]. Indexed [Lei82]. indices [Mos82]. indirection [Van88]. individual [RD87, Udd86]. induced [Miy88]. Induction [Mil83, Mor89, Rao88]. inductive [Tou84]. ineffable [Eme87]. inequalities [Mos82]. Inevitability [Och87, Eme87]. Inference [Row84, Tak88]. Infinitary [Sao84, Žák84]. Infinite [BP85a, Gra87, Jȩd88, Mey83, AFG87, Gir86, Mai85, Mos82, dPV88]. infinity [Sal89]. Inform [And80b, AJV86, AG88b, APT82, Avi80, Bal87, Bes84a, Bin82, Bv82, CIR88, CM80b, Day80a, Dun90, Flo91, Fou79, GPB83, Gro91, Hel81, Kri82, Kuč83, Llo80a, LMW88, Mei81, MT87b, Mor81, NI80, Ov80, Pet91, PG87, Sak88, SY93, TTv86, Tar86, TV87, TM89, VT90, Ver89a, Ver89b, Wyr84, YWL85, vW80a]. Information [AB81, CV89, HM88, TST83]. Inherently [Rei85]. Inhomogeneities [AS86b]. Init [Fre83b]. initial [Tau89, WM89]. initialization [Per83]. Input [Dym88, Ryt86a, Hoc87, Wri87, YYDH84]. Input-Driven [Ryt86a, Dym88]. insecure [Kaw87]. Insecurity [Sch84]. Insertion [KK83, CC86, Dob82, DLF87, Ov81a, Pan86b]. Insertion-Safeness [KK83]. instability [Bin82, Tsa81]. instance [HLY80]. Integer [Fre83a, LM85, LR84, SY93, Vys83]. Integers [Par84]. integral [Gun89, NI89]. Integration [WB86, RG89a]. intelligent [Bru81, Wol89]. interactions [Fra89]. interactive [BHZ87, FL82, FS88b]. Interchange [Mai82]. interconnection [BFJM87, DF88, LFQ^{+}81, Par86, Var89b]. interdistance [Sal89]. Interesting [And80a]. interlocking [Llo89]. Internal [Erk80, Day80b]. Interpolation [SS85a, BL80b, GP87b]. Interpolation-Binary [SS85a]. interpretations [Žák84]. Interpreting [FI87]. Interprocedural [RG89a]. Interprocess [Har81b]. interprocessor [RS87]. Intersection [Güt85, EM81, EMK82, ER80, Ima82, RDM88, Rei88, Wag86]. Interval [Ber83, HV87, Kei85, BL89a, BB86, Ber86c, Ber88a, Bod89, Gab81, Kei86, RR88a, RR88b, RR89a, Sha82a]. intervals [LB89]. intricate [SS81]. introducing [LBM88]. Introduction [Dv82, Wat80]. intuitionistic [Cia86]. Intuitive [Mag83]. invalidating [NM87]. invariance [Mor89, Pet91]. invariant [Pet91, VT90]. invariants [Rem84]. Inverse [KW84, GP89]. Inversion [BP85c, DK88, FVG87]. Inversions [Har82]. inverting [Hua81]. involving [Mak81]. IO [RG88, Rol89]. IO-complexity [RG88, Rol89]. irreducible [CL80]. isometric [NA81]. isomorphism [BMS80, Gro91, LK89, Mäk89, Mit88]. Isothetic [Güt85]. isotonic [Wan80]. issues [MGKL88]. items [KLV87]. Iterated [FR84, Ber86b, Chl89, Pan81b]. Iteration [Wal83]. iterative [DK88, Gin82, Jaz80].

J [AG88b, Bal87, Sar80]. Jobs [BB83, Ber80, CL88a]. jogger [Bir81]. Join [NP85, CP81, CK89, GT84, LD80, MGKL88, OT89]. joins [CGTW89]. joint [Sha82a]. Jordan [RTV89]. July [Hel81, Mei81]. Jump [GP87b, EY86]. Jumps [SV84, All89].

Kapur [PG87]. Karmarkar [Blu86]. Key [AS83, LLH89, SSA87, YH87]. keyless [AS83, Yun85]. keys [Bv82, CS83a, DS81, DMS81, Weg82b]. kinematic [Sch88]. Kleene [AB87a]. Knapsack [Sha83, Zem84]. Knapsacks [Wil83b]. knowledge [FI87, GS81, Won89]. Knuth [Bar81, Mét83, Pit87, Toy89]. Kochol [Dun90]. Krishnamoorthy [PG87]. Kronecker [SM83].

L [CIR88, Sar80, NS86]. L-attributed [NS86]. Labeling [Der85, Sys82]. ladder [Sha89]. Ladner [BD82]. LALR [BL89b]. Lamport [Vid88]. Language [Bro83, Gra87, JB84, Dai86, Dem82, Har81a, Lat83, MT86, MT87b, Pag80, Ryt81b, van89a]. Languages [ER83b, ER83a, GH81, Har85, KW84, LT86, Ryt84, Ryt86a, Sao84, Tak88, Wät85, YJ83, AFG87, Bar83, BU89, Cho81, Dym88, ER80, ER81, EH88a, FS88b, Fre81, Fre83b, Fro81, Hea89, IJR88, Jȩd83, MC89b, NA81, Nij82a, RG88, Rot89, Ryt81c, Sch86b, Sze88, Ten87b, Wag86, pW82, WW89b, YWL82, YWL85, Yu89, Žák84, dPV88, dV89]. LaPadula [McL85]. Large [Jin84, RSD81, BDQ82, Won89]. Latin [BMS80]. lattice [Pal87]. layer [CM89, Dev85]. Layers [Tho85]. layout [SR81]. layouts [BC87, Bre80, GH87, LS86b, Sch86a]. Lazy [Boo88, Han80]. LC [CCL89a]. Leader [Tau89, Bod88a, Lav89]. leader-finding [Lav89]. leading [Fus83]. least [Boo80, CC86]. least-cost [CC86]. Leeuwen [AG88b]. leftist [Kem87]. Legal [IN84]. lemma [BM82, Del83, Mor87a, Yu89, Mai82]. length [BB87, BWD84, Gre89, KPS88, NO88, Sch86a]. length-reducing [NO88]. lengths [BP89, BC87]. Lenstra [Paj80]. lET0L [WU88]. Lett [And80b, AJV86, AG88b, APT82, Avi80, Bal87, Bes84a, Bin82, Bv82, CIR88, CM80b, Day80a, Dun90, Flo91, Fou79, GPB83, Hel81, Kri82, Kuč83, Llo80a, LMW88, Mei81, MT87b, Mor81, Ov80, Pet91, PG87, Sak88, SY93, TTv86, Tar86, TV87, TM89, VT90, Ver89a, Ver89b, Wyr84, YWL85, vW80a]. Lett. [Gro91]. Letter [Car84, Kri82, Pet91, Žák84, Car83]. level [Gru84, TST83]. Levelling [Far80]. levels [PP80]. lexical [ISHY88]. Lexicographic [MO87b]. lexicographical [Roh83, Sem81]. Lexicographically [Boo80, KR85, Miy88, Sho89b]. liberal [Mor87b]. Like [Der85, Kri86a, Ten87b]. Limited [AR80a, DRT81, Mül87, Naw83]. Limits [AK86]. Line [Güt85, IMR81, Lum85, Mat88, Slu87, Tak86, Wel85, ARW89, Die89, Her89, IK83, ITT82, Lia80, Man82, Mat89a, Mel87, NI80, RW87, Sys82]. Linear [Auf88, Cla86, CRL85, CL86c, Fre84, GLPG82, HS86, Hro85b, IMM81, Jar81, JM86, LR84, Pan86a, RR89a, RSR80, Ryt84, SS87, Tak88, AK89, AC88, APT79, APT82, Ata83, Bur82b, CDL87, EW86, EMM89, EY87, Gab81, GT88, GJ82, GPS87, HK88, IJ89, KNT87, Klu84, Li88a, Lic86, LP82, Lyo88, Mak88a, MU84, Mei85, MZ82, Min88, Ola89, OS88, Pan88, Pep87, RW87, Sab81, Sch89, Sha88a, Sha89, Sho89a, Sla82, Sri87, TV81, Wag86, WW89b, Zem84, ten80]. linear-processor [Sha88a]. Linear-Time [Fre84, IMM81, APT79, APT82, Gab81, GT88, MU84, Min88, Ola89]. Linearity [YJ83]. Lines [Sto82]. linked [Weg82b]. LISP [HM81, TJP87, Sto86, YYS83]. Lisp-based [YYS83]. List [Llo80a, Llo80b, vW80a, Hem80, Lev82, TJP87, Weg82b]. list-searching [Lev82]. Listing [Hik83]. Lists [Hug86]. literal [BP88]. Liveness [AS85, RM83, Rub87]. LL [MF81, NS86]. Load [Bra88]. Local [Kaw87, RM85, VRL87, WM89]. locality [TJP87]. Locally [ER83a, CC86]. location [LP89a, WK88]. Log [Er83c, Wil83a, GL80]. Log-Logarithmic [Wil83a]. Logarithmic [Wil83a, BV88, BP87]. Logic [BT82, HP85, LBM88, SY84b, Sza87, Yas81, Cia86, Far82, FI87, Lei80, Sab80, Tak87, VD87, Weg82a, YD86, ZLS87]. logic-termal [Sab80]. Logical [Clo83, Mor85, Bur89, Ray87]. logics [FI87, McC89, Ten87a]. Long [LS86b]. Longest [AD86, Apo86, Apo87, EMM89, Mir87, RT85b, dI89a]. Longest-Common-Subsequence [AD86]. look [FV89b, Shi81]. look-ahead [FV89b]. lookahead [BL89b]. Loop [Kas84, Ryt83b, Rem84]. Loop-Free [Ryt83b]. Looping [HP85, SST88]. loops [Mor89, OP85]. loss [BF87a, BF87b, KT88]. lossless [LD80, Var84]. Lower [ARW89, BS86b, ĎSVT85, FG89, GK86b, Har85, Hro85b, JT85, LM85, Mar86a, Sei88, Tur89, AK81, Avi80, Ban80, BR89, Bod88a, Bro80a, BP88, DJ89, FL82, Hua88b, Kla87, Lev87, Lia80, Man84, Mor87a, Ó’D88, Pan81a, Ram87, SE87, SV87, Tie84, Vit85a, van80b]. LR [CC86, NS86, SIN87, SN87, SST88, Tar82a, Wit88]. LR-attributed [NS86, SIN87]. LR-parsers [SN87]. LRU [WFL83]. lT0L [WU88]. LTUR [Min88].

M [PG87, Sar80, TTv86]. Machine [GT83, AG84, Hou87, YYS83, SB89]. Machines [CC84, Hro85a, Kir88, RSD81, TS81, YG83, CIRB87, CIR88, Chr84, Die89, FG89, Gor80, ITT82, Par87, Pas80, Suz88, Van89b]. made [Kal86]. Maggiolo [Boo81b]. Maggiolo-Schettini [Boo81b]. Magic [CGTW89]. main [Mor87a]. majority [Sav88]. makes [Boo88]. Mäkinen [Gro91]. management [BDS82]. Manhattan [Lip84]. manipulating [Ryt81a]. Manipulation [SO87, Sak88]. many [Sch81a, SY84a]. map [Wil80b]. mapping [Li88a]. mappings [Spy82]. Maps [Sha82b, Wil80a]. March [Tar86, TV87]. marking [Der80, Lyo88]. Marriage [KK85, GI89, PG87]. Marriages [All83, Hul84]. Master [Bv82, DS81, DMS81]. Match [AMM84, CC88a]. Matching [GK86b, IMM81, IS86, KC87, LY86, Mey85, Tak86, AG84, BP83, CR87, CD89, DF84, GL89, HK81b, IM81, II86, ISHY88, LK89, PS89, SV82, SV87]. matchings [Ans87, Sav80]. Math [Žák84]. mathematics [Bli88]. Matijasevitch [BG89]. Matrices [Bit85, Cha86, Mak86a, MA85, PV80, Wis85, BES82, Col79, CM80b, CR89, IMR80, ZG89]. Matrix [BP85c, CRL85, JT85, Kir88, RH88, San80, Sch82b, SM83, AS88, Bou85, Coh83, CM80b, DK88, GP89, IM81, Lot84, Mak81, Mor81, Pan88, Pet80b, Rom80, Vys84]. matrix*vector [ZM87]. Mauceri [Kri82]. Maurer [Žák84]. max [NI89]. max-flow [NI89]. maxima [Buc89]. Maximal [FP82, IS86, RSR80, AM87, Avi80, Deh86a, Dev80, HL88, II86, Iwa89, JYP88, JM88, Mar86a, Miy88, van80b]. Maximal-Cut [FP82]. Maximum [Has81, KC87, Sav80, AH87, DJ89, HS89, IM81, JS87, Sch86a, SH89, Shi89, SV82, WW85, YG87]. May [VT90, EG89]. maze [Sze82]. mean [BF87a, BF87b, Win82]. measure [Sky83]. measures [BY87, CW82, Lis81, Oza83, Van89b]. mechanism [PP80, YYS83]. mechanisms [SR88a]. median [CY85]. Mehlhorn [Flo91, Ryt86a]. membership [LD88]. memories [CLW80b]. Memory [RSD81, Vit81, BJ80, BDS82, CLW80a, Kfo85, Kuč82, Kuč83]. Merge [NP85, van80a, Erk80]. Merging [DD81, EY86, HR89, BES81, Car86, MU84, Zhe83]. mergings [AC87]. Mesh [HS85, Deh86a, DS88, GS89, Sav88]. Mesh-Connected [HS85, Deh86a, GS89]. mesh-of-processors [DS88]. Message [RDM88, BS87, Ber87a, KT88, LMT87]. message-efficient [LMT87]. messages [Har81a, Lav89]. MetaSoft [Bli88]. Method [Der85, AG81, Bui77, Bui79, CP82, Day80a, Far82, Gol87, NM87, NA89, Pet80b, Wil80a, Wit88, YH87]. Methods [JB84, Day80b, Gun89, Ove81, Ov81a]. metric [GS83, MC88, MC89a]. Microcomputer [DRT81]. Microprogramming [LR80]. Microprograms [LR84, RM85]. Mikhaılovich [Zem82]. MIMD [MRS84, McK85]. min [NI89]. min-cut [NI89]. Minimal [CC88b, CH86, LS82a, LR86, NM82, RSR80, Akm87, DM89, Dub88, EM84, GP88, LP82, PY81b]. minimax [RW87]. Minimization [Jan89, LL84, WFL83, BM89b]. minimize [Ber80, BWD84]. Minimizing [BF87a, BF87b, FS84, BC87]. Minimum [CW84, CLW80a, DD81, IMM81, MP88, Tsi88, AC89, BP83, Dil87b, DF84, GT88, GI89, HL88, JK85, JM88, PR86, Sch81b, Sho86, Tar82c, Tar86, Ver88b, Ver89b, ZLS87]. Minimum-cost [MP88]. Minimum-Weight [IMM81, BP83]. miracles [Mor88b]. mismatches [GL89]. Misra [Ngu85]. missing [Sik81]. Mixed [Zan80]. modal [Auf88, Far82]. Model [Bis84, WFL83, ACK89, CM80a, Che82, Fal89, GP88, KPK89, NA89, PPK88, Ram87, RBP82b, San89a, SN89, TST83]. Modeling [DLF87]. Models [Jaj85, May83, Lot84]. Modern [And80a]. modification [Cla83, KK89b, Van89b]. modified [SSA89, Wil81a]. modular [BSM^{+}88, Dai86]. Module [AR80a, CCL89a, SSA89]. Modulo [BP85c, Cha85]. moments [Dob82, Pan86b]. monitor [KK89b]. monoid [BL87]. Monoids [Dub85, Bea89, CD86]. monotone [Dun90, Gol80, Koc89, LN88, Tie84, Weg82a, Weg83]. monotonicity [HY82, PS81b]. Montague [CK86a]. Monte [Ber82]. Morphic [KW84]. morphism [Ber86b, Car83]. morphisms [Pan81b, WS82]. Morris [Bar81, AJV86]. Morse [dV88, Pan81b]. motion [Pap85, Sha89]. movement [TV81]. moves [Vit80, Wal82]. moving [SY84a, Wal82]. MR [AG88b, APT82, Bal87, Bes84a, Bin82, Bv82, CIR88, CM80b, Flo91, Fou79, Gro91, Kri82, Kuč83, Llo80a, Mor81, NI80, Ov80, PG87, Sar80, SY93, TTv86, Ver89a, Ver89b, Wyr84, YWL85, Žák84, vW80a, Day80a]. MST [Ola89]. Much [Sch83b]. Multi [GS81, Alb89, Fle89, KO88, LD88, CL81]. multi-dimensional [KO88]. multi-parent [LD88]. multi-pass [Alb89]. Multi-problem-solving [GS81]. multi-processor [Fle89]. MULTI-RESET [CL81]. multicache [Lop89]. multicommodity [NI89]. Multicounter [Hro85a]. Multidimensional [Har82, Ore82]. multidisk [CC88a]. Multihead [Ryt84, SV84, CR87, Ryt81c]. Multilevel [Lei82]. multiparty [Fra89]. Multiple [Hoc87, LM81, Tho85, CGPD89, CGPdS89, CC87, DF88, Hel81, LM80b, Mei81, Ray89a, Zem84]. Multiplication [JT85, San80, Sch82b, AS88, Coh83, IM81, Lot84, Mak81, Mor81, Pet80b, Rom80, RH88, Vys84, ZM87]. Multiplications [Mak86a]. Multiply [Cha86]. Multiplying [Mak86a, PV80]. Multiprocessor [BB83, Bud87, BS87, KS87b, ÖA88]. Multiprogramming [Bro83]. multipushdown [Chr84]. multireader [Vid89]. multisection [BES82]. Multiserialization [FR84]. Multiset [JL82, Mar87]. multistage [BFJM87, Var89b]. multitape [Fre81, Kin80]. Multivalued [Kat84, TC84, GT84, Vid89, Weg82a, Zan80]. multivariate [Val80]. mutable [GS88b]. mutational [Kra80]. Mutual [Hoo86, KP84, Mag85, Mar86b, AJV86, DT80, FP87, MB85, Mor79, Pet81, Ray87, SM92]. Myths [Pet81].

N [Bin82, Sar80, TTv86]. Natural [JB84, Bal87, Dob85, Wil87]. NC [Ede87, LK89]. Nearest [EZ82, MS86, Eas81, GS83, PS88]. Nearly [Kor81, Gil88, Kor82]. negation [Sky83]. negative [Cou84, Yap83]. Neighbor [EZ82, Cha83b, Eas81]. Neighbors [MS86, AK89, GS83]. neighbourhood [KNT87]. nested [AGvG89, DM89, Gil88]. Nesting [Jȩd87]. net [Sch86b]. Nets [Mag84, Mag85, Yam81, Mag87, Pet80a]. Network [IM88, Jaj85, Les83, LMW86, LMW88, Van85b, BP85b, Bir85, CM89, GMZ89, GT89, Gol87, Kři86b, LPP88, SR81, Tie84]. Networks [Bro87, PV80, BFJM87, CK88b, Eic83, Gos89, GS84b, Has81, JK85, MW87, NI89, PP89, Var89b, YH87]. Newton [NA89]. next [Ram89b]. NLOG [Sho89b]. NLOG-complete [Sho89b]. NLogSpace [LS82b]. no [And80b, AG88b, APT82, Bal87, Bes84a, Bin82, Bv82, CIR88, CM80b, Day80a, Flo91, Fou79, GGPS89, GPB83, Gro91, Kri82, Kuč83, Llo80a, Mor81, NI80, Ov80, PG87, Sar80, SY93, Sze87, TTv86, Ver89a, Ver89b, Wyr84, YWL85, Žák84, vW80a]. node [Gos89, GP86, Iwa89, IRV88, KR84, Sch89]. node-disjoint [Iwa89]. node-rewriting [KR84]. nodes [CP81, Yap83]. Noisy [GT83]. Non [Avi84, Kat84, ZBS83, CP82, Dil87a, YD86]. non-additive [CP82]. Non-Conflict-Free [Kat84]. non-Hamiltonian [Dil87a]. non-Horn [YD86]. Non-Partitionable [Avi84]. Non-Uniform [ZBS83]. nonatomic [AG88a]. Noncommutative [Mak86a]. Nonconstructive [FL87]. nondegenerate [Dil87a]. Nondeterministic [Hol86, Coo88, MC80, Ryt81b, Ryt82a, Ryt89, Sis89, Upf82]. Nonisomorphic [Hat86]. nonlinearly [BU89]. nonoptimality [Lev87]. nonpolygonal [Lau87]. nonpreemptible [Klu84]. nonprimitive [Cal87]. nonredundant [Got87a]. nonrigidity [GK86a]. nonsequential [Eic83]. Nonserial [FB88]. nontraditional [BG89]. nonuniform [SSW87]. normal [ABC87, Bis89, KM80, YWL82, YWL85]. Normalized [KO88]. north [GS83]. north-east [GS83]. notation [Ber87b]. Note [AS86c, CK86b, FS86, GS87, KR84, LS86a, Mäk86b, Pit87, Ros86, SSS86, Yao80, Baa88, BD82, Blu83, BT89, Bv82, Bui80, Bur89, CCS89, Cha85, Chr84, CH88c, CL86a, Coh83, CW82, Dev80, Dev85, Dob85, DJ89, Ebe80, Eis89, Fie87, Flo91, Hol84, IMR80, IT80, ITT82, JK85, Jan85, Kem87, Kir80, Kob87, KV81, Kre87, Lei80, LM80a, LP89b, Lic84, LP83, Lou80, Mäk85a, MT87a, Mei81, Mor81, Mot82, Mou80, Nar89, Nij82a, OP85, Pan86b, Pan81c, Pet80a, Rob89a, RG89b, Rot89, Ryt82a, Ryt89, Sar80, SM92, Sha89, Slu85, Stå89, SY93, Sto86, Szw87, Ten87a, Var84, Var89a, Vys83, Vys84, Wri87, Yeh85, Yod84, Van87b]. Notes [dI89a, Sze89a]. Novel [Hug86]. November [Sak88]. NP [Ber81, Bod89, BHZ87, Dam89, FS88b, FPT81, Ko82, Kři86b, Rob80, SY84a, SV82, Tou84, Weg85]. NP-complete [Ber81, Bod89, Dam89, FPT81]. NP-completeness [Kři86b, SV82]. NP-hard [Ko82, Rob80, Weg85]. NP-hardness [SY84a]. NTS [Boo81a, Fro81]. Null [AM84, Kel86]. Number [Ber84a, Hro85a, SS85b, AC87, Bac82, Ber88a, Bod89, Buc89, Car87b, CP81, Che82, CLW80a, GP88, Kem87, Lav89, Pep87, RR89a, Sla82, Tou84, Zer89]. Number-Theoretic [SS85b, Tou84]. numbering [GM88]. Numbers [Er83c, Mül87, CH89, GJ82, GL80, GP87b, Kuč89, Mio84, Pet80b, Ray89b, Tsi89, WS80b]. numerical [Ber82, Bin82, Tsa81].

O [Dwe88, SM87]. Obituary [Zem82]. object [Bis89, Fei89, van89a]. object-oriented [van89a]. objects [EM81, Kel89, Rei88, SN89]. Oblivious [Kol83, Vit85b]. observation [Mal87]. observations [CIRB87, CIR88, Ko82]. Obstacle [Lau87]. obstacles [Fei89, Roh86]. Obtaining [Ram87]. Occam [BEHW87]. Occurring [LM81]. October [MT87b, TM89]. Odd [Der85, Ebe80]. Odd/Even [Der85]. Off [Pan86a, Slu87, Die89]. Off-Line [Slu87, Die89]. Office [EM85]. On-Line [Tak86, IK83, Mat89a, Mel87, ITT82, Lia80, NI80, RW87]. on-the-fly [RM83, Van87a]. Once [Tru80]. One [Hås87, LS86a, LY86, Mon84, Nak81, Ryt84, UMS82, Vit85b, Wal84, Aur86, CR87, Die89, Fre80, IW81, Jan85, Kem89, Li88a, Ryt81c, Sav88, Tie84, Var89a, Vit85a, Wat88]. one- [Fre80]. one-dimensional [Aur86, Li88a]. One-Head [Mon84, Vit85b]. one-output [Tie84]. One-Pebble [Nak81]. one-rule [Jan85]. one-tape [Die89]. One-To-One [LS86a, Kem89]. One-Way [LY86, Ryt84, UMS82, Hås87, CR87, IW81, Ryt81c, Sav88, Var89a, Wat88]. Ones [Sch83b]. only [NO88]. opaque [Akm87, Dub88]. open [Rao88]. Operating [MF84, Fre83b]. operation [Ida83, San89a]. Operations [Clo83, Ber87b, HM81, MB85]. Operator [Jȩd88, HL87a, Wil81b]. operator-precedence [HL87a]. operators [EPS82, GP88]. opposition [OR89]. Optimal [Alb89, CC88a, Eas81, Eic83, FPT81, GH87, HR89, HK81a, Hig86, IJ89, IRV88, Kor81, Kut88, Lei82, Leu82, MO81, PV80, RR89b, Sch89, Sil81, VRL87, Ver88b, Ver89b, Vit85b, Weg85, Akl84, AGS89, AC89, BDS82, Car87b, Cha83a, Cro81, CLW83, Gil88, GJ82, KST82, Kir80, Kor82, LMT87, Lic86, Mak81, Ov81a, Ryt89, SS89, SN88, Sha89, Sle80, Tsi85, Vog89, WW85, WW87, Wu89, ten80]. Optimality [Mag83, AT81, BP85b, GS84b]. optimally [Col88]. Optimization [FH87, DI89b, NU88, SS81]. optimizations [Fra83]. optimizing [Fal89]. Optimum [CN88, Shi83, Wit88]. oracles [Bab87a, HH88]. Order [CR88, Er83c, Sza87, vR88, Bur89, FH87, GT86, GMZ89, Gil88, IMR80, Pet80b, Ram87, Sem81, Sho89b, WS80b]. Order- [Er83c, Pet80b, WS80b]. Ordered [CC88b, KR85, NI83, Shi83, CS83a, Gro91, Kem89, Mäk89]. ordering [BGLY81, Fra83]. Orderings [JL82, Mar87]. orders [Fus83]. ordinals [Deh86b]. organization [KV81]. Organizations [PM84, CMT88]. organizing [CSO89]. Orientation [Ata84, Vis85, Tsi85]. Oriented [Güt83, dI89a, van89a]. orthogonal [EM81]. outer [Dev85]. outerplanar [EMM89]. output [Man82, Sys82, Tie84]. Overflow [Tam82]. Overlapping [Win81]. Overloading [WS80a].

P [AG88b, Kot89, TTv86]. paced [Man82]. Packing [KLV87, FV89a, FPT81, Fre80, Leu89, Lia80, Shi89, Sle80]. page [Wee88]. Paging [WFL83, HL87b]. pair [Dye80, HNS88, Mos82]. Pairs [Lau85, MT84, Ove81]. Papadimitriou [Sha89]. Papadimitriou-Silverberg [Sha89]. paper [Bin82, CM80b, Dun90]. paradox [Kel89]. Parallel [Ata84, BES82, Ber84a, Ber88b, BP85c, Chl89, Cza80a, Cza80b, Fri87, GP89, GIR88, GP87a, GLPG82, HS85, IS86, JM88, KNI86, KC87, Kuč82, Kuč83, LS86a, LP89a, MRS84, Mor88d, NM82, RS89, Sch83b, SB89, SV86, Tir83, Vis85, Wei83, Wis85, Akl84, AGS89, AC89, BES81, BP87, Blu83, Bop89, BM89b, BMRS88, CM89, Chl88, CY85, CL88b, Dai86, Deh86a, Dye80, EY86, GM88, GT89, Goo87, Hag88, HL88, HL89, Hul84, IMR80, II86, JK85, Kaw80, Lee89, LP89b, Mar86a, MG89, McK85, NA81, Nao87, OT89, PR86, PGR87b, PGR87a, PK88, RGG89, RR89b, Rub87, Ryt88, Ryt89, SS89, SV88, SR88b, Shi89, Sri87, Tsi85, Val80, WT87, Zhe83]. Parallelism [AM87, BDP88]. parallelizable [Miy88]. parallelized [Hag89]. Parameter [CC88b]. Parametric [GI89, Sal89]. parent [LD88]. parenthesis [Sem81]. Parikh [YJ83]. Parikh-Boundedness [YJ83]. Parse [Kri86a]. parser [Kru88]. parsers [MF81, SN87, SST88, Ukk85, Wit88, Wyr83, Wyr84]. Parsing [CK86a, Tar82a, Bar83, BCC83, HL87a, SB89]. part [SN87, Nja85]. Partial [AMM84, CD89, BGLY81, CC88a, McC89]. Partially [Dub85, EZ82, BL87, CD86]. partition [LP83]. Partitionable [Avi84]. Partitioned [GY87, HL87b]. partitioning [AN80, Arb88, Cha83a, DF84, GY87]. partitions [Kři88b]. parts [LN88]. pass [Alb89]. passing [BS87, Ber87a]. past [Fie87]. Path [Gun89, IN84, AM87, AAI87, BYV82, Ber81, Fei89, GST89, Lip84, MP88, MT84, Mor81, Nar89, Ó’D88, Pap85, Rom80, Rom88, Ryt82b, Tar82c, Tar86, Var89b, Yap83]. path-finding [MP88]. path-planning [Ó’D88]. Paths [Der85, EMM89, Ita88, Iwa89, Roh86, Roh88, Ryt88, Wat81, dI89a]. Pattern [AG84, Tak86, CD89, ISHY88]. patterns [CSO89, Par86]. Patterson [FZ82]. Paw [Ola88b]. Paw-Free [Ola88b]. PDL [PT85]. Pebble [Nak81, LT80, Lou80, Sze82]. Perception [Gei84]. Perception-based [Gei84]. Perfect [CC88b, CS83b, GR82, IMM81, HL89, LK89, Ste89]. Performance [HL87b, Mag84, Mag85, Mag87, Apo86, Fra86, Gun89, WI88]. perimeter [DJ89]. Periodic [BB83, Nja85, CC89, CCL89b, LM80a, Ver88b, Ver89b]. Periodically [LM81]. permanent [Mih89]. Permutation [Har82, Ros86, Sat86, Sup85, Bir85, CLW80a, CLW80b, FVG87, Hua81]. Permutations [Atk85, Mai82, McK84, BK88, Hås87]. permuting [TST83]. persistence [Gra80]. Persistency [Yam81]. persistent [Mag87]. Petri [Mag84, Mag85, Mag87, Pet80a, Sch86b, Yam81]. phase [BO86]. Physical [Sch83b]. piece [Leu89]. piecewise [Sha89]. piecewise-linear [Sha89]. pipelined [DK88, ISHY88]. pitfall [Kel89]. pitfalls [SN89]. placing [CK88a]. Planar [Fre84, Hig86, Lau85, Gol80, Gra72, Has81, JK85, KNT87, LP89a, MP87, Nao87, PY81a, Sha88a, Ste89, Tur89, Wil80a, RH87]. Plane [CC83, EM85, HNS88, IMM81, BP83, CH86, FPT81, Ima82, Rei88, RW87, Roh86, SSW87, WT87]. Plane-sweep [HNS88]. planning [Ó’D88, Sha89]. Plotkin [Iwa89]. Point [Avi84, BL89a, Hig86, AH89, AC88, Goo87, LNS82, LP89a, PS88, RW87]. Pointers [And80a, SO87, Sak88]. Points [EM85, Sup85, BR89, Byk78, CK88a, DS79, EK89, Fou79, GZW85, MC88, MC89a, Ov80, WT87, WM89]. Poisson [Pob86]. Policies [Mag83, MT87a]. policy [HK81a]. polygon [AK89, Akm87, Ber86a, BM89a, DM89, Dub88, Hig82, PS81b]. Polygonal [EMK82, Roh86]. Polygons [Güt83, AAI87, Ata83, CW84, Har89, Li88b, LN88, Ola88a, Ola89, Roh88, RD87, Sch81b, Woo84]. polyhedra [Kla87]. polyhedral [Sha88b]. polyhedron [Cha83a]. polyline [Mel87]. Polynomial [BP85c, Kri86a, McK84, Rys83, Son85, AS86b, Ans87, Arb88, Bab87a, BP87, BCW80, BT89, FL87, Hat86, Huỳ87, Lau83, Pan88, Rol89, Sab80, Sab81, SY88, Sli82, Zip89]. polynomial-size [Huỳ87]. polynomial-time [AS86b, Bab87a, BT89, FL87, Sab80, Sab81, Sli82]. Polynomially [GS88a]. Polynomials [BM81, CL80, Sch81a, Val80]. Positive [BB88, Weg85]. possibility [Cia86]. Possible [Wil83a, Weg82a]. Post [EM85, Pan81c]. Post-Office [EM85]. postponement [Wit88]. Powell [OR89]. Power [BP85c, CK86b, SV84, LC85, Sch82a, Vys83]. Powerful [Mon84, Sky83]. Powers [SM83]. pp [AJV86, Avi80, Dun90, Hel81, LMW88, Mei81, MT87b, Pet91, Sak88, Tar86, TV87, TM89, VT90]. Practical [IS87, Asa83, AS88, Leu82, SIN87, Wit88]. PRAM [HR89]. PRAMs [LPP88]. Pratt [Bar81]. precedence [HL87a, Wil81b, Wyr83, Wyr84]. Precise [Mar83]. preconditions [Mor87b]. Predicate [Ber87a, OR86, MC80]. Predicative [Bro87]. Preemptive [BB83, CL88a, LM80a, MT87a, RS87]. Prefix [EH88a, Kri82, MR81, Wyr83, Wyr84]. Prefixes [AFG87]. Preperfectness [NO88]. Presburger [Jaf81, Ros86]. presence [BO86, Tau89, Yap83]. presentation [MR84]. presented [MO87a]. Preservation [Mäk84, Tir83]. Presorted [Her83].

Prespecification [HHS87, Orł88, HH87]. prevent [Ray87]. Prime [Ray89b]. primitive [FR89, ZLS87]. Primitives [KF85, VU86, Mar85a, Mar85b]. principal [Mor88c]. principles [Ov81b]. priorities [Gos89]. Priority [Car87a, MT87a, MT84, Sun89]. privacy [Wil83d]. privileges [LJ86]. Probabilistic [CC84, FV89a, Fre80, IMR81, Bab87b, BDD^{+}82, CL80, Fra86, Fre81, IM81, Ko82, Lei80, LPP88, Man84, SS84a]. probabilistically [BCW80]. Probabilities [ZBS83, AS86a]. Probability [SM86, Sha82a, SY88]. Probable [Ham87]. probe [Mar85a, Mar85b]. probes [Ber86a]. probing [Ram89a]. Problem [AR80a, AB87b, CK86b, EM85, FS86, Far80, HV87, Jos85, KK85, KP84, LL84, Llo83, Ryt86b, Sav82, TS86, AN84, AM87, AJV86, Apo87, AC89, Ban80, BH89, BP89, BS89, Ber81, Bir81, Boo81b, BDP88, BL80a, CP82, Cha88, Cha83b, CCL89a, Cul89, Dam89, DM89, Deh86a, DS88, Dye80, ER80, Flo91, FV89a, Fra86, FP87, FL84, Gol80, Gro91, GS81, HNS88, HS89, Hua88b, Jan89, Jar81, JS87, Kes80, Kin80, Kla87, Kre87, Lin89, MO87a, MP88, Mäk85b, Mäk89, Maz88, Meh88, Mei85, Mir87, Miy88, MT84, Mor81, Mor79, Mot82, Nar89, Pan81c, PY81a, Pet81, PG87, RV87, RDM88, Ran83, RR89a, Rei87a, RT85b, Rom80, Shi81, Shi89, Sho89b, Slu85, SV82, Var81, WW89b, Wil84, Wil81a, Wil83d, Woo84, YG87]. problem [Zem84]. Problems [AS86c, AB81, AK85, BS86a, BM81, BO85, Cza80a, FS84, FP82, IMR81, IN84, Ili85, JT85, Kel86, Kri86a, MSR85, MRS84, NM82, Oya84, UMS82, Arb88, BG89, BL87, Bru81, DJ89, EO82, GS88a, GS89, Gir86, HK81b, Huỳ87, IT80, KO88, Ko82, Kři86b, MO81, Ove81, Ov81b, Ov81a, Pan81a, RR88b, RH87, Rys83, Sha88b, Sli82, SS81, WK88, Zem84, vW80b]. procedural [Lau80]. Procedure [AS81, Cha83a, RG89a]. Procedures [GPB82, Cia86, GPB83, LBM88]. Process [And80b, AJV86, AG88b, APT82, Avi80, Bal87, Bes84a, Bin82, Bv82, CIR88, CM80b, Day80a, Dun90, Flo91, Fou79, GPB83, Gro91, Hel81, Kri82, Kuč83, Llo80a, LMW88, Mei81, MT87b, Mor81, Ov80, Pet91, PG87, Sak88, SY93, TTv86, Tar86, TV87, TM89, VT90, Ver89a, Ver89b, Wyr84, YWL85, vW80a, AR80b, BMRS88, FR89, PPK88, Sza88]. Processes [Bro83, Les83, Mag85, May83, RT85a, Cza89, Dai86, Gos89, Hol84]. processing [BT87, GS84b, Hem80, MPY87, PK88]. processor [BWD84, Fle89, Jan89, Kub87, Lee89, LFQ^{+}81, McK85, SSA89, Sha88a, Tsi85]. Processors [Ber84a, Jin84, LM81, Van85b, Wis85, BF87a, BF87b, Bod88a, CYC89, DF88, DS88, Kub87, Lav89, Pep87, RTV89]. PROCOL [van89a]. produce [Gru84]. produced [Mét83]. Product [MSR85]. productions [AN86]. Program [AK85, BH87, BW80, Cha81, WFL83, AS86a, Dud83, Fal89, Hem80, Kor87a, KL88, Sis89, TC83, Upf82]. Programmed [CK84]. programmers [CLM89]. Programming [AS81, Cla86, GH81, LM85, LR84, Dai86, FB88, Heh89, LBM88, LT88b, McC89, Ten87a]. Programs [Bro80b, Bro87, Con83, Hol86, IMR81, Las86, May83, Mil83, Tra84, Abi89, AG88a, ABC87, Bro86, Bur82b, Dwe88, Eic83, Fra89, Gin82, GS88b, Kaw80, Kfo85, Lif84, MC80, Pas80, Ryt81a]. Projections [Fre81, Wei83, GG82, Li88b]. Prolog [CW86, Fal89, KM88, MTK89, Sto86]. Proof [DK84, Orł88, BA80, Cia86, Mis89, Ngu85, RM83, VT90, Ver87, WSA84, Yoc89]. proofs [BHZ87, Con83, Fra89, Upf82]. propagation [LJ86]. Proper [Ber83]. Properties [Dub85, Tir83, Tra84, Lis81, Muk89, NI89, NI80, Rao88, Suz88, Tur84]. property [Cho81, EM84, HY82, Kfo85, LD80, MC89a, Mor88c, RM83, Rub87, Taz84, WSA84, Wyr83, Wyr84, de 81]. Propositional [AB87b, HS88b, Auf88, Coo88, FI87, YD86]. Protection [AB81, Bis84, Bud85, San89a, San89b]. protocol [Lop89, van89a]. protocol-constrained [van89a]. Protocols [Win81, Bab87b, FS88b, KT88]. prove [KOR87b]. Proving [Rao88, Suz88, Kes88, Kři86b]. pseudoforest [GT88]. public [SSA87]. pumping [Yu89]. Pure [Mäk86b, BP88, Stå89, HM81]. Purging [SM80]. purpose [AFK88, McK85]. Pushdown [IR81, Mon84, Ryt83b, Vit85b, Ryt81b, Ryt81c, Ryt82a, Ryt83a, Vit85a, Yeh85]. Putnam [GPB83, GPB82]. puzzle [SS84a]. puzzles [VD87]. Pyramid [GT83, Sto82].

quadratic [IT89, PS80, PS81a, Sni81]. quadratically [BU89, dV89]. Quadtrees [SM86, Wis85]. Quantification [CK86a, Ten87b]. quantified [APT79, APT82]. quantities [Pan81a]. quantization [Wu89]. quaternary [OIM84]. Queen [FS86, Rei87a]. queens [Roh83]. Queries [Wil83a, DI89b, Fie87, FMNT87, GS84b, LW82, Taz84]. Query [HV87, Lei86, CM80a, EW86, Sik81]. Queue [Clo83, Van85a, Vit85b, HM81, MT87a, MT84, Sun89, Vit85a]. Queueing [Les83]. Queues [Car87a, Bur82a]. Quickly [Wal84, CH89]. quicksort [MG89, Röh82]. quiescence [Mat89c]. quindiagonal [BES82]. quotients [BSZ89].

R [AG88b, TTv86]. Rabin

[CL80, Mos82, SS84b]. radius [Cha83b]. RAM [Man82]. Rana [AG88b, TTv86]. Random [Bab87a, Sat86, BK88, CH88c, Ev89, Fri87, Fri88, GZW85, Mye83, RGG89, Ram89a, Sha82a, Tsi89, Zip89]. random-access [Mye83]. randomised [Zer89]. randomized [II86]. Randomly [TS86]. Range [Wil83a, EW86, EO87, FMNT87, LW82, Mat88]. Rank [IMR81, IMR80, Lic84]. ranking [IRV88, Maz88, Sch89, Zak82]. ratio [KS87b]. Rational [CK86b, DK84, EH88a, Mio84, Rom88, vN82]. rationality [Sch86b]. Raymond [SM92]. Rayward [WI88]. Rayward-Smith [WI88]. Razor [BEHW87]. readers [Kes80]. Real [Clo83, Heh89, HM81, Jos85, SV84, Son85, UMS82, CC89, LO89, LM80a, LL88]. Real-Time [Clo83, Jos85, SV84, UMS82, Heh89, HM81, CC89, LO89, LM80a, LL88]. realization [MP87, SN87]. reasoning [Pet91]. Recognition [GR82, Ryt86a, Ede87, Gab81, vN82]. recognizable [Fre81]. recognize [Sys82]. Recognized [Sao84, Ryt81b, Ryt81c]. Recognizing [Sav88, JF82, SV88]. reconfiguration [ÖA88]. reconnection [Bra88]. Reconstructing [Ski89, Wei83]. Reconstruction [Li88b, SPL^{+}89]. Record [ZBS83]. recording [VRL87]. records [Che82]. Recovery [Llo83, CK88a, HK81a, Mat89c, VRL87]. rectangle [EO82]. rectangles [WW87]. Rectangular [Nak81, Lot84]. Rectilinear [GK86b, HK88, Woo84]. Recurrence [GLPG82]. recurrences [GPS87]. recurrent [Par86]. recursion [CLM89, Wat80, VV87]. Recursive [Bar83, Tra84, Yoc89, Abi89, Cal87, Kru88, Par86, Rob89a, Ryt81a, Sab81, Sha82a]. Recursively [CC84, RS89]. reduce [CYC89]. reducible [LK89]. Reducing [SM86, NO88, SM87, TV81]. Reduction [BP85c, Tha85, Tir83, Var89a]. reductions [Van88]. redundancy [ten80]. redundant [Cha85]. reevaluation [TW89]. Reference [CK84, KNI86, TJP87]. references [NM87]. refinement [HHS87, Jos88, Mor88a, Mor88b]. reflected [San89b]. Reflexive [Bud85]. regions [SS81]. Register [LR84, Vid88, Vid89]. regular [BA80, Hea89, Jȩd83, OSW81, Rot89, Ryt89, SN87, Vid88]. regularity [GP88]. Related [UMS82, Sha88b, Zem84]. Relating [Weg83]. Relation [Cza80b, AGvG89, JF82, MC80, Mäk85a, MGKL88]. Relational [Jaj85, Kel86, MST84, Cas83, Che82, CLM89, Dem82, DI89b, GM82, Kan80, LD80, MC80, Sik81]. Relations [AM84, LS82a, LB89, Zan80, ZLS87, vN82]. relationship [Nij82b, Oli80]. relative [KNT87, MC89b]. relativised [Zip89]. relativization [KL87]. Release [Fre83a, SY93]. Remark [Apo87, NI80]. Remarks [CR87, Miy84, Nak81, Oya84, Sze88, Cia86, Sze89b]. removing [OP85]. renaming [LS89, MM85]. Renewal [Les83]. renumbering [Boo88]. reorganisation [Leu82]. repair [CC86]. repeated [EO87]. Repeating [HP85]. repetition [Hoc87]. repetitions [Cro81]. replacement [HRY87]. replication [Muk89]. Reply [Bin82, Mor81]. represent [AGvG89, Coo88, Ev89]. Representation [AT87, Hug86, BL89a, GS81, KOR87b, Per83]. representation-independent [KOR87b]. representations [LS82a]. Representing [Wis85]. required [Dev85]. Requirements [FP83, CS82, ten80]. Requires [Kir88]. requiring [Bar83]. rescaling [AMA88]. reset [Wag86, CL81]. residue [ABM84]. residuosity [IT89]. Resolution [YD86, Auf88, BA80, Cia86, Eis89, MZ82, Min88]. Resource [MF84, CP82, Jan89]. Resources [Llo80b, Win81, Ber80, Klu84, Llo80a, Llo81, vW80a]. Response [DEGS84]. Restivo [Kri82]. restricted [DL88, Leu89, YWL82, YWL85]. Result [JM86, Ban80, Ryt83a, Sza88]. Results [Har85, JV86, Vol85, Cou84, Kři86b, KOR87b, MW87, Mur83, Wan80]. Retrieval [AMM84, Bit85, Eva86, Taz84, Tru80, ten80]. reversal [Bir85, Chr84, GST89, Par87]. Reversals [Hro85a, Kir88]. Reverse [Hug86]. Revisited [Hol86, Der80, Hag89, Kel87, LC85, RR88a, Van87a, Wal82]. rewrite [Van88]. Rewriting [BL87, KR84, MO87a, Mét83, Pla85, Rus87, Toy87, Toy89]. right [SN87]. Rights [Bud85]. Ring [HS88a, TM89, RTV89, Suz88]. rings [Bod88a, Lav89, Wal82]. Risks [Lei86]. RNG [Ola89]. Robinson [CDS89]. Robust [RBP82a, BL80b]. rollback [Tsi89, VRL87]. ROM [IT88]. Romani [Mor81, Mor81]. root [Sys82]. Rosen [Boo81b]. Rosenbrock [Day80b]. Rosser [Boo81a]. rotation [LP89c, Mäk88b, Pal87]. Rotations [Tar83b]. routes [CN88]. routing [AT81, Bir85, SS81, Var89b]. Routings [IM88]. Row [Bit85]. Rows [MA85]. rudimentary [Cal87]. rule [Abi89, BP88, GPS87, Jan85, KM88, NO88, Tak87, Yoc89]. Rules [BS86a, BO85, WS80a, NO88]. running [Vit80]. Russians [San80]. Ruzzo [KL87].

S [AG88b, Kri82, PG87, TTv86]. Safe [Lau80]. Safeness [KK83]. Safety [ADS86]. salesman [AN84, BH89, Dil87b, Kan83]. Salomaa [Žák84]. same [Hat86]. sampling [CH88c, GZW85, RGG89]. Sardinas [FZ82]. Satisfiability [AB87b, FB88, Fra86, GS88a, JS87, YD86]. satisfies [Sis89]. satisfying [CL88b]. saves [Kie86]. saving [Hag88]. Savitch [Tom81]. scaling [KO88]. scanner [ISHY88]. SCCS [GP88]. schedule [BWD84]. Scheduling [BB83, BWD84, CC89, DL88, FS84, FP83, Fre83a, GS84a, LM81, Llo80b, SY93, Ber80, CM80a, CL88a, Jan89, Kub87, KS87b, LM80a, Llo80a, Llo81, RS87, Vog89, vW80a]. schema [WSA84]. Schemas [Cza80a, Cza80b]. schemata [Sab81]. Schematic [Tra84, San89a]. Scheme [AT87, CC88b, SS84b, Vit81, BJ80, Hua88a, IW81, LLH89, MTK89, TC83]. Schemes [AK85, Cha81, AT81, CH88a, HL89, JF82, OS88]. Schettini [Boo81b]. Schneider [Bv82]. Schorr [Ham86, Der80]. Sci [NI80, Sar80]. science [Heh81]. scores [Vit80]. Search [HC85, Kor81, Kri86a, LBB81, LL85, Man86, Reg81, Rei85, SS85a, Sav82, Sil81, Tar83b, Awe85, BY87, BG89, Bru81, BL80b, CGPD89, CGPdS89, Cid88, EW86, Gor86, GP87b, Kor82, LMT87, Mak88a, Mat88, Sal89, SV88, Sha88a, SIM89, Sze82, Ver88b, Ver89b, Weg85, WM89, YH87, ZC87, dI89a]. Searches [EZ82]. Searching [Güt85, MNS84, Ore82, BD80, Bar84, BGLY81, Deh86a, DS88, Eas81, EMK82, EY86, Lev82, MO81, Ov81b, Ov81a, YH87, vW80b]. section [Ray89a]. secure [Yun85]. Security [Lei86, McL85]. seed [Bou85]. Segmentation [Naw83]. Segments [Güt85, Lum85, Ros86, Wel85, Her89, Jaf81]. Selection [AN80, MA85, Akl84, Col88, CA86, Dob85, Erk80, FL84, MZ82, Mot82, Sal89, UY88, Wey84]. Selector [TB84]. Self [Pag85, Wog89, CGR87, CSO89, Eis89, Fal89, Kes88, Pla85]. Self-Adjusting [Pag85, Wog89]. self-embedding [Pla85]. self-optimizing [Fal89]. self-organizing [CSO89]. self-resolution [Eis89]. self-stabilization [CGR87, Kes88]. semantic [BSM^{+}88, Koc87]. Semantics [Bro80b, Bro83, MT86, MT87b, All89, AG88a, Bro86, EPS82, Gor80, LNS82, MC80]. Semaphore [KF85]. semaphores [FP87, Udd86]. Semi [MO87b, YJ83, CGTW89, Jan85]. Semi-Commutations [MO87b]. semi-joins [CGTW89]. Semi-Linearity [YJ83]. semi-Thue [Jan85]. semidynamic [AT81]. semigroup [CYC89]. Sense [LMW86, LMW88]. sensitive [CC87, Ebe81]. Sensitivity [Tar82c, Tar86]. Sentences [AB87b, Gru84, YYDH84, YD86]. separable [CW84]. separate [Bab87a]. Separating [Rob89b, HH88]. Separation [AK85, KL87, Zip89]. Separator [RH87, Li88a]. September [Pet91].

Sequence [LL85, Fus83, Pan81b, Wog89]. Sequences [Her83, Kal85, RSR80, Har89, KPS88]. Sequential [Lei82, Mag85, Man86, Rei85, Sch83b, Wei83, Bro86, Gor80]. serial [GS84b]. serializability [Kel87]. Series [CK86b, BI88]. Set [Car84, Cza80a, Güt85, Hig86, Kel86, Lei86, Sch84, Sup85, Ban80, Ber80, BR89, BDP88, Buc89, Byk78, CDL87, DS79, EK89, Fou79, Goo87, Gra72, Heh81, LS89, MM85, Mar86a, Mei85, Ov80, RDM88, RW87, Ski89, WT87, WW87, ZLS87]. Set-Theoretic [Cza80a, Kel86]. set-up [Ber80]. Sets [Avi84, Kat84, NI83, Tak88, Win81, AS86b, BYV82, BL89b, Ber84b, BT89, CLW83, Har83, JYP88, JM88, Llo89, Man84, McC88, Per83, Whi81, dPV88]. Several [Oya84]. Shaker [WS88]. Shaker-sort [WS88]. Shape [Ric83, Ber86a, SS81]. Shared [Vit81, IW81, SN89]. Shared-Memory [Vit81]. Sharing [AB81, Bv82, DS81, DMS81, San89b]. Shellsort [IS87]. Shifting [HM88]. shop [Jan89]. Short [Coo88, BHZ87, SN88, Nja85]. short-circuit [SN88]. Shortest [AAI87, Der85, Mor81, Pro89, Roh86, Rom80, Rom88, Fei89, Lin89, MT84, Pap85, Roh88, Sha88b, Tar82c, Tar86, Wat81, Yap83]. Shortest-path [Mor81, Rom80, Pap85]. Should [Hof83, Wal84]. Shuffle [Jȩd88, Bir85, Jȩd83, Jȩd87, LS86b, SR81]. shuffle-exchange [Bir85, LS86b, SR81]. side [Lam88]. sided [Ber86a, FMNT87]. Signals [Nja85]. Signature [SS84b, Lei80]. SIL [And80a]. Silverberg [Sha89]. similarity [BM89a, CW82, Oza83, ZG89]. similarly [GL80]. Simon [KL87]. Simple [BL89b, Fro81, GL89, HF86, Llo83, Alb89, AAI87, BM82, CM80a, CL88b, Far82, Fre80, GS84b, Heh81, II86, Kla87, Leu84, MC80, MU84, Mel87, Mis89, Ola89, PS81b, SN87]. Simplified [GPB82, BA80, GPB83, Min88, Rei87a]. simplifier [Sch88]. simplifying [Sni81]. simulate [Vit85a]. simulated [FS88a]. Simulating [TW89]. Simulation [Tir83, UMS82, Vit85b, IJ89, LPP88, Par87, Ryt81a, Ryt83a, Tsi89]. simultaneity [GY87]. simultaneous [Cou81]. Single [Cha81, CC88b, Abi89, NO88, Wag86, Wri87]. single-reset [Wag86]. sink [KST82]. sink-finding [KST82]. sinking [Mos82]. Situ [Dv82, FVG87, Han86, MU84]. Size [Hik83, WB86, Car83, EM84, Eas81, FL84, GG82, Got87a, Huỳ87, SSW87, SE87, Ukk85, Weg83, Van87b]. sized [NU88]. Sizes [Bou84, BS86b, Leu89, Ski89, Wil80a]. SJRPs [GS84b]. skewed [CSO89]. Skolemizing [McC88]. slicing [KL88]. Small [Ber84a, BDS82, CS82, CIRB87, CIR88, Kuč89, Mor87a, Rob89b, Tom81]. smart [Ram89b]. Smith [WI88]. Smoothsort [Her83]. snapshots [Hua89, LY87]. Software [Lop89, DT80, Kra84, Yoo85]. Software-controlled [Lop89]. Sokołowski [Nij82a]. Solution [KP84, AJV86, Bac82, CP82, DT80, EY87, FP87, Kel89, Mei85, Mor79, PS80, PS81a, Ran83, Rei87a, Sri87, Urb80, VD87]. Solvability [Maz88]. solvable [GS88a, GS89]. solve [AC88]. Solver [CL86c]. Solvers [CRL85]. solves [FR89, HNS88]. Solving [Bru81, EM85, FY80, GS81, Huỳ87, KO88, Lic86, MO87a, Pan88]. Some [BY87, BS86a, BM81, Bis84, CIRB87, CIR88, Cia86, Cou84, Dub85, Gil88, IMR81, IJR88, Ida83, IT88, JT85, Ko82, Muk89, NI80, Nak81, Ov81b, Oya84, Sze89a, Sze89b, Tag86, Tar82a, Tra84, Vol85, Wan80, AMA88, Arb88, CW82, EO82, Leu82, MC89b, Pan81a, SR88a, SV82, Szp88]. son [Oli80]. son-trees [Oli80]. Sort [Dob80, HS85, NP85, Chl88, Chl89, CK89, Dob82, Han86, HL87b, Pan86b, WS88]. sort-based [CK89]. Sorted [MA85, Goo87]. sorters [AFK88, PP89]. Sorting [AFK88, Dv82, GJ82, HR89, MA80, MG80, Mül87, Weg82b, van80a, AK81, BP85b, Bop89, Erk80, LP89b, PP89, Ric86, Wri87]. sources [Sha82a]. Space [Far83, MS86, Ryt86a, Ryt86b, SS87, Wil83a, Van89b, BP83, BDS82, Bur82b, CS82, CDL87, CIRB87, CIR88, Che86, EW86, EK89, Gol89, Gol80, GJ82, Hag88, Kaw80, KS87a, LT80, Nos85, Par87, Roh88, Sze87, Sze88, Sze89a, Sze89b, Tom81, WW87, ten80]. space-optimal [WW87]. spaces [Kob87, MC88, MC89a]. Spanning [LR86, NM82, Sto81, Tsi88, CFHP89, GT88, Gus83, JM88, Kut88, LP82, MP88, PR86, Tar82c, Tar86]. Sparse [JM86, Sch82b, Sha82b, AS86b, BT89, Coh83, Har83, HH88]. special [AFK88, Jan85, McK85]. special-purpose [AFK88]. specifiable [Tag86]. specification [Sis89]. Specifications [Bro87, BM82, Jos88]. Specified [EZ82]. specifying [Hol84]. speed [Die89]. speedup [CLM89]. Split [LL85, Ber84b, Wog89]. Splitmerge [Car86]. sporadic [CC89]. squad [Cul89]. Square [Car84, CD86, SV86, Vit85b, Car83, Gre89, Mai85]. Square-Free [Car84, CD86, Car83, Mai85]. squares [Bac82, BMS80, Ima82]. SSS [McG83]. Stabbing [Güt83, ARW89]. Stability [Kob83, Ber82, Day80b, MT87a]. stabilization [CGR87, Kes88]. Stable [All83, DD81, KK85, SS87, Bui77, Bui79, Car86, Day80a, GI89, Hul84, PG87]. Stables [CS83b]. Stack [WFL83, App87, Mye83, Ryt81a]. star [GMZ89, Lat83]. starry [LT88a]. starvation [AJV86, Mor79, Udd86]. starvation-free [AJV86, Mor79]. State [YG83, AK86, FG89, SR88a, Suz88]. state-dependent [SR88a]. Statement [Wil83c]. statements [Lif84]. states [Cha88]. static [Kor87a]. stationary [FS88a]. Statistical [Lei86, Row84, Sch84, Che82, Fal89]. statistics [CR88, GMZ89]. Steady [Man82]. Steady-paced-output [Man82]. Steiner [BP89, CH86, Flo91, Meh88, WI88]. Step [WB86]. Step-Size [WB86]. steps [CLW80a]. stiff [Bui77, Bui79, Day80a]. Stopping [Bab87b]. Storage [AR80a, DD81, Eva86, FP83, Rob80, Slu87, FL84, GS86, LS82a, Leu84, NM87, Slu85, TV81, Tru80, TST83, ten80, Van89b]. storage-size [FL84]. Store [Vit85b]. stores [Vit85a]. Storing [Bit85]. Straight [IMR81, Sto82, Dob82, Pan86b]. Straight-Line [IMR81]. strategies [RTV89]. strategy [Apo86, Auf88, BS89]. Streams [AB81, TJP87]. Strict [de 87, Kel87]. Strictness [Mis88]. strikes [Wal83]. String [AD86, LY86, Mey85, MNS84, BD80, Bar84, CC86, CR87, CC87, GL89, MO87a, SV87]. String-Matching [LY86, CR87]. string-rewriting [MO87a]. Strings [Par84, Apo86, CD89, Rob89b, Sem81]. Strong [Ata84, GM82, SY84a, Vis85, EM84, Sze89a, Tar82b, Tar83a, Tsi85, Boo81b, MF81]. structural [TJP87, UY88]. Structure [And80a, Gun82, Mil83, Wil83b, CCS89, Ev89, Fie87, FMNT87, LW82, Sch83a]. Structured [Pas80, DL88]. Structures [AT87, AMM84, Cza88, Cza89, Ev89, HS88b, KV81, LO89, MPY87, OS88, SO87, Sak88, Sun89, Szp88]. Study [CSO89, WK88, Wee88]. stuttering [ADS86, Mir87]. Sub [OSW81]. Sub-regular [OSW81]. Subclass [Oya84]. Subclasses [Kri86a, IJR88, Sli82]. subforests [Sla82]. subgraph [Miy88, Shi81, YG87]. subgraphs [Dil87b, Kan83]. subkeys [BD80]. sublogarithmic [Sze89b]. Subsequence [AD86, Apo86, Apo87, Mir87, RT85b]. Subsequences [Kal85]. Subset [AB87b, MSR85, TS86, Cas83]. subsets [EM84]. substitution [Boo88, Mor88c]. substrings [Boo80]. Subsumption [Got87b]. subtext [BD80]. Subtraction [Vol85]. Subtree [LK89, Gro91, Mäk89]. Subtrees [Hik83, Lyo88]. Subword [ER83b, ER83a, BU89, ER81]. succinct [Mic81]. sufficiency [Nij82a]. Sufficient [Kol83, Wil83c, MC89a]. Sum [TS86, Bac82, Rus87, Szp88, Toy87]. Sums [Er83c]. Super [Cal87, AS86b]. Super-exponentials [Cal87]. superconcentrator [BKV^{+}81]. superimposed [Won89]. Superincreasing [Wil83b]. surrogate [Won89]. sweep [HNS88]. sweeping [Mic81]. switch [Weg85]. switched [PM88]. switching [PS80, PS81a]. symbolic [BT87]. symbols [dV88]. Symmetric [BSZ89, BES82, BW87, GGPS89, GP87b, Oli80]. symmetrical [Kob87]. Symmetries [Hig86]. symmetrizability [CM80b]. symmetrizing [Col79, CM80b]. synchronization [FR89, Gos89, WU88]. Synchronizing [Dwe88]. Synchronous [Van81, Tur89]. synopsis [Con83]. syntactical [DI89b]. Syntax [JB84, EPS82]. syntax-free [EPS82]. Synthesis [GS84a, BSM^{+}88, Gin82, Mak81]. System [Cza80b, DK84, MF84, Orł88, Sar80, CP82, CC89, GS81, Jan85, Kaw80, Lei80, Toy89]. systematic [Wil81b]. Systems [AR80a, BB83, Bud85, CRL85, CL86c, DRT81, Llo80b, Mag84, Mag85, RAK85, Sha83, Tha85, Tir83, Žák84, ABM84, AK86, AC88, AR80b, BSM^{+}88, Bar80, Ber82, Ber87a, Boo81a, BL87, CM80a, Cha88, EY87, Fle89, Gei84, Gra80, HRY87, Kob83, KW82, Llo80a, Llo81, Lop89, Mag87, Mét83, Muk89, NO88, Ngu85, Och87, Pan88, Pla85, RBP82b, RH88, Rus87, Ryt82b, SSA89, SM87, Toy87, WU88, Yoc89, vW80a, Van88]. Systolic [CC83, Mül87, CK88b, LT88b, RT85b, SM87, ZM87]. Szymanski [Apo86].

T [Day80a]. T0L [SSA87]. table [OP85]. table-driven [OP85]. Tables [BDQ82, JV86, Pag85, AG81, Wog89]. tactile [Ber86a]. Tagless [Ham86, Lyo88]. tail [Kie86]. Take [Bis84]. Take-Grant [Bis84]. tale [LNS82]. Tampering [Kak85]. Tan [AG88b]. Tape [Vit85b, Die89, EL81, Vit85a]. tardiness [HS89]. Tarjan [Ban80]. task [Llo81]. Tasks [Fre83a, LM81, BWD84, CC89, DL88, Kub87, LM80a, SY93]. Tausworthe [Fus83]. technique [Jak82, KO88, Mir87, Rem84]. techniques [CLM89]. Tel [AG88b]. Temporal [Sza87, ZLS87]. term [Pla85, Rus87, Toy87, Toy89]. term-rewriting [Rus87]. termal [Sab80]. terminals [WW85]. terminating [Bar80]. Termination [DS80, DFv83, Hua89, Top84, ARG86, AG88b, BO86, HS88a, HZ87, Hua88a, KT88, Mat89b, Ran83, Rus87, TTv86, TV87, TM89, Toy87]. terms [Oza83, Pan81a]. ternary [PW87]. Terrain [Far80]. terrains [Sha88b]. tessellation [NI80]. Test [dPV88, AG84, CLW83, Mat89b, UY88, Wey84, Wil81b]. Testing [CH88a, HLY80, LD88, Mäk83, PS81b, WB86, APT79, APT82, Bha80, BKV^{+}81, Kor87a, LD80, ZG89]. th [Urb80]. Their [Cza80b, Mil83, Tra84, Wät85, Cza89, Pob86, ten80]. theorem [AB87a, BD82, BI88, Li88a, McL85, Mor89, Spy82, Tom81, YWL82, YWL85, RH87]. Theorems [BT82, ABC87, LNS82, Paj80]. Theoretic [Cza80a, Kel86, SS85b, Tou84]. Theory [Žák84, Cas83, Ham87, Heh81, Kie86, Rys83]. There [GGPS89, Sze87, FS88b]. thesis [Blu83]. threaded [Gor86]. Three [Dv82, EM85, Car83, Cha88, dV88, FMNT87, GH87, IT80, Pap85, YWL82, YWL85]. three-dimensional [Cha88, GH87]. three-restricted [YWL82, YWL85]. three-sided [FMNT87]. three-way [IT80]. threshold [Dun90, Koc89, LLH89].

Throughput [DEGS84]. Thue [dV88, Jan85, NO88]. Tight [ĎSVT85, SV87, GST89, Ó’D88, SE87]. Time [Ber84a, Clo83, DEGS84, Er83c, FS84, Fre83a, Fre84, GPB82, Gun82, IMM81, Jos85, LM85, Mor85, PV80, Roh88, Ryt81c, Ryt82b, Ryt83b, SS87, SN88, SV84, UMS82, Vit85b, Wel85, WW87, ZM87, AK89, AS86b, APT79, APT82, Ata83, Bab87a, BF87a, BP87, BF87b, BCW80, BT89, CS82, CYC89, CC89, Cla86, CL86b, DS88, Dev85, Dud83, EW86, EMM89, EL81, Far83, FL87, FL82, Gab81, GT88, GIR88, GPB83, GJ82, GL80, Heh89, Her89, HK88, HM81, KNT87, Kaw80, Kie86, LMT87, LO89, LM80a, LL88, Lot84, MU84, Min88, Mur83, Ola89, Pep87, Sab80, Sab81, Sch89, Sha82a, Sho89a, Sli82, SY93, Vit85a, Vř’t87, WT87, Wil84, Zip89]. Time- [WW87]. Time-optimal [SN88, LMT87]. time-space [CS82]. timeout [Ber87a]. Times [Fre83a, Bab87b, DL88, Sle80, SY93, Vit80]. timestamps [HM88]. Toeplitz [BP85c]. tolerance [ÖA88]. Tolerant [IM88, BJ80, DF88, Kut88, Var89b]. Tompa [KL87]. too [Jar81]. tool [Kři86b, Ray89b, Sli82]. Top [FV89b]. Top-down [FV89b]. topological [Rei87b, Sho89b]. Total [Ber86c, Kei86, RR88a, RSR80]. totalistic [CK88b]. tour [Bli88]. Towers [Atk81, BL80a, SS84a, Wal82, Wal83]. Trade [Pan86a]. Trade-Off [Pan86a]. tradeoff [Lot84]. tradeoffs [CS82, CL86b]. trails [Ebe88]. transaction [Kre87, Wol87]. Transactions [FR84]. Transducers [IR81, FV89b, Yeh85]. transductions [Hea89, LT88a]. transfer [BM82, Ber87a, Del83]. Transfers [ZBS83]. transform [Pob86]. Transformation [BH87, Bea89, Wit88]. Transformational [Bro80b]. Transformations [Cha86, ĎSVT85, Tra84, Zöb87, Ryt89]. transformer [MC80]. Transforms [Nja85]. transition [AS86a, AC88]. Transitive [JM86, Ebe81, IK83]. transitivity [Zan80]. Translation [Nos85, RSD81, Bur82b]. Transportation [Kir88, Yoo85]. Transshipment [Far80]. transversals [AB87c]. trap [IW81]. trap-door [IW81]. Trapdoor [Wil83b]. Trapdoors [Sha83]. Traveling [Dil87b, Kan83, AN84]. travelling [BH89]. traversal [KPS88]. traversals [CYL88]. Tree [HS86, Hik83, Kri86a, LR86, NM82, Pit87, Sab81, Sao84, Sil81, Tar83b, TS81, YJ83, Bre80, CS83a, CYL88, CFHP89, CA86, CW82, DLF87, DL88, Eme87, FV89b, IJ89, MP88, OIM84, PK88, Ram87, San88, San89b, SM92, Sla82, Szp86, WI88, Wil81a, VV87]. Tree-Like [Kri86a]. tree-structured [DL88]. tree-to-tree [Wil81a]. tree-type [Szp86]. Trees [CH86, EZ82, Eva86, Far80, Kor81, KK83, Lea88, MS86, Reg81, Ric83, Sei88, Shi83, Sto81, Tsi88, vR88, BY87, BI88, Che86, CR89, Eas81, FZ82, GIR88, Gor86, Gre89, Gro91, GP87b, GS86, GH87, Gus83, HL88, HL89, HY82, IRV88, Jar81, JM88, Kem87, Kem89, Kor82, Lei88, LT80, Leu84, LP89c, MS81, Mäk88b, Mäk89, Mos82, Oli80, Pal87, PR86, PW87, Sav80, Sch89, Sch86a, Tam83, Tar82c, Tar86, Zak82]. triangle [De 82]. Triangular [BP85c, KLV87]. triangulating [WT87]. Triangulation [Dil87a, Gol89, Lev87]. triangulations [Dil87b, Kir80, Lin86]. Tridiagonal [CL86c, EY87]. Tries [Ore82]. trigonometric [Sch88]. Trinity [Hou87]. true [Pet91, VT90]. truth [APT79, APT82, CL88b]. Tsao [Bin82]. TSP [BH89]. Tuning [LR80]. Tuples [KR85]. Turing [CIRB87, CIR88, CC84, Deh86b, Die89, ITT82, Kir88, Par87, Pas80]. Turner [Bur82b, Nos85]. Two [ABC87, BT82, Car84, CYL88, GT83, Gir86, Ham86, Hof83, JV86, LM85, LY86, Lis81, Mic81, Miy84, Mon84, Ryt83b, SV84, Tha85, UMS82, AN86, And79, And80b, Apo86, AAI87, Bac82, Bar84, Byk78, CDS89, Cou81, DS79, ER80, Fou79, Fra83, Fre80, Gru84, IT80, Jan89, Kem89, Kub87, Lic86, LT88b, NI80, Ov80, Paj80, Ryt81b, Ryt82a, Ryt83a, SSA89, Sch81b, Sle80, TST83, Var89a, Vit85a, pW82, WK88, Yap83]. two-dimensional [AN86, Fre80, IT80, LT88b, NI80, pW82]. Two-Head [LY86]. Two-Headed [Hof83]. two-level [Gru84, TST83]. two-processor [Jan89, Kub87, SSA89]. Two-Way [Miy84, Mon84, Ryt83b, SV84, UMS82, Mic81, Ryt81b, Ryt82a, Ryt83a, Var89a]. Type [Gun82, Lei86, Baa88, Szp86]. types [GS88b, Lau80, Lis81]. typing [Mar83].

UET [Llo80a, Llo80b, Llo81, vW80a]. ultrametric [Kři88b]. Un-Skolemizing [McC88]. Unambiguous [Mäk83, Ryt82b]. Unary [Cha81]. unavoidable [CLW83]. Unbounded [Hro85b]. Unboundedness [YG83, CH88a]. uncertainty [KPK89]. Undecidability [WW89b, Boo81b, Pla85]. undecidable [Abi89, Mäk85b, NO88]. undefined [Ten87a]. undercover [Mäk85a]. underlying [Fal89]. understanding [ACK89]. Undirected [Ata84, Ebe80, KK89a, Tsi85]. unfair [MB85]. Unifiability [LS82b]. Unification [MN82, Baa88, CDS89, EL81, LBM88]. unified [RR88b]. Uniform [BB83, Eme87, ZBS83, BF87a, BF87b, Bod88b, GZW85, Par87]. Uniformly [LBB81, LN88]. unifying [CMT88]. unimodal [Ola89]. unimodality [Ola88a]. union [Ban80, Mei85]. Unique [Ryt86b, AG84, Var89b]. unique-path [Var89b]. uniqueness [MC88, MC89a]. Unit [Fre83a, Gre89, Vit85b, BDS82, Kři86b, Min88, SY93, YYDH84]. Unit-length [Gre89]. Unit-Time [Fre83a, SY93]. Universal [KPS88, Łuk82, Sar80, HLY80, MGKL88].

Unmerging [SS87]. unordered [SV88]. unreliable [BO86]. unstructuredness [Tse87]. Untyped [Mis88]. unwind [Kfo85]. Update [HV87, PR86, Vos88]. Update/Query [HV87]. Updating [AD83, Tar83b, Tsi88, JM88]. Upper [Ukk85, BU89, BS86b, CP81, Her89, LP89c, Mot82, dV89]. Use [BK84, KF85, AFK88, Bar83, Mur83, Tsi89]. Used [Ore82]. useful [Yun85]. user [IW81]. Using [Ber84a, CYC89, Clo83, EZ82, Ev89, MO87a, Mag84, Mag85, Mak81, Mak86a, Mül87, MS86, AS83, Bar83, Ber86a, BGLY81, CV89, CGPD89, CGPdS89, CIRB87, CIR88, FP87, GS88b, Hua89, Mag87, PM88, Ram87, Tar82b, Tar83a, Udd86, Wri87]. utilisation [Leu84]. utilization [GS86].

Vaidya [Iwa89]. Valuation [HV87]. Value [CW86, Gol80]. Values [AM84, Sik81, Ten87a]. Variable [Clo83, KS87a, WB86, Boo88]. variable-renumbering [Boo88]. Variables [LM85, Bur89, CDS89, McC89, Mor88a]. Variant [Sei88, Car87b, Kes88]. Variants [Bis84, DT80, SS84b]. Variation [Dob80, BL80b]. Variations [Cul89, Ham86, IS87]. Varieties [BI88, Mor87b]. various [Mos82]. Vector [TS81, Avi80, Cha88, Gra80, HRY87, Jak82, van80b]. vectors [Buc89, Dev80]. verification [AK86, Lif84]. Verifying [GK86a, Mil83, Hol84, Sis89, WS82]. versus [Bur89]. Vertex [Sav82, Tsi88, BDD^{+}82, BS89, BDP88, CW84, Cla83, DM89]. vertically [LJ86]. vertices [PS88]. Very [Mon84, Won89]. via [BP85c, Dev80, Han80, IM81, LR80, Pan88, Tsi89]. Victor [Zem82]. view [Hul84, Muk89]. view-based [Muk89]. Virtual [RSD81, BJ80, Kel89]. Visibility [Wel85, MKK89, SE87]. VLSI [ABM84, BSM^{+}88, BC87, BP85b, CRL85, CL86b, CL86a, CL86c, Hoc87, Jin84, Joh80, Lot84, PW87, PV80, Sch86a, Tho85, Vř’t87]. Voronoı [Lin89, OIM84, Kan83]. Voronoi-diagram [OIM84]. Voronoıdiagram [Aur86]. Vs [HP85, RG89a].

wait [FR89]. wait-free [FR89]. Waite [Der80, Ham86]. walks [Pro89]. Walsh [ĎSVT85]. Walsh-Hadamard [ĎSVT85]. watchman [CN88]. watchmen [Nta86]. watchtower [Sha88b]. Way [LY86, Miy84, Mon84, Ryt83b, Ryt84, SV84, UMS82, AGvG89, CR87, Hås87, IW81, IT80, Mic81, Ryt81b, Ryt81c, Ryt82a, Ryt83a, Sav88, Var89a, Wat88]. Weak [Miy84, Yam81, EL81, FP87, Jar81, Sze89a, Udd86]. Weak-Counters [Miy84]. Weakest [Orł88, HH87, Mor87b]. Weakly [Car84]. Wegman [Sar80]. Weight [IMM81, BP83, Dil87b, SH89]. Weighted [AB87a, Shi83, Aur86, BF87a, BF87b, DF84, EMM89, HS89]. Well [Mor89]. Well-founded [Mor89]. Welsh [OR89]. Welsh-Powell [OR89]. Where [Kol83]. Where-Oblivious [Kol83]. whether [BKV^{+}81]. Which [Gra87, SS85b, Mur83, Sen81, Sky83]. whose [dV89]. Width [Yas81, Par87]. Wiltink [Bal87]. window [MKK89]. wire [BC87]. without [ADS86, BK84, CK84, Eis89, IT88, Lat83, SO87, Sak88, Wil83b]. Wood [Žák84]. Word [BS86a, BO85, Gra87, BB87, Boo81b, BL87, Cro81, dV88, MO87a, Rot89, Won89]. Words [BP85a, Car84, AFG87, CD86, CLW83, Gir86, de 81, dPV88]. world [Lau87]. Worst [JM86, KK85, Ov81a, PG87, Sun89, WI88, Wil83a, Apo86, KS87b, Röh82]. Worst-Case [JM86, KK85, Wil83a, Ov81a, PG87, Sun89, WI88, Apo86, KS87b]. writer [Vid89]. writers [Kes80]. Writing [Bac82].

Yes [Hof83]. yields [Ber86b]. zero [Baa88]. Zigzag [Atk85]. Zooming [EO87].

References

Asano:1987:SPB

[AAI87] Takao Asano, Tetsuo Asano, and Hiroshi Imai. Shortest path between two simple polygons. Information Processing Letters, 24(5): 285–288, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ancilotti:1981:ISS

[AB81] P. Ancilotti and M. Boari. Information streams sharing a finite buffer: Protection problems. Information Processing Letters, 12(3): 151–159, June 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alexandrakis:1987:WGK

[AB87a] Athanasios Alexandrakis and Symeon Bozapalidis. Weighted grammars and Kleene’s theorem. Information Processing Letters, 24(1):1–4, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arvind:1987:ASP

[AB87b] V. Arvind and S. Biswas. An O(n^{2}) algorithm for the satisfiability problem of a subset of propositional sentences in CNF that includes all Horn sentences. Information Processing Letters, 24(1):67–69, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atallah:1987:EAC

[AB87c] Mikhail Atallah and Chanderjit Bajaj. Efficient algorithms for common transversals. Information Processing Letters, 25(2):87–91, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apt:1987:TNF

[ABC87] K. R. Apt, Luc Bougé, and Ph. Clermont. Two normal form theorems for CSP programs. Information Processing Letters, 26(4):165–171, December 4, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abiteboul:1989:BUD

[Abi89] Serge Abiteboul. Boundedness is undecidable for Datalog programs with a single recursive rule. Information Processing Letters, 32(6): 281–287, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alia:1984:FVC

[ABM84] G. Alia, F. Barsi, and E. Martinelli.

A fast VLSI conversion between binary and residue systems. Information Processing Letters, 18(3):141– 145, March 30, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1987:CNM

[AC87] M. D. Atkinson and H. W. Chang.

Computing the number of mergings with constraints. Information Processing Letters, 24(5):289–292, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arnold:1988:LAS

[AC88] André Arnold and Paul Crubille. A linear algorithm to solve fixed-point equations on transition systems. Information Processing Letters, 29(2): 57–66, September 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atallah:1989:OPA

[AC89] Mikhail J. Atallah and Danny Z. Chen. An optimal parallel algorithm for the minimum circle-cover problem. Information Processing Letters, 32(4):159–165, September 1, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1989:GMU

[ACK89] Alok Aggarwal, Don Coppersmith, and Dan Kleitman. A generalized model for understanding evasiveness. Information Processing Letters, 30(4):205–208, February 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Agrawal:1983:UHD

[AD83] Rakesh Agrawal and David J. Dewitt. Updating hypothetical data bases. Information Processing Letters, 16(3):145–146, April 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1986:BLA

[AD86] Lloyd Allison and Trevor I. Dix. A bit-string longest-common-subsequence algorithm. Information Processing Letters, 23(6):305– 310, December 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alpern:1986:SS

[ADS86] Bowen Alpern, Alan J. Demers, and Fred B. Schneider. Safety without stuttering. Information Processing Letters, 23(4):177–180, November 8, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Autebert:1987:PIW

[AFG87] Jean-Michel Autebert, Philippe Flajolet, and Joaquim Gabarró. Prefixes of infinite words and ambiguous context-free languages. Information Processing Letters, 25(4): 211–216, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atallah:1988:SEU

[AFK88] Mikhail J. Atallah, Greg N. Frederickson, and S. Rao Kosaraju. Sorting with efficient use of specialpurpose sorters. Information Processing Letters, 27(1):13–15, February 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ancona:1981:NMI

[AG81] M. Ancona and V. Gianuzzi. A new method for implementing LR(k) tables. Information Processing Letters, 13(4–5):171–176, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apostolico:1984:PMM

[AG84] A. Apostolico and R. Giancarlo.

Pattern matching machine implementation of a fast test for unique decipherability. Information Processing Letters, 18(3):155–158, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anderson:1988:ASN

[AG88a] James H. Anderson and Mohamed G. Gouda. Atomic semantics of nonatomic programs. Information Processing Letters, 28(2): 99–103, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arora:1988:MCD

[AG88b] R. K. Arora and M. N. Gupta. More comments on: “Distributed termination detection algorithm for distributed computations” [Arora, Gupta and S. P. Rana, Inform. Process. Lett. 22 (1986), no. 6, 311– 314; R. B. Tan, G. Tel and J. van Leeuwen, ibid. 23 (1986), no. 3, 163; MR 87j:68034a b]. Information Processing Letters, 29(1):53– 55, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [TTv86, ARG86].

Akl:1989:OPA

[AGS89] Selim G. Akl, David Gries, and Ivan Stojmenovíc. An optimal parallel algorithm for generating combinations. Information Processing Letters, 33(3):135–139, November 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Abiteboul:1989:AWR

[AGvG89] Serge Abiteboul, Marc Gyssens, and Dirk van Gucht. An alternative way to represent the cogroup of a relation in the context of nested databases. Information Processing Letters, 32(6):317–324, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apostolico:1987:FMC

[AH87] Alberto Apostolico and Susanne E.

Hambrusch. Finding maximum cliques on circular-arc graphs. Information Processing Letters, 26(4): 209–215, December 4, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1989:CCB

[AH89] Alok Aggarwal and Michael Hawrylycz. On computing the closest boundary point on the convex hull. Information Processing Letters, 31 (6):311–314, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Annot:1986:CMS

[AJV86] J. K. Annot, M. D. Janssens, and A. J. Van De Goor. Comments on Morris’s starvation-free solution to the mutual exclusion problem [Inform. Process. Lett. 8(2), 15 February 1979, pp. 76–80]. Information Processing Letters, 23(2):91–97, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Mor79].

Atallah:1981:ABL

[AK81] Mikhail J. Atallah and S. Rao Kosaraju. An adversary-based lower bound for sorting. Information Processing Letters, 13(2):55– 57, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Auzins:1985:SEE

[AK85] A. J. Auzins and E. B. Kinber. On separation of the emptiness and equivalence problems for program schemes. Information Processing Letters, 20(2):91–93, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apt:1986:LAV

[AK86] Krzysztof R. Apt and Dexter C. Kozen. Limits for automatic verification of finite-state concurrent systems. Information Processing Letters, 22(6):307–309, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aggarwal:1989:LTA

[AK89] Alok Aggarwal and Dina Kravets. A linear time algorithm for finding all farthest neighbors in a convex polygon. Information Processing Letters, 31(1):17–20, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Akl:1984:OAP

[Akl84] Selim G. Akl. An optimal algorithm for parallel selection. Information Processing Letters, 19 (1):47–50, July 26, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/0020019084901285.

Akman:1987:ADO

[Akm87] Varol Akman. An algorithm for determining an opaque minimal forest of a convex polygon. Information Processing Letters, 24(3):193– 198, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alblas:1989:OIS

[Alb89] Henk Alblas. Optimal incremental simple multi-pass attribute evaluation. Information Processing Letters, 32(6):289–295, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1983:SMC

[All83] Lloyd Allison. Stable marriages by coroutines. Information Processing Letters, 16(2):61–65, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1989:DSE

[All89] Lloyd Allison. Direct semantics and exceptions define jumps and coroutines. Information Processing Letters, 31(6):327–330, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atzeni:1984:FDR

[AM84] Paolo Atzeni and Nicola M. Morfuni. Functional dependencies in relations with null values. Information Processing Letters, 18(4):233– 238, May 14, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Anderson:1987:PMP

[AM87] Richard Anderson and Ernst W.

Mayr. Parallelism and the maximal path problem. Information Processing Letters, 24(2):121–126, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aguilar-Martin:1988:CSR

[AMA88] J. Aguilar-Martin and Claudi Alsina. Characterizations of some rescaling functions. Information Processing Letters, 28(3):127–132, July 4, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alt:1984:PMR

[AMM84] Helmut Alt, Kurt Mehlhorn, and J. Ian Munro. Partial match retrieval in implicit data structures. Information Processing Letters, 19 (2):61–65, August 31, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1980:SDP

[AN80] D. C. S. Allison and M. T. Noga. Selection by distributive partitioning. Information Processing Letters, 11 (1):7–8, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Allison:1984:TSP

[AN84] D. C. S. Allison and M. T. Noga. The L_{1} traveling salesman problem. Information Processing Letters, 18 (4):195–199, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aizawa:1986:DIA

[AN86] Kunio Aizawa and Akira Nakamura. Direction-independent application of productions on twodimensional arrays. Information Processing Letters, 22(6):295–301, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andrew:1979:AEA

[And79] A. M. Andrew. Another efficient algorithm for convex hulls in two dimensions. Information Processing Letters, 9(5):216–219, December 16, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [And80b].

Anderson:1980:EPI

[And80a] Bruce Anderson. Encoded pointers — an interesting data-structure for modern SIL’s. Information Processing Letters, 10(2):47–50, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Andrew:1980:CAE

[And80b] A. M. Andrew. Corrigendum: “Another efficient algorithm for convex hulls in two dimensions” (Inform. Process. Lett. 9 (1979), no. 5, 216–219). Information Processing Letters, 10(3):168, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [And79].

Anstee:1987:PAM

[Ans87] Richard P. Anstee. A polynomial algorithm for b-matchings: an alternative approach. Information Processing Letters, 24(3):153–157, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apostolico:1986:IWC

[Apo86] Alberto Apostolico. Improving the worst-case performance of the Hunt-Szymanski strategy for the longest common subsequence of two strings.

Information Processing Letters, 23 (2):63–69, August 20, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Apostolico:1987:RHN

[Apo87] Alberto Apostolico. Remark on the Hsu–Du new algorithm for the longest common subsequence problem. Information Processing Letters, 25(4):235–236, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Appel:1987:GCC

[App87] Andrew W. Appel. Garbage collection can be faster than stack allocation. Information Processing Letters, 25(4):275–279, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aspvall:1979:LTA

[APT79] Bengt Aspvall, Michael F. Plass, and Robert Endre Tarjan. A lineartime algorithm for testing the truth of certain quantified Boolean formulas. Information Processing Letters, 8(3):121–123, March 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [APT82].

Aspvall:1982:ELT

[APT82] Bengt Aspvall, Michael F. Plass, and Robert Endre Tarjan. Erratum: “A linear-time algorithm for testing the truth of certain quantified Boolean formulas” [Inform. Process. Lett. 8 (1979), no. 3, 121–123; MR 80b:68050]. Information Processing Letters, 14(4):195, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [APT79].

Arora:1980:AMA

[AR80a] R. K. Arora and S. P. Rana. Analysis of the module assignment problem in distributed computing systems with limited storage. Information Processing Letters, 10(3):111– 115, April 18, 1980. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arora:1980:HAP

[AR80b] R. K. Arora and S. P. Rana. Heuristic algorithms for process assignment in distributed computing systems. Information Processing Letters, 11(4–5):199–203, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arbib:1988:PCS

[Arb88] Claudio Arbib. A polynomial characterization of some graph partitioning problems. Information Processing Letters, 26(5):223–230, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arora:1986:DTD

[ARG86] R. K. Arora, S. P. Rana, and M. N.

Gupta. Distributed termination detection algorithm for distributed computations. Information Processing Letters, 22(6):311–314, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [TTv86, AG88b].

Avis:1989:LBL

[ARW89] D. Avis, J. M. Robert, and R. Wenger. Lower bounds for line stabbing. Information Processing Letters, 33(2):59–62, November 10, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Arora:1981:GPD

[AS81] R. K. Arora and N. K. Sharma. Guarded procedure: a distributed programming concept. Information Processing Letters, 13(4–5):199–203, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alpern:1983:KEU

[AS83] B. Alpern and F. B. Schneider. Key exchange using keyless cryptography. Information Processing Letters, 16(2):79–81, February 26, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alpern:1985:DL

[AS85] Bowen Alpern and Fred B. Schneider. Defining liveness. Information Processing Letters, 21(4):181– 185, October 7, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alanko:1986:CCT

[AS86a] Timo O. Alanko and R. L. Smelianski. On the calculation of control transition probabilities in a program. Information Processing Letters, 22(5):273–276, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ambos-Spies:1986:IPT

[AS86b] Klaus Ambos-Spies. Inhomogeneities in the polynomial-time degrees: the degrees of super sparse sets. Information Processing Letters, 22(3):113–117, March 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ambos-Spies:1986:NCP

[AS86c] Klaus Ambos-Spies. A note on complete problems for complexity classes. Information Processing Letters, 23(5):227–230, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1988:PAB

[AS88] Michael D. Atkinson and N. Santoro. A practical algorithm for Boolean matrix multiplication. Information Processing Letters, 29(1): 37–38, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Asai:1983:CPI

[Asa83] Hitohisa Asai. A consideration of a practical implementation for a new convergence division. Information Processing Letters, 17(5):273– 281, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Agrawala:1981:OSR

[AT81] A. K. Agrawala and S. K. Tripathi.

On the optimality of semidynamic routing schemes. Information Processing Letters, 13(1):20–22, October 27, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Alexandridis:1987:ESE

[AT87] N. A. Alexandridis and P. D. Tsanakas. An encoding scheme for the efficient representation of hierarchical image structures. Information Processing Letters, 25(3):199– 206, May 29, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atallah:1983:LTA

[Ata83] Mikhail J. Atallah. A linear time algorithm for the Hausdorff distance between convex polygons. Information Processing Letters, 17(4):207– 209, November 8, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atallah:1984:PSO

[Ata84] Mikhail J. Atallah. Parallel strong orientation of an undirected graph. Information Processing Letters, 18 (1):37–39, January 20, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1981:CTH

[Atk81] M. D. Atkinson. The cyclic Towers of Hanoi. Information Processing Letters, 13(3):118–119, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Atkinson:1985:ZPC

[Atk85] M. D. Atkinson. On zigzag permutations and comparisons of adjacent elements. Information Processing Letters, 21(4):187–189, October 7, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Auffray:1988:LSP

[Auf88] Yves Auffray. Linear strategy for propositional modal resolution. Information Processing Letters, 28(2): 87–92, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Aurenhammer:1986:ODW

[Aur86] Franz Aurenhammer. The onedimensional weighted Voronoĭdiagram. Information Processing Letters, 22 (3):119–123, March 3, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Avis:1980:CLB

[Avi80] David Avis. Comments on a lower bound for convex hull determination: “On the Ω(n log n) lower bound for convex hull and maximal vector determination” by van Emde Boas [Inform. Process. Lett. 10(3), 18 April 1980, pp. 132–136]. Information Processing Letters, 11(3): 126, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [van80b].

Avis:1984:NPS

[Avi84] David Avis. Non-partitionable point sets. Information Processing Letters, 19(3):125–129, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Awerbuch:1985:NDD

[Awe85] Baruch Awerbuch. A new distributed depth-first-search algorithm. Information Processing Letters, 20(3):147–150, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ben-Ari:1980:SPR

[BA80] Mordechai Ben-Ari. A simplified proof that regular resolution is exponential. Information Processing Letters, 10(2):96–98, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baader:1988:NUT

[Baa88] Franz Baader. A note on unification type zero. Information Processing Letters, 27(2):91–93, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Babai:1987:ROS

[Bab87a] László Babai. Random oracles separate PSPACE from the polynomial-time hierarchy. Information Processing Letters, 26(1):51– 53, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Babaoglu:1987:STD

[Bab87b] Ozalp Babaoglu. Stopping times of distributed consensus protocols: A probabilistic analysis. Information Processing Letters, 25(3):163– 169, May 29, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Backhouse:1982:WNS

[Bac82] Ronald Backhouse. Writing a number as the sum of two squares: a new solution. Information Processing Letters, 14(1):15–17, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Balcazar:1986:I

[Bal86] José L. Balcázar. On Δ_{2}^{P}-immunity.

Information Processing Letters, 23 (1):25–28, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Balsters:1987:CDN

[Bal87] H. Balsters. Comments on: “A deficiency of natural deduction” [Inform. Process. Lett. 25 (1987), no. 4, 233–234; MR 88i:03020a] by J. G. Wiltink. Information Processing Letters, 26(3):163–164, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Wil87].

Banachowski:1980:CTR

[Ban80] Lech Banachowski. A complement to Tarjan’s result about the lower bound on the complexity of the set union problem. Information Processing Letters, 11(2):59–65, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barnden:1980:CSD

[Bar80] J. A. Barnden. A characterization of systems derived from terminating concurrent histories. Information Processing Letters, 10(3): 148–152, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barth:1981:AIK

[Bar81] Gerhard Barth. An alternative for the implementation of the Knuth– Morris–Pratt algorithm. Information Processing Letters, 13(4–5):

134–137, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barnard:1983:RDP

[Bar83] D. T. Barnard. Recursive descent parsing using implementation languages requiring definition before use. Information Processing Letters, 17(5):255–258, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barth:1984:ACT

[Bar84] Gerhard Barth. An analytical comparison of two string searching algorithms. Information Processing Letters, 18(5):249–256, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1983:PSP

[BB83] Alan A. Bertossi and Maurizio A. Bonuccelli. Preemptive scheduling of periodic jobs in uniform multiprocessor systems. Information Processing Letters, 16(1):3–6, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1986:HCI

[BB86] Alan A. Bertossi and Maurizio A. Bonuccelli. Hamiltonian circuits in interval graph generalizations. Information Processing Letters, 23(4): 195–200, November 8, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berstel:1987:LWC

[BB87] Jean Berstel and Srečko Brlek. On the length of word chains. Information Processing Letters, 26(1):23– 28, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berman:1988:PBD

[BB88] Joel Berman and Willem J. Blok.

Positive Boolean dependencies. Information Processing Letters, 27(3): 147–150, March 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bhatt:1987:CMW

[BC87] Sandeep N. Bhatt and Stavros S.

Cosmadakis. The complexity of minimizing wire lengths in VLSI layouts. Information Processing Letters, 25(4):263–267, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bossi:1983:DCA

[BCC83] A. Bossi, N. Cocco, and L. Colussi.

A divide-and-conquer approach to general context-free parsing. Information Processing Letters, 16(4): 203–208, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blum:1980:EFB

[BCW80] Manuel Blum, Ashok K. Chandra, and Mark N. Wegman. Equivalence of free Boolean graphs can be decided probabilistically in polynomial time. Information Processing Letters, 10(2):80–82, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bailey:1980:FSS

[BD80] T. A. Bailey and R. G. Dromey. Fast string searching by finding subkeys in subtext. Information Processing Letters, 11(3):130–133, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Balcazar:1982:NTL

[BD82] J. L. Balcazar and J. Diaz. A note on a theorem by Ladner. Information Processing Letters, 15(2):84– 86, September 6, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Becker:1982:PAV

[BDD^{+}82] M. Becker, W. Degenhardt, J. Doenhardt, S. Hertel, G. Kaninke, W. Keber, K. Mehlhorn, S. Näher, H. Rohnert, and T. Winter. A probabilistic algorithm for vertex connectivity of graphs. Information Processing Letters, 15(3):135– 136, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bovet:1988:PFV

[BDP88] Daniel P. Bovet, S. De Agostino, and R. Petreschi. Parallelism and the feedback vertex set problem. Information Processing Letters, 28(2): 81–85, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bermond:1982:TLG

[BDQ82] J.-C. Bermond, C. Delorme, and J.- J. Quisquater. Tables of large graphs with given degree and diameter. Information Processing Letters, 15(1): 10–13, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bron:1982:MMU

[BDS82] C. Bron, E. J. Dijkstra, and S. D.

Swierstra. A memory management unit for the optimal exploitation of a small address space. Information Processing Letters, 15(1):20–22, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Beaudry:1989:CIT

[Bea89] Martin Beaudry. Characterization of idempotent transformation monoids. Information Processing Letters, 31(3):163–166, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blumer:1987:ORL

[BEHW87] Anselm Blumer, Andrzej Ehrenfeucht, David Haussler, and Manfred K. Warmuth. Occam’s razor. Information Processing Letters, 24 (6):377–380, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1980:CSJ

[Ber80] Alan A. Bertossi. On the complexity of scheduling jobs on dedicated resources to minimize set-up costs. Information Processing Letters, 10(4– 5):173–177, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1981:EHP

[Ber81] Alan A. Bertossi. The edge Hamiltonian path problem is NP-complete. Information Processing Letters, 13 (4–5):157–159, End ??, 1981. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Beretta:1982:MCE

[Ber82] G. Beretta. Monte Carlo estimation of numerical stability in fast algorithms for systems of bilinear forms. Information Processing Letters, 14 (4):162–167, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1983:FHC

[Ber83] Alan A. Bertossi. Finding Hamiltonian circuits in proper interval graphs. Information Processing Letters, 17(2):97–101, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berkowitz:1984:CDS

[Ber84a] Stuart J. Berkowitz. On computing the determinant in small parallel time using a small number of processors. Information Processing Letters, 18(3):147–150, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1984:DSS

[Ber84b] Alan A. Bertossi. Dominating sets for split and bipartite graphs. Information Processing Letters, 19 (1):37–40, July 26, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/0020019084901261.

Bernstein:1986:DSC

[Ber86a] Herbert J. Bernstein. Determining the shape of a convex n-sided polygon by using 2n + k tactile probes. Information Processing Letters, 22 (5):255–260, April ??, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berstel:1986:EIM

[Ber86b] Jean Berstel. Every iterated morphism yields a co-CFL. Information Processing Letters, 22(1):7–9, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1986:TDI

[Ber86c] Alan A. Bertossi. Total domination in interval graphs. Information Processing Letters, 23(3):131–134, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bernstein:1987:PTT

[Ber87a] A. J. Bernstein. Predicate transfer and timeout in message passing systems. Information Processing Letters, 24(1):43–52, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berztiss:1987:NDO

[Ber87b] Alfs T. Berztiss. A notation for distributed operations. Information Processing Letters, 26(1):19– 21, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1988:DNI

[Ber88a] Alan A. Bertossi. On the domatic number of interval graphs. Information Processing Letters, 28(6):275– 280, August 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertossi:1988:PCA

[Ber88b] Alan A. Bertossi. Parallel circlecover algorithms. Information Processing Letters, 27(3):133–139, March 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barlow:1981:PMA

[BES81] R. H. Barlow, D. J. Evans, and J. Shanehichi. A parallel merging algorithm. Information Processing Letters, 13(3):103–106, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barlow:1982:PMD

[BES82] R. H. Barlow, D. J. Evans, and J. Shanehchi. Parallel multisection for the determination of the eigenvalues of symmetric quindiagonal matrices. Information Processing Letters, 14(3):117–118, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Best:1984:EFC

[Bes84a] Eike Best. Erratum: “Fairness and conspiracies” [Inform. Process. Lett. 18 (1984), no. 4, 215–220; MR 85h:68053]. Information Processing Letters, 19(3):162, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Bes84b].

Best:1984:FC

[Bes84b] Eike Best. Fairness and conspiracies.

Information Processing Letters, 18 (4):215–220, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [Bes84a].

Bazewicz:1987:MMW

[BF87a] J. Bazewicz and G. Finke. Minimizing mean weighted execution time loss on identical and uniform processors. Information Processing Letters, 24(4):A259–A263, March ??, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blazewicz:1987:MMW

[BF87b] J. Błażewicz and G. Finke. Minimizing mean weighted execution time loss on identical and uniform processors. Information Processing Letters, 24(4):259–263, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bermond:1987:EMI

[BFJM87] Jean Claude Bermond, Jean Michel Fourneau, and Alain Jean-Marie. Equivalence of multistage interconnection networks. Information Processing Letters, 26(1):45– 50, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blass:1989:MNA

[BG89] Andreas Blass and Yuri Gurevich.

On Matijasevitch’s nontraditional approach to search problems. Information Processing Letters, 32(1):41– 45, July 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Borodin:1981:ESU

[BGLY81] A. Borodin, L. J. Guibas, N. A. Lynch, and A. C. Yao. Efficient searching using partial ordering. Information Processing Letters, 12(2): 71–75, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bird:1987:AAE

[BH87] R. S. Bird and John Hughes. The alpha-beta algorithm: an exercise in program transformation. Information Processing Letters, 24(1):53– 57, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Basart:1989:AAT

[BH89] J. M. Basart and L. Huguet. An approximation algorithm for the TSP (travelling salesman problem). Information Processing Letters, 31(2): 77–81, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bhat:1980:CTG

[Bha80] Kabekode V. S. Bhat. On the complexity of testing a graph for ncube. Information Processing Letters, 11(1):16–19, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boppana:1987:DCN

[BHZ87] Ravi B. Boppana, Johan Håstad, and Stathis Zachos. Does co-NP have short interactive proofs? Information Processing Letters, 25(2): 127–132, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Beynon:1985:CBF

[BI85] W. M. Beynon and C. S. Iliopoulos. Computing a basis for a finite Abelian p-group. Information Processing Letters, 20(3):161–163, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bozapalidis:1988:VFS

[BI88] Symeon Bozapalidis and Stavros Ioulidis. Varieties of formal series on trees and Eilenberg’s theorem. Information Processing Letters, 29(4): 171–175, November 12, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bini:1982:RPN

[Bin82] Dario Bini. Reply to the paper: “The numerical instability of Bini’s algorithm” [Inform. Process. Lett. 12 (1981), no. 1, 17–19; MR 82i:65029] by N. K. Tsao. Information Processing Letters, 14(3):144– 145, May 16, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Tsa81].

Bird:1981:JP

[Bir81] R. S. Bird. The jogger’s problem. Information Processing Letters, 13(3):114–117, December 13, 1981.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Birkhoff:1985:DRA

[Bir85] Ingrid J. M. Birkhoff. A direct routing algorithm for the bit-reversal permutation on a shuffle-exchange network. Information Processing Letters, 21(5):259–268, November 18, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biskup:1984:SVT

[Bis84] Joachim Biskup. Some variants of the take-grant protection model. Information Processing Letters, 19(3): 151–156, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Biskup:1989:BCN

[Bis89] Joachim Biskup. Boyce-Codd normal form and object normal forms. Information Processing Letters, 32 (1):29–33, July 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bitner:1985:SMD

[Bit85] James R. Bitner. Storing matrices on disk for efficient row and column retrieval. Information Processing Letters, 20(5):249–254, June 12, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandwajn:1980:SFT

[BJ80] Alexandre Brandwajn and Rene Joly. A scheme for a fault-tolerant virtual memory. Information Processing Letters, 10(2):99–103, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bechtold:1984:UEH

[BK84] Ulrich Bechtold and Klaus Küspert.

On the use of extendible hashing without hashing. Information Processing Letters, 19(1):21– 26, July 26, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/0020019084901236.

Brassard:1988:GRP

[BK88] Gilles Brassard and Sampath Kannan. The generation of random permutations on the fly. Information Processing Letters, 28(4):207–212, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blum:1981:CTW

[BKV^{+}81] M. Blum, R. M. Karp, O. Vornberger, C. H. Papadimitriou, and M. Yannakakis. The complexity of testing whether a graph is a superconcentrator. Information Processing Letters, 13(4–5):164–167, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buneman:1980:THP

[BL80a] Peter Buneman and Leon Levy. The Towers of Hanoi problem. Information Processing Letters, 10(4–5): 243–244, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burton:1980:RVI

[BL80b] F. W. Burton and G. N. Lewis. A robust variation of interpolation search. Information Processing Letters, 10(4–5):198–201, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Book:1987:RSW

[BL87] Ronald V. Book and Hai-Ning Liu. Rewriting systems and word problems in a free partially commutative monoid. Information Processing Letters, 26(1):29–32, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Barkley:1989:PRH

[BL89a] Ronald E. Barkley and T. Paul Lee. Point representation and hashing of an interval. Information Processing Letters, 30(4):201–203, February 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bermudez:1989:SCL

[BL89b] Manuel E. Bermudez and George Logothetis. Simple computation of LALR(1) lookahead sets. Information Processing Letters, 31(5):233– 238, June 12, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blikle:1988:GTM

[Bli88] Andrzej Blikle. A guided tour of the mathematics of MetaSoft ’88. Information Processing Letters, 29(2): 81–86, September 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blum:1983:NPC

[Blu83] Norbert Blum. A note on the “parallel computation thesis”. Information Processing Letters, 17(4):203– 205, November 8, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blum:1986:TAA

[Blu86] Lenore Blum. Towards an asymptotic analysis of Karmarkar’s algorithm. Information Processing Letters, 23(4):189–194, November 8, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertoni:1981:ECC

[BM81] Alberto Bertoni and Giancarlo Mauri. On efficient computation of the coefficients of some polynomials with applications to some enumeration problems. Information Processing Letters, 12(3):142–145, June 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bergstra:1982:STL

[BM82] J. A. Bergstra and J.-J. Ch. Meyer.

A simple transfer lemma for algebraic specifications. Information Processing Letters, 14(2):80–85, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boreddy:1989:AFP

[BM89a] Jayaramaiah Boreddy and R. N.

Mukherjee. An algorithm to find polygon similarity. Information Processing Letters, 33(4):205–206, December 21, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boxer:1989:PCM

[BM89b] Laurence Boxer and Russ Miller. A parallel circle-cover minimization algorithm. Information Processing Letters, 32(2):57–60, July 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burton:1988:PAP

[BMRS88] F. Warren Burton, G. P. McKeown, and V. J. Rayward-Smith. On process assignment in parallel computing. Information Processing Letters, 29(1):31–34, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bohm:1980:CBE

[BMS80] Corrado Böhm, Antonio Machì, and Giovanna Sontacchi. Complexity bounds for equivalence and isomorphism of Latin squares. Information Processing Letters, 10(4–5):231–233, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Book:1985:CRE

[BO85] Ronald V. Book and Friedrich Otto. Cancellation rules and extended word problems. Information Processing Letters, 20(1):5–11, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baumgarten:1986:TPC

[BO86] Bernd Baumgarten and Peter Ochsenschläger. On termination and phase changes in the presence of unreliable communication. Information Processing Letters, 22(1):15–20, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bodlaender:1988:BLB

[Bod88a] Hans L. Bodlaender. A better lower bound for distributed leader finding in bidirectional, asynchronous rings of processors. Information Processing Letters, 27(6):287–290, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bodlaender:1988:CFU

[Bod88b] Hans L. Bodlaender. The complexity of finding uniform emulations on fixed graphs. Information Processing Letters, 29(3):137–141, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bodlaender:1989:ANN

[Bod89] Hans L. Bodlaender. Achromatic number is NP-complete for cographs and interval graphs. Information Processing Letters, 31(3):135–138, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Booth:1980:LLC

[Boo80] Kellogg S. Booth. Lexicographically least circular substrings. Information Processing Letters, 10(4– 5):240–242, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Book:1981:NGC

[Boo81a] Ronald V. Book. NTS grammars and Church–Rosser systems. Information Processing Letters, 13(2):73– 76, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Book:1981:UWP

[Boo81b] Ronald V. Book. The undecidability of a word problem: on a conjecture of Strong, Maggiolo-Schettini and Rosen. Information Processing Letters, 12(3):121–122, June 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boom:1988:LVR

[Boo88] H. J. Boom. Lazy variablerenumbering makes substitution cheap. Information Processing Letters, 29(5):229–232, November 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Boppana:1989:ACP

[Bop89] Ravi B. Boppana. The average-case parallel complexity of sorting. Information Processing Letters, 33(3): 145–146, November 30, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bourret:1984:HES

[Bou84] Paul Bourret. How to estimate the sizes of domains. Information Processing Letters, 19(5):237– 243, November 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bounas:1985:DDS

[Bou85] Adam C. Bounas. Direct determination of a “seed” binary matrix. Information Processing Letters, 20(1): 47–50, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bartholdi:1983:FHB

[BP83] John J. Bartholdi, III and Loren K.

Platzman. A fast heuristic based on space filling curves for minimumweight matching in the plane. Information Processing Letters, 17(4): 177–180, November 8, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Beauquier:1985:CAI

[BP85a] D. Beauquier and D. Perrin. Codeterministic automata on infinite words. Information Processing Letters, 20(2):95–98, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bilardi:1985:VOA

[BP85b] G. Bilardi and F. P. Preparata. The VLSI optimality of the AKS sorting network. Information Processing Letters, 20(2):55–59, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bini:1985:FPP

[BP85c] Dario Bini and Victor Ya. Pan. Fast parallel polynomial division via reduction to triangular Toeplitz matrix inversion and to polynomial inversion modulo a power. Information Processing Letters, 21(2):79–81, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bini:1987:LBT

[BP87] Dario Bini and Victor Ya. Pan. A logarithmic Boolean time algorithm for parallel polynomial division. Information Processing Letters, 24(4): 233–237, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bugrara:1988:ELB

[BP88] Khaled M. Bugrara and Paul W. Purdom. An exponential lower bound for the pure literal rule. Information Processing Letters, 27(4): 215–219, April 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bern:1989:SPE

[BP89] Marshall Bern and Paul Plassmann. The Steiner problem with edge lengths 1 and 2. Information Processing Letters, 32(4):171– 176, September 1, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bloniarz:1989:LBD

[BR89] Peter A. Bloniarz and S. S. Ravi. An Ω(n log n) lower bound for decomposing a set of points into chains. Information Processing Letters, 31(6): 319–322, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brandwajn:1988:LID

[Bra88] Alexandre Brandwajn. Load imbalance in DASD dynamic reconnection. Information Processing Letters, 28(3):111–119, July 4, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brent:1980:ABT

[Bre80] R. P. Brent. On the area of binary tree layouts. Information Processing Letters, 11(1):46–48, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brown:1980:IBL

[Bro80a] Donna J. Brown. An improved BL lower bound. Information Processing Letters, 11(1):37–39, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Broy:1980:TSC

[Bro80b] M. Broy. Transformational semantics for concurrent programs. Information Processing Letters, 11(2): 87–91, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Broy:1983:DSC

[Bro83] Manfred Broy. Denotational semantics of communicating processes based on a language for applicative multiprogramming. Information Processing Letters, 17(1):29–35, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Broy:1986:DSC

[Bro86] Manfred Broy. Denotational semantics of communicating sequential programs. Information Processing Letters, 23(5):253–259, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Broy:1987:PSF

[Bro87] Manfred Broy. Predicative specifications for functional programs describing communicating networks. Information Processing Letters, 25 (2):93–101, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bruynooghe:1981:SCS

[Bru81] M. Bruynooghe. Solving combinatorial search problems by intelligent backtracking. Information Processing Letters, 12(1):36–39, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Benois:1986:CSE

[BS86a] Michèle Benois and Jacques Sakarovitch. On the complexity of some extended word problems defined by cancellation rules. Information Processing Letters, 23(6):281–287, December 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bourret:1986:LUB

[BS86b] Paul Bourret and R. Souza de Oliveira. Lower and upper bounds of the sizes of domains: estimates and experiments. Information Processing Letters, 22(5):247–253, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bajantri:1987:FMM

[BS87] M. Bajantri and David B. Skillicorn. A fast multiprocessor message passing implementation. Information Processing Letters, 24(6): 381–389, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bertolazzi:1989:DSV

[BS89] Paola Bertolazzi and Antonio Sassano. A decomposition strategy for the vertex cover problem. Information Processing Letters, 31(6): 299–304, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Balakrishnan:1988:SAM

[BSM^{+}88] M. Balakrishnan, S. Sutarwala, A. K. Majumdar, D. K. Banerji, J. G. Linders, and J. C. Majithia. A semantic approach for modular synthesis of VLSI systems. Information Processing Letters, 27(1):1– 7, February 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Berghammer:1989:SQD

[BSZ89] Rudolf Berghammer, Günther Schmidt, and Hans Zierer. Symmetric quotients and domain constructions. Information Processing Letters, 33(3):163–168, November 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bergstra:1982:TTA

[BT82] J. A. Bergstra and J. V. Tucker.

Two theorems about the completeness of Hoare’s logic. Information Processing Letters, 15(4):143– 149, October 31, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baleanu:1987:ASP

[BT87] C. Baleanu and D. Tomescu. An architecture for symbolic processing. Information Processing Letters, 26(4):217–222, December 4, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Book:1989:NSS

[BT89] Ronald V. Book and Shou Wen Tang. A note on sparse sets and the polynomial-time hierarchy. Information Processing Letters, 33(3): 141–143, November 30, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baron:1989:FLQ

[BU89] Gerd Baron and Friedrich Urbanek. Factorial languages with quadratically upper bounded growth functions and nonlinearly upper bounded subword complexities. Information Processing Letters, 32(5): 267–269, September 22, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buchta:1989:ANM

[Buc89] Christian Buchta. On the average number of maxima in a set of vectors. Information Processing Letters, 33(2):63–65, November 10, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Budd:1985:CRR

[Bud85] Timothy A. Budd. Creation and reflexive rights in grammatical protection systems. Information Processing Letters, 21(3):141–145, September 5, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Buda:1987:MA

[Bud87] Anatoliĭ O. Buda. Multiprocessor automata. Information Processing Letters, 25(4):257–261, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bui:1977:SMS

[Bui77] T. D. Bui. On an L-stable method for stiff differential equations. Information Processing Letters, 6(5): 158–161, October ??, 1977. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See erratum [Bui79] and comments [Day80a].

Bui:1979:ESM

[Bui79] T. D. Bui. Erratum: “On an Lstable method for stiff differential equations”. Information Processing Letters, 8(4):218, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Bui77, Day80a].

Bui:1980:NIB

[Bui80] T. D. Bui. A note on an improved bisection algorithm. Information Processing Letters, 10(1):35–36, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burton:1982:EFI

[Bur82a] F. Warren Burton. An efficient functional implementation of FIFO queues. Information Processing Letters, 14(5):205–206, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burton:1982:LST

[Bur82b] F. Warren Burton. A linear space translation of functional programs to Turner combinators. Information Processing Letters, 14(5):201– 204, July 23, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Burton:1989:NHF

[Bur89] F. Warren Burton. A note on higherorder functions versus logical variables. Information Processing Letters, 31(2):91–95, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brouwer:1982:NMK

[Bv82] Andries E. Brouwer and Peter van Emde Boas. A note on: “Master keys for group sharing” [Inform. Process. Lett. 12 (1981), no. 1, 23– 25; MR 82d:94046] by D. E. Denning and F. B. Schneider. Information Processing Letters, 14(1):12–14, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [DS81].

Bezem:1988:ECL

[BV88] Marc Bezem and Jan Van Leeuwen. On estimating the complexity of logarithmic decompositions. Information Processing Letters, 26(6):321– 324, January 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Broy:1980:PDE

[BW80] M. Broy and M. Wirsing. Program development: from enumeration to backtracking. Information Processing Letters, 10(4–5):193–197, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Brustmann:1987:CSF

[BW87] Bettina Brustmann and Ingo Wegener. The complexity of symmetric functions in bounded-depth circuits. Information Processing Letters, 25(4):217–219, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Blazewicz:1984:SIP

[BWD84] J. Blazewicz, J. Weglarz, and M. Drabowski. Scheduling independent 2-processor tasks to minimize schedule length. Information Processing Letters, 18(5):267–273, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Baeza-Yates:1987:SAM

[BY87] Ricardo A. Baeza-Yates. Some average measures in m-ary search trees. Information Processing Letters, 25 (6):375–382, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Bykat:1978:CHF

[Byk78] A. Bykat. Convex hull of a finite set of points in two dimensions. Information Processing Letters, 7(6): 296–298, October ??, 1978. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [DS79, Fou79, Ov80].

Bar-Yehuda:1982:CFP

[BYV82] Reuven Bar-Yehuda and Uzi Vishkin. Complexity of finding k-path-free dominating sets in graphs. Information Processing Letters, 14(5): 228–232, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cooper:1986:ESB

[CA86] Jack Cooper and Selim G. Akl. Efficient selection on a binary tree. Information Processing Letters, 23(3): 123–126, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Calude:1987:SEN

[Cal87] Cristian Calude. Super-exponentials nonprimitive recursive, but rudimentary. Information Processing Letters, 25(5):311–315, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carpi:1983:SSF

[Car83] Arturo Carpi. On the size of a square-free morphism on a three letter alphabet. Information Processing Letters, 16(5):231–235, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carpi:1984:CSW

[Car84] Arturo Carpi. On the centers of the set of weakly square-free words on a two letter alphabet. Information Processing Letters, 19(4):187– 190, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carlsson:1986:SFS

[Car86] Svante Carlsson. Splitmerge: a fast stable merging algorithm. Information Processing Letters, 22(4): 189–192, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carlsson:1987:DDH

[Car87a] Svante Carlsson. The Deap — a double-ended heap to implement double-ended priority queues. Information Processing Letters, 26(1):33– 36, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carlsson:1987:VHA

[Car87b] Svante Carlsson. A variant of Heapsort with almost optimal number of comparisons. Information Processing Letters, 24(4):247–250, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Casanova:1983:TFS

[Cas83] Marco A. Casanova. The theory of functional and subset dependencies over relational expressions. Information Processing Letters, 16(3): 153–160, April 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choffrut:1983:FPD

[CC83] C. Choffrut and K. Culik, II. Folding of the plane and the design of systolic arrays. Information Processing Letters, 17(3):149–153, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:1984:PTM

[CC84] M. Chrobak and B. S. Chlebus. Probabilistic Turing machines and recursively enumerable Dedekind cuts. Information Processing Letters, 19(4):167–171, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choe:1986:ECL

[CC86] Kwang-Moo M. Choe and Chun-Hyon H. Chang. Efficient computation of the locally least-cost insertion string for the LR error repair. Information Processing Letters, 23 (6):311–316, December 3, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1987:HSS

[CC87] Tae-Choong Chung and Jung-Wan Cho. History sensitive string for multiple alphabets. Information Processing Letters, 25(3):183–188, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:1988:OMP

[CC88a] M. Y. Chan and W. L. Chung. Optimal multidisk partial match file designs. Information Processing Letters, 28(3):149–155, July 4, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1988:OMP

[CC88b] C. C. Chang and C. H. Chang. An ordered minimal perfect hashing scheme with single parameter. Information Processing Letters, 27(2): 79–83, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chetto:1989:SPS

[CC89] Houssine Chetto and Maryline Chetto. Scheduling periodic and sporadic tasks in a real-time system. Information Processing Letters, 30(4):177–184, February 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chern:1989:LBB

[CCL89a] Maw-Sheng S. Chern, G. H. Chen, and Pangfeng Liu. An LC branchand-bound algorithm for the module assignment problem. Information Processing Letters, 32(2):61–71, July 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1989:FCP

[CCL89b] Kuo Liang Chung, Wen Chin Chen, and Ferng-Ching Lin. Fast computation of periodic continued fractions. Information Processing Letters, 33 (2):67–72, November 10, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carlsson:1989:NCD

[CCS89] Svante Carlsson, Jingsen Chen, and Thomas Strothotte. A note on the construction of the data structure “deap”. Information Processing Letters, 31(6):315–317, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carpi:1986:SWP

[CD86] Arturo Carpi and Aldo De Luca.

Square-free words of partially commutative free monoids. Information Processing Letters, 22(3):125– 131, March 3, 1986. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Consel:1989:PEP

[CD89] Charles Consel and Olivier Danvy. Partial evaluation of pattern matching in strings. Information Processing Letters, 30(2):79–86, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Castan:1987:SCA

[CDL87] M. Castan, M.-H. Durand, and M. Lemaî tre. A set of combinators for abstraction in linear space. Information Processing Letters, 24(3): 183–188, February 13, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Casas:1989:ACA

[CDS89] R. Casas, J. Diaz, and J. M. Steyaert. Average-case analysis of Robinson’s unification algorithm with two different variables. Information Processing Letters, 31(5): 227–232, June 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheston:1989:CST

[CFHP89] Grant A. Cheston, Art Farley, S. T. Hedetniemi, and Andrzej Proskurowski. Centering a spanning tree of a biconnected graph. Information Processing Letters, 32 (5):247–250, September 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chakrabarti:1989:ISEa

[CGPD89] P. P. Chakrabarti, S. Ghose, A. Pandey, and S. C. De Sarkar. Increasing search efficiency using multiple heuristics. Information Processing Letters, 30(1):33–36, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chakrabarti:1989:ISEb

[CGPdS89] P. P. Chakrabarti, S. Ghose, A. Pandey, and S. C. de Sarkar. Increasing search efficiency using multiple heuristics. Information Processing Letters, 32(5):275, September 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1987:CS

[CGR87] Ernest J. H. Chang, Gaston H.

Gonnet, and Doron Rotem. On the costs of self-stabilization. Information Processing Letters, 24(5): 311–316, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ceri:1989:MSJ

[CGTW89] Stefano Ceri, Georg Gottlob, Letizia Tanca, and Gio Wiederhold. Magic semi-joins. Information Processing Letters, 33(2):97– 107, November 10, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cormack:1984:AAH

[CH84] Gordon V. Cormack and R. Nigel Horspool. Algorithms for adaptive Huffman codes. Information Processing Letters, 18(3):159–165, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cockayne:1986:ECS

[CH86] E. J. Cockayne and D. E. Hewgill. Exact computation of Steiner minimal trees in the plane. Information Processing Letters, 22(3):151– 156, March 3, 1986. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:1988:TUD

[CH88a] Edward P. F. Chan and Hector J. Hernandez. Testing unboundedness of database schemes and functional dependencies. Information Processing Letters, 28(6):317–326, August 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cho:1988:CHB

[CH88b] Sang Cho and Dũng T. Hùynh. On a complexity hierarchy between L and NL. Information Processing Letters, 29(4):177–182, November 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:1988:NRS

[CH88c] Marek Chrobak and Richard Harter. A note on random sampling. Information Processing Letters, 29(5): 255–256, November 24, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cull:1989:CFN

[CH89] Paul Cull and James L. Holloway. Computing Fibonacci numbers quickly. Information Processing Letters, 32(3):143–149, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:1981:DFP

[Cha81] Tat Hung Chan. Deciding freeness for program schemes with a single unary function. Information Processing Letters, 13(3):98–102, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chazelle:1983:DPO

[Cha83a] Bernard Chazelle. A decision procedure for optimal polyhedron partitioning. Information Processing Letters, 16(2):75–78, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chazelle:1983:IAF

[Cha83b] Bernard Chazelle. An improved algorithm for the fixed-radius neighbor problem. Information Processing Letters, 16(4):193–198, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:1985:NRD

[Cha85] Mee Yee Chan. A note on redundant Disk Modulo allocation. Information Processing Letters, 20(3): 121–123, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chatelin:1986:TAM

[Cha86] Philippe Chatelin. On transformations of algorithms to multiply 2 × 2 matrices. Information Processing Letters, 22(1):1–5, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chan:1988:BPT

[Cha88] Tat Hung Chan. The boundedness problem for three-dimensional vector addition systems with states. Information Processing Letters, 26(6): 287–289, January 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cheung:1982:SME

[Che82] To-Yat Cheung. A statistical model for estimating the number of records in a relational database. Information Processing Letters, 15(3):115– 118, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1986:SCD

[Che86] Lin Chen. O(1) space complexity deletion for AVL trees. Information Processing Letters, 22(3):147– 149, March 3, 1986. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chlebus:1988:PBS

[Chl88] Bogdan S. Chlebus. A parallel bucket sort. Information Processing Letters, 27(2):57–61, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chlebus:1989:PIB

[Chl89] Bogdan S. Chlebus. Parallel iterated bucket sort. Information Processing Letters, 31(4):181–183, May 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choffrut:1981:CPD

[Cho81] C. Choffrut. A closure property of deterministic context-free languages. Information Processing Letters, 12(1):13–16, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chor:1982:AFF

[Cho82] Ben-Zion Chor. Arithmetic of finite fields. Information Processing Letters, 14(1):4–6, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:1984:NBR

[Chr84] Marek Chrobak. A note on bounded-reversal multipushdown machines. Information Processing Letters, 19(4):179–180, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cialdea:1986:SRP

[Cia86] Marta Cialdea. Some remarks on the possibility of extending resolution proof procedures to intuitionistic logic. Information Processing Letters, 22(2):87–90, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cidon:1988:YAD

[Cid88] Isreal Cidon. Yet another distributed depth-first-search algorithm. Information Processing Letters, 26(6):301–305, January 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chang:1988:ESO

[CIR88] Jik H. Chang, Oscar H. Ibarra, and Bala Ravikumar. Erratum: “Some observations concerning alternating Turing machines using small space” [Inform. Process. Lett. 25 (1987), no. 1, 1–9; MR 88e:68026] by the authors and L. Berman. Information Processing Letters, 27(1):53, February 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [CIRB87].

Chang:1987:SOC

[CIRB87] Jik H. Chang, Oscar H. Ibarra, Bala Ravikumar, and Leonard Berman. Some observations concerning alternating Turing machines using small space. Information Processing Letters, 25(1):1–9, April 20, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [CIR88].

Cioni:1984:PDD

[CK84] Gianna Cioni and Antoni Kreczmar. Programmed deallocation without dangling reference. Information Processing Letters, 18(4):179–187, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choi:1986:CQP

[CK86a] Key-Sun S. Choi and Gil Chang Kim. A controlled quantification in parsing of Montague grammar. Information Processing Letters, 22(4): 207–216, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culik:1986:NEP

[CK86b] Karel Culik, II and Juhani Karhumäki.

A note on the equivalence problem of rational formal power series. Information Processing Letters, 23 (1):29–31, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1988:APR

[CK88a] Moon Jung Chung and M. S. Krishnamoorthy. Algorithms of placing recovery points. Information Processing Letters, 28(4):177–181, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culik:1988:TSN

[CK88b] Karel Culik, II and Juhani Karhumäki.

On totalistic systolic networks. Information Processing Letters, 26(5): 231–236, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Choi:1989:HJI

[CK89] Hwang Kyu Choi and Myunghwan Kim. Hybrid join: an improved sort-based join algorithm. Information Processing Letters, 32(2):51–56, July 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cybenko:1987:FHE

[CKV87] George Cybenko, David W. Krumme, and K. N. Venkataraman. Fixed hypercube embedding. Information Processing Letters, 25(1):35–39, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Calmet:1980:IRP

[CL80] J. Calmet and R. Loos. An improvement of Rabin’s probabilistic algorithm for generating irreducible polynomials over GF(p). Information Processing Letters, 11(2):94–95, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clerbout:1981:IDM

[CL81] Mireille Clerbout and Michel Latteux. The inclusion of D0L in MULTI-RESET. Information Processing Letters, 13(2):45–47, November 13, 1981. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Codenotti:1986:NVC

[CL86a] B. Codenotti and G. Lotti. A note on the VLSI counter. Information Processing Letters, 22(4):193–195, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Codenotti:1986:ATB

[CL86b] Bruno Codenotti and Grazia Lotti. Area-time tradeoffs for bilinear forms computations in VLSI. Information Processing Letters, 23(2): 107–109, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Codenotti:1986:VFS

[CL86c] Bruno Codenotti and Grazia Lotti. A VLSI fast solver for tridiagonal linear systems. Information Processing Letters, 23(3):111–114, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1988:PSI

[CL88a] Guan-Ing Chen and Ten Hwang Lai.

Preemptive scheduling of independent jobs on a hypercube. Information Processing Letters, 28(4): 201–206, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cook:1988:SPA

[CL88b] Stephen A. Cook and Michael Luby.

A simple parallel algorithm for finding a satisfying truth assignment to a 2-CNF formula. Information Processing Letters, 27(3):141–145, March 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clarkson:1983:MGA

[Cla83] Kenneth L. Clarkson. A modification of the greedy algorithm for vertex cover. Information Processing Letters, 16(1):23–25, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clarkson:1986:LPT

[Cla86] Kenneth L. Clarkson. Linear programming in O(n × 3^{d^{2}}) time. Information Processing Letters, 22(1): 21–24, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clouatre:1989:GIR

[CLM89] A. Clouatre, N. Laliberte, and T. H.

Merrett. A general implementation of relational recursion with speedup techniques for programmers. Information Processing Letters, 32(5): 257–262, September 22, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Clocksin:1983:RFQ

[Clo83] W. F. Clocksin. Real-time functional queue operations using the logical variable. Information Processing Letters, 17(4):173–175, November 8, 1983. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1980:MNS

[CLW80a] K. M. Chung, F. Luccio, and C. K. Wong. Minimum number of steps for permutation in a bubble memory. Information Processing Letters, 11(2):81–83, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chung:1980:NPA

[CLW80b] K. M. Chung, F. Luccio, and C. K. Wong. A new permutation algorithm for bubble memories. Information Processing Letters, 10(4–5): 226–230, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1983:OTF

[CLW83] Max Crochemore, Michel Le Rest, and Philippe Wender. An optimal test on finite unavoidable sets of words. Information Processing Letters, 16(4):179–180, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cellary:1980:SMQ

[CM80a] Wojciech Cellary and Daniel Mayer.

A simple model of query scheduling in distributed data base systems. Information Processing Letters, 10(3): 137–147, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colbourn:1980:CCP

[CM80b] Charles J. Colbourn and Brendan D.

McKay. A correction to Colbourn’s paper on the complexity of matrix symmetrizability: “The complexity of symmetrizing matrices” [Inform. Process. Lett. 9 (1979), no. 3, 108– 109; MR 81a:68045]. Information Processing Letters, 11(2):96–97, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Col79].

Cheriyan:1989:PCF

[CM89] J. Cheriyan and S. N. Maheshwari.

The parallel complexity of finding a blocking flow in a 3-layer network. Information Processing Letters, 31 (3):157–161, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ciaccia:1988:UAE

[CMT88] Paolo Ciaccia, Dario Maio, and Paolo Tiberio. A unifying approach to evaluating block accesses in database organizations. Information Processing Letters, 28(5):253– 257, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chin:1988:OWR

[CN88] Wei-Pang Chin and Simeon Ntafos. Optimum watchman routes. Information Processing Letters, 28(1):39– 44, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cohen:1983:NFA

[Coh83] Jacques Cohen. A note on a fast algorithm for sparse matrix multiplication. Information Processing Letters, 16(5):247–248, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Colbourn:1979:CSM

[Col79] Charles J. Colbourn. The complexity of symmetrizing matrices. Information Processing Letters, 9(3): 108–109, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also correction [CM80b].

Cole:1988:OES

[Col88] Richard John Cole. An optimally efficient selection algorithm. Information Processing Letters, 26(6):295– 299, January 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Constable:1983:PPS

[Con83] Robert L. Constable. Programs as proofs: a synopsis. Information Processing Letters, 16(3):105–112, April 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cook:1988:SPF

[Coo88] Stephen A. Cook. Short propositional formulas represent nondeterministic computations. Information Processing Letters, 26(5):269– 270, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Courcelle:1981:SAT

[Cou81] Bruno Courcelle. The simultaneous accessibility of two configurations of two equivalent DPDAs. Information Processing Letters, 12(3): 111–114, June 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Courcelle:1984:SNR

[Cou84] Bruno Courcelle. Some negative results concerning DPDAs. Information Processing Letters, 18(5):285– 289, June 18, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ceri:1981:UBN

[CP81] S. Ceri and G. Pelagatti. An upper bound on the number of execution nodes for a distributed join. Information Processing Letters, 12(1): 46–48, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ceri:1982:SMN

[CP82] S. Ceri and G. Pelagatti. A solution method for the non-additive resource allocation problem in distributed system design. Information Processing Letters, 15(4):174– 178, October 31, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chrobak:1987:RSM

[CR87] Marek Chrobak and Wojciech Rytter. Remarks on string-matching and one-way multihead automata. Information Processing Letters, 24 (5):325–329, March 16, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chandran:1988:OSH

[CR88] Sharat Chandran and Azriel Rosenfeld. Order statistics on a hypercube. Information Processing Letters, 27(3):129–132, March 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culberson:1989:FAC

[CR89] Joseph C. Culberson and Piotr Rudnicki. A fast algorithm for constructing trees from distance matrices. Information Processing Letters, 30(4):215–220, February 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Codenotti:1985:VIF

[CRL85] B. Codenotti, F. Romani, and G. Lotti. VLSI implementation of fast solvers for band linear systems with constant coefficient matrix. Information Processing Letters, 21(3): 159–163, September 5, 1985. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Crochemore:1981:OAC

[Cro81] Max Crochemore. An optimal algorithm for computing the repetitions in a word. Information Processing Letters, 12(5):244–250, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carlson:1982:ETS

[CS82] David A. Carlson and John E. Savage. Extreme time-space tradeoffs for graphs with small space requirements. Information Processing Letters, 14(5):223–227, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cesarini:1983:ACC

[CS83a] F. Cesarini and G. Soda. An algorithm to construct a compact B-tree in case of ordered keys. Information Processing Letters, 17(1):13–16, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Croitoru:1983:PSG

[CS83b] Cornelius Croitoru and Emilian Suditu. Perfect stables in graphs. Information Processing Letters, 17 (1):53–56, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chng:1989:SSH

[CSO89] Hock Thiam Ch’ng, B. Srinivasan, and Beng Chin Ooi. Study of selforganizing heuristics for skewed access patterns. Information Processing Letters, 30(5):237–244, March 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culik:1989:VFS

[Cul89] Karel Culik, II. Variations of the firing squad problem and applications. Information Processing Letters, 30(3):153–157, February 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Carson:1989:EBD

[CV89] Scott D. Carson and Paul Vongsathorn. Error bounds on disk arrangement using frequency information. Information Processing Letters, 31 (4):209–213, May 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Culik:1982:NST

[CW82] Karel Culik, II and Derick Wood. A note on some tree similarity measures. Information Processing Letters, 15(1):39–42, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chin:1984:MVD

[CW84] Francis Chin and Cao An Wang. Minimum vertex distance between separable convex polygons. Information Processing Letters, 18(1):41– 45, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1986:VAF

[CW86] G. Chen and M. H. Williams. The value of an array facility in Prolog. Information Processing Letters, 23(5):247–251, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Cole:1985:PMA

[CY85] Richard Cole and Chee K. Yap. A parallel median algorithm. Information Processing Letters, 20(3):

137–139, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1989:UFP

[CYC89] Y. C. Chen, Z. C. Yeh, and G. H.

Chen. Using fewer processors to reduce time complexities of semigroup computations. Information Processing Letters, 32(2):89–93, July 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Chen:1988:TAC

[CYL88] Gen-Huey Chen, M. S. Yu, and L. T.

Liu. Two algorithms for constructing a binary tree from its traversals. Information Processing Letters, 28 (6):297–299, August 29, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czaja:1980:DFP

[Cza80a] Ludwik Czaja. Deadlock and fairness in parallel schemas: a settheoretic characterization and decision problems. Information Processing Letters, 10(4–5):234–239, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czaja:1980:PSS

[Cza80b] Ludwik Czaja. Parallel system schemas and their relation to automata. Information Processing Letters, 10(3):153–158, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czaja:1988:CS

[Cza88] Ludwik Czaja. Cause-effect structures. Information Processing Letters, 26(6):313–319, January 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Czaja:1989:FPC

[Cza89] Ludwik Czaja. Finite processes in cause-effect structures and their composition. Information Processing Letters, 31(6):305–310, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dai:1986:ELM

[Dai86] Ke-Chang Dai. EDISON-80, a language for modular programming of parallel processes. Information Processing Letters, 22(2):61–72, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Damaschke:1989:HCP

[Dam89] Peter Damaschke. The Hamiltonian circuit problem for circle graphs is NP-complete. Information Processing Letters, 32(1):1–2, July 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Darondeau:1989:BE

[Dar89] Ph. Darondeau. Bisimulation and effectiveness. Information Processing Letters, 30(1):19–20, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Day:1980:CSM

[Day80a] J. D. Day. Comments on: “On an L-stable method for stiff differential equations” [Inform. Process. Lett. 6 (1977), no. 5, 158–161; MR 56 #10015] by T. D. Bui. Information Processing Letters, 11(1):31–32, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Bui77, Bui79].

Day:1980:ISR

[Day80b] J. D. Day. On the internal Sstability of Rosenbrock methods. Information Processing Letters, 11(1): 27–30, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dudzinski:1981:SMS

[DD81] Krzysztof Dudziński and Andrzej Dydek. On a stable minimum storage merging algorithm. Information Processing Letters, 12(1):5–8, February 13, 1981. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deLuca:1981:CPF

[de 81] Aldo de Luca. A combinatorial property of the Fibonacci words. Information Processing Letters, 12(4): 193–195, ???? ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeFelice:1982:TC

[De 82] Clelia De Felice. On the triangle conjecture. Information Processing Letters, 14(5):197–200, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deLemosMeira:1987:SCF

[de 87] Silvio Romero de Lemos Meira.

Strict combinators. Information Processing Letters, 24(4):255–258, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deb:1980:CFA

[Deb80] Ashoke Deb. Conflict-free access of arrays — a counter example. Information Processing Letters, 10(1): 20, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dowdy:1984:TCR

[DEGS84] Lawrence W. Dowdy, Derek L. Eager, Karen D. Gordon, and Lawrence V. Saxton. Throughput concavity and response time convexity. Information Processing Letters, 19(4):209–212, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dehne:1986:AME

[Deh86a] Frank Dehne. O(n^{1/2}) algorithms for the maximal elements and ECDF searching problem on a meshconnected parallel computer. Information Processing Letters, 22(6): 303–306, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dehornoy:1986:TCO

[Deh86b] Patrick Dehornoy. Turing complexity of the ordinals. Information Processing Letters, 23(4):167– 170, November 8, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Deleage:1983:ATL

[Del83] J. L. Deléage. An application of a transfer lemma. Information Processing Letters, 16(4):161–163, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demolombe:1982:GDR

[Dem82] R. Demolombe. Generalized division for relational algebraic language. Information Processing Letters, 14(4): 174–178, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dershowitz:1980:SMA

[Der80] N. Dershowitz. The Schorr–Waite marking algorithm revisited. Information Processing Letters, 11(3): 141–143, November 18, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Derigs:1985:EDL

[Der85] Ulrich Derigs. An efficient Dijkstralike labeling method for computing shortest odd/even paths. Information Processing Letters, 21(5):253– 258, November 18, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devroye:1980:NFC

[Dev80] Luc Devroye. A note on finding convex hulls via maximal vectors. Information Processing Letters, 11(1): 53–56, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devroye:1985:NET

[Dev85] Luc Devroye. A note on the expected time required to construct the outer layer. Information Processing Letters, 20(5):255–257, June 12, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dyer:1984:PAM

[DF84] M. E. Dyer and A. M. Frieze. A partitioning algorithm for minimum weighted Euclidean matching. Information Processing Letters, 18(2): 59–62, February 28, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Das:1988:FFI

[DF88] P. K. Das and D. Q. M. Fay. Faulttolerant and flexible interconnection of multiple processors. Information Processing Letters, 28(5):259– 268, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dijkstra:1983:DTD

[DFv83] Edsger W. Dijkstra, W. H. J. Feijen, and A. J. M. van Gasteren. Derivation of a termination detection algorithm for distributed computations. Information Processing Letters, 16 (5):217–219, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deFraysseix:1989:NOD

[dI89a] Hubert de Fraysseix and Hiroshi Imai. Notes on oriented depth-first search and longest paths. Information Processing Letters, 31(1):53–56, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Demolombe:1989:HSO

[DI89b] Robert Demolombe and Arantza Illarramendi. Heuristics for syntactical optimization of relational queries. Information Processing Letters, 32(6):313–316, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dietzfelbinger:1989:SCO

[Die89] Martin Dietzfelbinger. The speed of copying on one-tape off-line Turing machines. Information Processing Letters, 33(2):83–89, November 10, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dillencourt:1987:NHN

[Dil87a] Michael B. Dillencourt. A non-Hamiltonian, nondegenerate Delaunay triangulation. Information Processing Letters, 25(3):149–151, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dillencourt:1987:TSC

[Dil87b] Michael B. Dillencourt. Traveling salesman cycles are not always subgraphs of Delaunay triangulations or of minimum weight triangulations. Information Processing Letters, 24(5):339–342, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Drysdale:1989:NLB

[DJ89] Robert L. Drysdale, III and Jerzy W. Jaromczyk. A note on lower bounds for the maximum area and maximum perimeter k-gon problems. Information Processing Letters, 32(6):301–303, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Darondeau:1984:TFP

[DK84] Ph. Darondeau and L. Kott. Towards a formal proof system for ω-rational expressions. Information Processing Letters, 19(4):173– 177, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeVel:1988:IPA

[DK88] O. Y. De Vel and E. V. Krishnamurthy. An iterative pipelined array architecture for the generalized matrix inversion. Information Processing Letters, 26(5):263–267, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Du:1988:STS

[DL88] Jianzhong Du and Joseph Y.-T. T. Leung. Scheduling tree-structured tasks with restricted execution times. Information Processing Letters, 28(4):183–188, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Driscoll:1987:MBI

[DLF87] James R. Driscoll, Sheau-Dong Lang, and LeRoy A. Franklin. Modeling B-tree insertion activity. Information Processing Letters, 26(1): 5–18, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

dAtri:1988:HAG

[dM88] Alessandro d’Atri and Marina Moscarini. On hypergraph acyclicity and graph chordality. Information Processing Letters, 29(5):271– 274, November 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dasgupta:1989:AAM

[DM89] Bhaskar Dasgupta and C. E. Veni Madhavan. An approximate algorithm for the minimal vertex nested polygon problem. Information Processing Letters, 33(1):35–44, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Denning:1981:MMK

[DMS81] D. E. Denning, H. Meijer, and F. B.

Schneider. More on master keys for group sharing. Information Processing Letters, 13(3):125–126, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobosiewicz:1980:EVB

[Dob80] Włodzimierz Dobosiewicz. An efficient variation of bubble sort. Information Processing Letters, 11(1): 5–6, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doberkat:1982:AEH

[Dob82] Ernst-E. E. Doberkat. Asymptotic estimates for the higher moments of the expected behavior of straight insertion sort. Information Processing Letters, 14(4):179–182, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dobosiewicz:1985:NNS

[Dob85] Włodzimierz Dobosiewicz. A note on natural selection. Information Processing Letters, 21(5):239–243, November 18, 1985. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeBra:1983:AHD

[DP83] Paul De Bra and Jan Paredaens. An algorithm for horizontal decompositions. Information Processing Letters, 17(2):91–95, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deLuca:1988:TSL

[dPV88] Aldo de Luca, Mariacristina Pelagalli, and Stefano Varricchio. Test sets for languages of infinite words. Information Processing Letters, 29 (2):91–95, September 30, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Davis:1981:DCL

[DRT81] R. H. Davis, C. Rinaldi, and C. J. Trebilcock. Data compression in limited capacity microcomputer systems. Information Processing Letters, 13(4–5):138–141, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Devai:1979:CCH

[DS79] F. Dévai and T. Szendrényi. Comments on convex hull of a finite set of points in two dimensions. Information Processing Letters, 9(3): 141–142, October 5, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Byk78, Fou79, Ov80].

Dijkstra:1980:TDD

[DS80] Edsger W. Dijkstra and C. S. Scholten. Termination detection for diffusing computations. Information Processing Letters, 11(1):1–4, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Denning:1981:MKG

[DS81] Dorothy E. Denning and Fred B.

Schneider. Master keys for group sharing. Information Processing Letters, 12(1):23–25, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also note [Bv82].

Dehne:1988:TAE

[DS88] Frank Dehne and Ivan Stojmenović.

An O(\sqrt{n}) time algorithm for the ECDF searching problem for arbitrary dimensions on a mesh-of-processors. Information Processing Letters, 28(2):67–70, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duris:1985:TCA

[ĎSVT85] Pavol Ďuriš, Ondrej Sýkora, Imrich Vř’to, and Clark D. Thompson. Tight chip area lower bounds for discrete Fourier and Walsh-Hadamard transformations. Information Processing Letters, 21(5): 245–247, November 18, 1985. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Doran:1980:VSS

[DT80] R. W. Doran and L. K. Thomas.

Variants of the software solution to mutual exclusion. Information Processing Letters, 10(4–5):206–208, July 5, 1980. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Duboc:1985:SPC

[Dub85] Christine Duboc. Some properties of commutation in free partially commutative monoids. Information Processing Letters, 20(1):1–4, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dublish:1988:AFM

[Dub88] Pratul Dublish. An O(n^{3}) algorithm for finding the minimal opaque forest of a convex polygon. Information Processing Letters, 29(5):275– 276, November 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Duda:1983:ECP

[Dud83] Andrzej Duda. The effects of checkpointing on program execution time. Information Processing Letters, 16 (5):221–229, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dunne:1990:CKP

[Dun90] P. E. Dunne. Comment on Kochol’s paper “Efficient monotone circuits for threshold functions” [Inform. Process. Lett. 32(3), 24 August 1989, pp. 121–122]. Information Processing Letters, 34(5):221– 222, May 7, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Koc89].

Dijkstra:1982:ITA

[Dv82] Edsger W. Dijkstra and A. J. M. van Gasteren. An introduction to three algorithms for sorting in situ. Information Processing Letters, 15(3): 129–134, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

DeLuca:1988:FTM

[dV88] Aldo de Luca and Stefano Varricchio. On the factors of the Thue– Morse word on three symbols. Information Processing Letters, 27(6): 281–285, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

deLuca:1989:FLW

[dV89] Aldo de Luca and Stefano Varricchio. Factorial languages whose growth function is quadratically upper bounded. Information Processing Letters, 30(6):283–288, March 28, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dwelly:1988:SIB

[Dwe88] Andrew Dwelly. Synchronizing the I/O behavior of functional programs with feedback. Information Processing Letters, 28(1):45–51, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dyer:1980:FPA

[Dye80] Charles R. Dyer. A fast parallel algorithm for the closest pair problem. Information Processing Letters, 11(1):49–52, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Dymond:1988:IDL

[Dym88] Patrick W. Dymond. Input-driven languages are in log n depth. Information Processing Letters, 26(5): 247–250, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eastman:1981:OBS

[Eas81] C. M. Eastman. Optimal bucket size for nearest neighbor searching in k-d trees. Information Processing Letters, 12(4):165–167, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ebert:1980:NOE

[Ebe80] Jürgen Ebert. A note on odd and even factors of undirected graphs. Information Processing Letters, 11 (2):70–72, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ebert:1981:STC

[Ebe81] Jürgen Ebert. A sensitive transitive closure algorithm. Information Processing Letters, 12(5):255–258, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ebert:1988:CET

[Ebe88] Jürgen Ebert. Computing Eulerian trails. Information Processing Letters, 28(2):93–97, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edenbrandt:1987:CGR

[Ede87] Anders Edenbrandt. Chordal graph recognition is in NC. Information Processing Letters, 24(4):239–241, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Everett:1989:ADH

[EG89] H. Everett and A. Gupta. Acyclic directed hypercubes may have exponential diameter. Information Processing Letters, 32(5):243–245, September 22, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Engelfriet:1988:PEL

[EH88a] Joost Engelfriet and Hendrik Jan Hoogeboom. Prefix and equality languages of rational functions are co-context-free. Information Processing Letters, 28(2):77–79, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Esfahanian:1988:CCE

[EH88b] Abdol Hossein Esfahanian and S. Louis Hakimi. On computing a conditional edge-connectivity of a graph. Information Processing Letters, 27(4):195–199, April 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eichholz:1983:OND

[Eic83] S. Eichholz. Optimal networks for distributing nonsequential programs. Information Processing Letters, 16(2):71–74, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eisinger:1989:NCR

[Eis89] Norbert Eisinger. A note on the completeness of resolution without self-resolution. Information Processing Letters, 31(6):323–326, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Egecioglu:1989:ADS

[EK89] Ömer Eğecioğlu and Bahman Kalantari. Approximating the diameter of a set of points in the Euclidean space. Information Processing Letters, 32(4):205–211, September 1, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Erni:1981:TTC

[EL81] W. Erni and R. Lapsien. On the time and tape complexity of weak unification. Information Processing Letters, 12(3):146–150, June 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edelsbrunner:1981:IOO

[EM81] H. Edelsbrunner and H. A. Maurer. On the intersection of orthogonal objects. Information Processing Letters, 13(4–5):177–181, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eades:1984:AGS

[EM84] Peter Eades and Brendan McKay. An algorithm for generating subsets of fixed size with a strong minimal change property. Information Processing Letters, 19(3):131–133, October 19, 1984. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Edelsbrunner:1985:FEP

[EM85] H. Edelsbrunner and H. A. Maurer.

Finding extreme points in three dimensions and solving the post-office problem in the plane. Information Processing Letters, 21(1):39–47, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Emerson:1987:UIT

[Eme87] E. Allen Emerson. Uniform inevitability is tree automaton ineffable. Information Processing Letters, 24(2):77–79, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edelsbrunner:1982:PIS

[EMK82] H. Edelsbrunner, H. A. Maurer, and D. G. Kirkpatrick. Polygonal intersection searching. Information Processing Letters, 14(2):74–79, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ellis:1989:LTA

[EMM89] John A. Ellis, Manrique Mata, and Gary MacGillivray. A linear time algorithm for longest (s, t)-paths in weighted outerplanar graphs. Information Processing Letters, 32(4): 199–204, September 1, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edelsbrunner:1982:ESR

[EO82] Herbert Edelsbrunner and Mark H.

Overmars. On the equivalence of some rectangle problems. Information Processing Letters, 14(3):124– 127, May 16, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edelsbrunner:1987:ZRR

[EO87] Herbert Edelsbrunner and Mark H. Overmars. Zooming by repeated range detection. Information Processing Letters, 24(6):413–417, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Elgot:1982:SSA

[EPS82] Calvin C. Elgot, Alan J. Perlis, and Lawrence Snyder. A syntax-free semantics for the APL operators. Information Processing Letters, 14(3): 128–131, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:1980:EIT

[ER80] A. Ehrenfeucht and G. Rozenberg. On the emptiness of the intersection of two D0S languages problem. Information Processing Letters, 10(4– 5):223–225, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:1981:SCD

[ER81] A. Ehrenfeucht and G. Rozenberg. On the subword complexity of D0L languages with a constant distribution. Information Processing Letters, 13(3):108–113, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:1983:SCL

[ER83a] A. Ehrenfeucht and G. Rozenberg.

On the subword complexity of locally catenative DOL languages. Information Processing Letters, 16(1): 7–9, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehrenfeucht:1983:SCF

[ER83b] A. Ehrenfeucht and G. Rozenberg.

On the subword complexity of mfree D0L languages. Information Processing Letters, 17(3):121–124, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [ER83a].

Er:1983:CSO

[Er83c] M. C. Er. Computing sums of orderk Fibonacci numbers in log time. Information Processing Letters, 17 (1):1–5, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Erkio:1980:IMS

[Erk80] H. Erkio. Internal merge sorting with delayed selection. Information Processing Letters, 11(3):137– 140, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ehlers:1989:URC

[Ev89] E. M. Ehlers and S. H. von Solms. Using random context structure grammars to represent chemical structures. Information Processing Letters, 30(3):159–166, February 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Evans:1986:ETS

[Eva86] John B. Evans. Experiments with trees for the storage and retrieval of future events. Information Processing Letters, 22(5):237–242, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Edelsbrunner:1986:HRS

[EW86] Herbert Edelsbrunner and Emo Welzl. Halfplanar range search in linear space and O(n^{0.695}) query time. Information Processing Letters, 23(6):289–293, December 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Evans:1986:MPJ

[EY86] David John Evans and Nadia Y. Yousif. Merging by the parallel jump searching algorithm. Information Processing Letters, 23(5):239– 246, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Evans:1987:ESB

[EY87] D. J. Evans and W. S. Yousif. Explicit solution of block tridiagonal systems of linear equations. Information Processing Letters, 24(3): 207–209, February 13, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Eastman:1982:PSN

[EZ82] C. M. Eastman and M. Zemankova. Partially specified nearest neighbor searches using k-d trees. Information Processing Letters, 15(2):53– 56, September 6, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Falkowski:1989:SPP

[Fal89] Bernd-Jürgen J. Falkowski. A selfoptimizing Prolog program and the underlying statistical model. Information Processing Letters, 31(5): 273–276, June 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Farley:1980:LTT

[Far80] Arthur M. Farley. Levelling terrain trees: a transshipment problem. Information Processing Letters, 10(4– 5):189–192, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

FarinasDelCerro:1982:SDM

[Far82] L. Fariñas Del Cerro. A simple deduction method for modal logic. Information Processing Letters, 14 (2):49–51, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Farinas:1983:STT

[Far83] Luis Fariñas. Space as time. Information Processing Letters, 17(3): 113–115, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fernandez-Baca:1988:NDP

[FB88] David Fernández-Baca. Nonserial dynamic programming formulations of satisfiability. Information Processing Letters, 27(6):323–326, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feinberg:1989:CSP

[Fei89] Ellen B. Feinberg. Characterizing the shortest path of an object among obstacles. Information Processing Letters, 31(5):257–264, June 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Foster:1989:LBA

[FG89] M. J. Foster and Ronald I. Greenberg. Lower bounds on the area of finite-state machines. Information Processing Letters, 30(1):1–7, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraser:1987:OAE

[FH87] Christopher W. Fraser and David R. Hanson. Optimization of argument evaluation order. Information Processing Letters, 24(6):391–395, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:1987:ILK

[FI87] Michael J. Fischer and Neil Immerman. Interpreting logics of knowledge in propositional dynamic logic with converse. Information Processing Letters, 25(3):175–181, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Field:1987:NND

[Fie87] Dan Field. A note on a new data structure for in-the-past queries. Information Processing Letters, 24(2): 95–96, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:1982:LBT

[FL82] Michael J. Fischer and Nancy A.

Lynch. A lower bound for the time to assure interactive consistency. Information Processing Letters, 14(4): 183–186, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Friesen:1984:SSS

[FL84] D. K. Friesen and M. A. Langston.

A storage-size selection problem. Information Processing Letters, 18(5): 295–296, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fellows:1987:NAP

[FL87] Michael R. Fellows and Michael A.

Langston. Nonconstructive advances in polynomial-time complexity. Information Processing Letters, 26(3):157–162, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fleischer:1989:CCM

[Fle89] Rudolf Fleischer. Communication complexity of multi-processor systems. Information Processing Letters, 30(2):57–65, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Floren:1991:NFA

[Flo91] Rolf Floren. A note on: “A faster approximation algorithm for the Steiner problem in graphs” [Inform. Process. Lett. 27 (1988), no. 3, 125–128; MR 89d:68031] by K. Mehlhorn. Information Processing Letters, 38(4):177–178, May 31, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Meh88].

Finkel:1987:EDA

[FM87] Raphael Finkel and Hari H. Madduri. An efficient deadlock avoidance algorithm. Information Processing Letters, 24(1):25–30, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fries:1987:DST

[FMNT87] O. Fries, K. Mehlhorn, S. Näher, and A. Tsakalidis. A log log n data structure for three-sided range queries. Information Processing Letters, 25(4):269–273, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fournier:1979:CCH

[Fou79] Alain Fournier. Comments on convex hull of a finite set of points in two dimensions [Inform. Process. Lett. 7 (1978), no. 6, 296–298; MR 80b:68041]. Information Processing Letters, 8(4):173, April 30, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Byk78, DS79, Ov80].

Farley:1982:DMC

[FP82] Arthur M. Farley and Andrzej Proskurowski. Directed maximalcut problems. Information Processing Letters, 15(5):238–241, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fischer:1983:SRF

[FP83] Michael J. Fischer and Michael S.

Paterson. Storage requirements for fair scheduling. Information Processing Letters, 17(5):249–250, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Friedberg:1987:ESM

[FP87] Stuart A. Friedberg and Gary L.

Peterson. An efficient solution to the mutual exclusion problem using weak semaphores. Information Processing Letters, 25(5):343–347, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fowler:1981:OPC

[FPT81] Robert J. Fowler, Michael S. Paterson, and Steven L. Tanimoto. Optimal packing and covering in the plane are NP-complete. Information Processing Letters, 12(3):133– 137, June 13, 1981. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fle:1984:MIT

[FR84] M. P. Flé and G. Roucairol. Multiserialization of iterated transactions. Information Processing Letters, 18 (5):243–247, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feitelson:1989:IWF

[FR89] Dror G. Feitelson and Larry Rudolph. Implementation of a waitfree synchronization primitive that solves n-process consensus. Information Processing Letters, 32(2):81– 83, July 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraser:1983:GTC

[Fra83] Christopher W. Fraser. A generalization of two code ordering optimizations. Information Processing Letters, 16(2):67–70, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Franco:1986:PPA

[Fra86] John Franco. On the probabilistic performance of algorithms for the satisfiability problem. Information Processing Letters, 23(2):103– 106, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Francez:1989:CPD

[Fra89] Nissim Francez. Cooperating proofs for distributed programs with multiparty interactions. Information Processing Letters, 32(5):235–242, September 22, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frederickson:1980:PAS

[Fre80] Greg N. Frederickson. Probabilistic analysis for simple one- and twodimensional bin packing algorithms. Information Processing Letters, 11 (4–5):156–161, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Freivalds:1981:PLR

[Fre81] Rūsinš Freivalds. Projections of languages recognizable by probabilistic and alternating finite multitape automata. Information Processing Letters, 13(4–5):195–198, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frederickson:1983:SUT

[Fre83a] Greg N. Frederickson. Scheduling unit-time tasks with integer release times and deadlines. Information Processing Letters, 16(4):171–173, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [SY93].

Freund:1983:IAO

[Fre83b] Rudolf Freund. Init and Anf operating on ω-languages. Information Processing Letters, 16(5):265– 269, June 10, 1983. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frederickson:1984:LTA

[Fre84] Greg N. Frederickson. On lineartime algorithms for five-coloring planar graphs. Information Processing Letters, 19(5):219–224, November 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frieze:1987:PAF

[Fri87] A. M. Frieze. Parallel algorithms for finding Hamilton cycles in random graphs. Information Processing Letters, 25(2):111–117, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frieze:1988:RCH

[Fri88] Alan M. Frieze. On the random construction of heaps. Information Processing Letters, 27(2):103– 109, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Frougny:1981:SDN

[Fro81] C. Frougny. Simple deterministic NTS languages. Information Processing Letters, 12(4):174–178, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Faigle:1984:MCT

[FS84] Ulrich Faigle and Rainer Schrader. Minimizing completion time for a class of scheduling problems. Information Processing Letters, 19 (1):27–29, July 26, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/0020019084901248.

Falkowski:1986:NQP

[FS86] Bernd-Jürgen J. Falkowski and Lothar Schmitz. A note on the queens’ problem. Information Processing Letters, 23(1):39–46, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Faigle:1988:CSD

[FS88a] Ulrich Faigle and Rainer Schrader. On the convergence of stationary distributions in simulated annealing algorithms. Information Processing Letters, 27(4):189–194, April 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fortnow:1988:TIP

[FS88b] Lance Fortnow and Michael Sipser.

Are there interactive protocols for CO-NP languages? Information Processing Letters, 28(5):249–251, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fushimi:1983:IOE

[Fus83] Masanori Fushimi. Increasing the orders of equidistribution of the leading bits of the Tausworthe sequence. Information Processing Letters, 16(4):189–192, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fagin:1983:ADF

[FV83] Ronald Fagin and Moshe Y. Vardi.

Armstrong databases for functional and inclusion dependencies. Information Processing Letters, 16(1):13– 19, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Foster:1989:PAH

[FV89a] Dean P. Foster and Rakesh V.

Vohra. Probabilistic analysis of a heuristics for the dual bin packing problem. Information Processing Letters, 31(6):287–290, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fulop:1989:TTT

[FV89b] Z. Fülöp and S. Vágvölgyi. Topdown tree transducers with deterministic top-down look-ahead. Information Processing Letters, 33(1): 3–5, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Feijen:1987:IIC

[FVG87] W. H. J. Feijen, A. J. M. Van Gasteren, and David Gries. Insitu inversion of a cyclic permutation. Information Processing Letters, 24(1):11–14, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Fraenkel:1980:CSA

[FY80] Aviezri S. Fraenkel and Yaacov Yesha. Complexity of solving algebraic equations. Information Processing Letters, 10(4–5):178–179, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Franchi-Zannettacci:1982:ETS

[FZ82] Paul Franchi-Zannettacci. An extension to trees of the Sardinas and Patterson algorithm. Information Processing Letters, 14(4):168–173, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gabow:1981:LTR

[Gab81] Harold N. Gabow. A lineartime recognition algorithm for interval dags. Information Processing Letters, 12(1):20–22, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Geist:1984:PCD

[Gei84] R. Geist. Perception-based configuration design of computer systems. Information Processing Letters, 18(1):55–57, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gelenbe:1982:SP

[GG82] E. Gelenbe and D. Gardy. On the size of projections. Information Processing Letters, 14(1):18–21, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gereb-Graus:1989:TNC

[GGPS89] Mihály Geréb-Graus, Ramamohan Paturi, and Endre Szemerédi. There are no p-complete families of symmetric Boolean functions. Information Processing Letters, 30(1):47– 49, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gunn:1981:DCP

[GH81] Hamish I. E. Gunn and David M.

Harland. Degrees of constancy in programming languages. Information Processing Letters, 13(1):35– 38, October 27, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1987:OTD

[GH87] Ajay K. Gupta and Susanne E.

Hambrusch. Optimal threedimensional layouts of complete binary trees. Information Processing Letters, 26(2):99–104, October 19, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gusfield:1989:PSM

[GI89] Dan Gusfield and Robert W. Irving. Parametric stable marriage and minimum cuts. Information Processing Letters, 30(5):255–259, March 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gilbert:1988:SND

[Gil88] John Russell Gilbert. Some nested dissection order is nearly optimal. Information Processing Letters, 26 (6):325–328, January 25, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gini:1982:ASI

[Gin82] Guiseppina Gini. The automatic synthesis of iterative programs. Information Processing Letters, 14(2): 67–73, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gire:1986:TDP

[Gir86] F. Gire. Two decidability problems for infinite words. Information Processing Letters, 22(3):135– 140, March 3, 1986. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gibbons:1988:PTE

[GIR88] Alan M. Gibbons, Amos Israeli, and Wojciech Rytter. Parallel O(log n) time edge-colouring of trees and Halin graphs. Information Processing Letters, 27(1):43–51, February 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gonzalez:1982:SNL

[GJ82] Teofilo F. Gonzalez and Donald B.

Johnson. Sorting numbers in linear expected time and optimal extra space. Information Processing Letters, 15(3):119–124, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goralcik:1986:VN

[GK86a] Pavel Goralčík and Václav Koubek.

Verifying nonrigidity. Information Processing Letters, 22(2):91– 95, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grigoriadis:1986:LBC

[GK86b] M. D. Grigoriadis and B. Kalantari.

A lower bound to the complexity of Euclidean and rectilinear matching algorithms. Information Processing Letters, 22(2):73–76, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gries:1980:CFN

[GL80] David Gries and Gary Levin. Computing Fibonacci numbers (and similarly defined functions) in log time. Information Processing Letters, 11 (2):68–69, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grossi:1989:SES

[GL89] R. Grossi and F. Luccio. Simple and efficient string matching with k mismatches. Information Processing Letters, 33(3):113–120, November 30, 1989. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Greenberg:1982:EPA

[GLPG82] Albert C. Greenberg, Richard E. Ladner, Michael S. Paterson, and Zvi Galil. Efficient parallel algorithms for linear recurrence computation. Information Processing Letters, 15(1):31–35, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Graham:1982:SER

[GM82] M. H. Graham and A. O. Mendelzon. Strong equivalence of relational expressions under dependencies. Information Processing Letters, 14(2): 57–62, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gazit:1988:IPA

[GM88] Hillel Gazit and Gary L. Miller. An improved parallel algorithm that computes the BFS numbering of a directed graph. Information Processing Letters, 28(2):61–65, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gerstel:1989:BCO

[GMZ89] O. Gerstel, Y. Mansour, and S. Zaks. Bit complexity of order statistics on a distributed star network. Information Processing Letters, 30(3):127–132, February 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldschlager:1980:SEA

[Gol80] Leslie M. Goldschlager. A space efficient algorithm for the monotone planar circuit value problem. Information Processing Letters, 10(1): 25–27, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Golumbic:1987:GMA

[Gol87] Martin Charles Golumbic. A general method for avoiding cycling in a network. Information Processing Letters, 24(4):251–253, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldman:1989:SEG

[Gol89] Sally A. Goldman. A space efficient greedy triangulation algorithm. Information Processing Letters, 31(4): 191–196, May 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goodrich:1987:FCH

[Goo87] Michael T. Goodrich. Finding the convex hull of a sorted point set in parallel. Information Processing Letters, 26(4):173–179, December 4, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gordon:1980:DSS

[Gor80] Michael J. C. Gordon. The denotational semantics of sequential machines. Information Processing Letters, 10(1):1–3, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gordon:1986:EFT

[Gor86] Dan Gordon. Eliminating the flag in threaded binary search trees. Information Processing Letters, 23(4): 209–214, November 8, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goscinski:1989:SAP

[Gos89] A. Goscinski. A synchronization algorithm for processes with dynamic priorities in computer networks with node failures. Information Processing Letters, 32(3):129–136, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gottlob:1987:SNF

[Got87a] Georg Gottlob. On the size of nonredundant FD-covers. Information Processing Letters, 24(6):355– 360, April 6, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gottlob:1987:SI

[Got87b] Georg Gottlob. Subsumption and implication. Information Processing Letters, 24(2):109–111, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gusfield:1986:EAA

[GP86] Dan Gusfield and Leonard Pitt. Equivalent approximation algorithms for node cover. Information Processing Letters, 22(6):291–294, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldberg:1987:PCC

[GP87a] Andrew V. Goldberg and Serge A.

Plotkin. Parallel (δ + 1)-coloring of constant-degree graphs. Information Processing Letters, 25(4):241– 245, June 17, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guntzer:1987:JIS

[GP87b] U. Guntzer and M. Paul. Jump interpolation search trees and symmetric binary numbers. Information Processing Letters, 26(4):193– 204, December 4, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guessarian:1988:MNO

[GP88] Irène Guessarian and Lutz Priese.

On the minimal number of × operators to model regularity in fair SCCS. Information Processing Letters, 29(6):297–300, December 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galil:1989:PED

[GP89] Zvi Galil and Victor Pan. Parallel evaluation of the determinant and of the inverse of a matrix. Information Processing Letters, 30(1): 41–45, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldberg:1982:ATA

[GPB82] Allen Goldberg, Paul Purdom, and Cynthia Brown. Average time analyses of simplified Davis–Putnam procedures. Information Processing Letters, 15(2):72–75, September 6, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [GPB83].

Goldberg:1983:CAT

[GPB83] Allen Goldberg, Paul Purdom, and Cynthia Brown. Corrigendum: “Average time analyses of simplified Davis–Putnam procedures” [Inform. Process. Lett. 15 (1982), no. 2, 72– 75]. Information Processing Letters, 16(4):213, May 13, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [GPB82].

Gries:1987:HRC

[GPS87] David Gries, Adriano Pascoletti, and Luigi Sbriz. Horner’s rule and the computation of linear recurrences. Information Processing Letters, 25(4):237–240, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goh:1982:RPE

[GR82] L. Goh and D. Rotem. Recognition of perfect elimination bipartite graphs. Information Processing Letters, 15(4):179–182, October 31, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Graham:1972:EAD

[Gra72] R. L. Graham. An efficient algorithm for determining the convex hull of a finite planar set. Information Processing Letters, 1(4): 132–133, June ??, 1972. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [GS87].

Grabowski:1980:DPV

[Gra80] Jan Grabowski. The decidability of persistence for vector addition systems. Information Processing Letters, 11(1):20–23, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grazon:1987:IWL

[Gra87] Anne Grazon. An infinite word language which is not co-CFL. Information Processing Letters, 24(2): 81–85, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gregori:1989:ULE

[Gre89] Angelo Gregori. Unit-length embedding of binary trees on a square grid. Information Processing Letters, 31 (4):167–173, May 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grossi:1991:FCS

[Gro91] Roberto Grossi. Further comments on the subtree isomorphism for ordered trees: “On the subtree isomorphism problem for ordered trees” [Inform. Process. Lett. 32 (1989), no. 5, 271–273; MR 90k:68139] by E. Mäkinen. Information Processing Letters, 40(5):255–256, December 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Mäk89].

Grune:1984:HPA

[Gru84] Dick Grune. How to produce all sentences from a two-level grammar. Information Processing Letters, 19(4):181–185, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Grune:1987:HCI

[Gru87] Dick Grune. How to compare the incomparable. Information Processing Letters, 24(3):177–181, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guida:1981:MKR

[GS81] G. Guida and M. Somalvico. Multiproblem-solving: knowledge representation and system architecture. Information Processing Letters, 13 (4–5):204–214, ???? ??, 1981. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guibas:1983:CAN

[GS83] Leo J. Guibas and Jorge Stolfi. On computing all north-east nearest neighbors in the L_{1} metric. Information Processing Letters, 17(4):219– 223, November 8, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1984:SSA

[GS84a] R. G. Gupta and V. S. P. Srivastava. On synthesis of scheduling algorithms. Information Processing Letters, 19(3):147–150, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gursel:1984:AOS

[GS84b] Goker Gursel and Peter Scheuermann. Asserting the optimality of serial SJRPs in processing simple queries in chain networks. Information Processing Letters, 19(5):255– 260, November 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gupta:1986:ASU

[GS86] G. K. Gupta and B. Srinivasan. Approximate storage utilization of Btrees. Information Processing Letters, 22(5):243–246, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gries:1987:NGC

[GS87] David Gries and Ivan Stojmenovic.

A note on Graham’s convex hull algorithm. Information Processing Letters, 25(5):323–327, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Gra72].

Gallo:1988:PSS

[GS88a] Giorgio Gallo and Maria Grazia Scutellà. Polynomially solvable satisfiability problems. Information Processing Letters, 29(5):221–227, November 24, 1988. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gopalakrishnan:1988:IFP

[GS88b] Ganesh C. Gopalakrishnan and Mandayam K. Srivas. Implementing functional programs using mutable abstract data types. Information Processing Letters, 26(6):277– 286, January 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gibbons:1989:CPE

[GS89] A. M. Gibbons and Y. N. Srikant.

A class of problems efficiently solvable on mesh-connected computers including dynamic expression evaluation. Information Processing Letters, 32(6):305–311, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ginat:1989:TAB

[GST89] David Ginat, Daniel D. Sleator, and Robert E. Tarjan. A tight amortized bound for path reversal. Information Processing Letters, 31(1):3– 5, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gangolli:1983:TPM

[GT83] Anil R. Gangolli and Steven L. Tanimoto. Two pyramid machine algorithms for edge detection in noisy binary images. Information Processing Letters, 17(4):197–202, November 8, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goodman:1984:CMD

[GT84] Nathan Goodman and Y. C. Tay. A characterization of multivalued dependencies equivalent to a join dependency. Information Processing Letters, 18(5):261–266, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gajewska:1986:DHO

[GT86] Hania Gajewska and Robert E. Tarjan. Deques with heap order. Information Processing Letters, 22(4): 197–200, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gabow:1988:LTA

[GT88] Harold N. Gabow and Robert E.

Tarjan. A linear-time algorithm for finding a minimum spanning pseudoforest. Information Processing Letters, 27(5):259–263, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Goldberg:1989:PAF

[GT89] Andrew V. Goldberg and Robert E.

Tarjan. A parallel algorithm for finding a blocking flow in an acyclic network. Information Processing Letters, 31(5):265–271, June 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gunn:1982:CTT

[Gun82] Hamish I. E. Gunn. Compile time type checking of structure field accessing. Information Processing Letters, 14(1):22–25, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gunther:1989:PIM

[Gun89] Neil Gunther. Path integral methods for computer performance analysis. Information Processing Letters, 32(1):7–13, July 3, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Gusfield:1983:CED

[Gus83] Dan Gusfield. Connectivity and edge-disjoint spanning trees. Information Processing Letters, 16(2):87– 89, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guting:1983:SOP

[Güt83] Ralf Hartmut Güting. Stabbing Coriented polygons. Information Processing Letters, 16(1):35–40, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guting:1985:FDI

[Güt85] Ralf Hartmut Güting. Fast dynamic intersection searching in a set of isothetic line segments. Information Processing Letters, 21(4):165– 171, October 7, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Galil:1987:PEA

[GY87] Zvi Galil and Moti Yung. Partitioned encryption and achieving simultaneity by partitioning. Information Processing Letters, 26(2):81– 88, October 19, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Guralnik:1985:AUR

[GZW85] Gerald Guralnik, Charles Zemach, and Tony Warnock. An algorithm for uniform random sampling of points in and on a hypersphere. Information Processing Letters, 21 (1):17–21, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hagerup:1988:SSP

[Hag88] Torben Hagerup. On saving space in parallel computation. Information Processing Letters, 29(6):327– 329, December 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hagerup:1989:HRP

[Hag89] Torben Hagerup. Hybridsort revisited and parallelized. Information Processing Letters, 32(1):35–39, July 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hamburg:1986:TTV

[Ham86] Mark C. Hamburg. Two tagless variations on the Deutsch–Schorr– Waite algorithm. Information Processing Letters, 22(4):179–183, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hamlet:1987:PCT

[Ham87] Richard G. Hamlet. Probable correctness theory. Information Processing Letters, 25(1):17–25, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hanson:1980:CIL

[Han80] D. R. Hanson. Code improvement via lazy evaluation. Information Processing Letters, 11(4–5):163–167, December 12, 1980. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Handley:1986:SDS

[Han86] C. C. Handley. An in situ distributive sort. Information Processing Letters, 23(5):265–270, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harland:1981:CLE

[Har81a] D. M. Harland. Concurrency in a language employing messages. Information Processing Letters, 12(2): 59–62, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harland:1981:FIC

[Har81b] David M. Harland. On facilities for interprocess communication. Information Processing Letters, 12(5): 221–226, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hart:1982:PIM

[Har82] Johnson M. Hart. Permutation inversions and multidimensional cumulative distribution functions. Information Processing Letters, 14(5): 218–222, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hartmanis:1983:SS

[Har83] Juris Hartmanis. On sparse sets in NP − P. Information Processing Letters, 16(2):55–60, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hartmanis:1985:IRA

[Har85] Juris Hartmanis. Independence results about context-free languages and lower bounds. Information Processing Letters, 20(5):241–248, June 12, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hartley:1989:DPG

[Har89] Richard I. Hartley. Drawing polygons given angle sequences. Information Processing Letters, 31(1): 31–33, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hassin:1981:MFP

[Has81] Refael Hassin. Maximum flow in (s, t) planar networks. Information Processing Letters, 13(3):107, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Haastad:1987:OWP

[Hås87] Johan Håstad. One-way permutations in NC^{0}. Information Processing Letters, 26(3):153–155, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hattori:1986:NGS

[Hat86] Yoshio Hattori. Nonisomorphic graphs with the same T -polynomial. Information Processing Letters, 22 (3):133–134, March 3, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Herman:1985:DS

[HC85] Ted Herman and K. Mani Chandy.

On distributed search. Information Processing Letters, 21(3):129– 133, September 5, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Head:1989:DIR

[Hea89] Tom Head. Deciding the immutability of regular codes and languages under finite transductions. Information Processing Letters, 31(5):239– 241, June 12, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hehner:1981:BTS

[Heh81] Eric C. R. Hehner. Bunch theory: a simple set theory for computer science. Information Processing Letters, 12(1):26–30, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hehner:1989:RP

[Heh89] Eric C. R. Hehner. Real-time programming. Information Processing Letters, 30(1):51–56, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hellman:1981:ACA

[Hel81] M. E. Hellman. Another cryptanalytic attack on “A cryptosystem for multiple communication” [Inform. Process. Lett. 10(4–5), 5 July 1980, pp. 180–183]. Information Processing Letters, 12(4):182–183, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LM80b, Mei81].

Hemerik:1980:FDL

[Hem80] C. Hemerik. Formal derivation of a list processing program. Information Processing Letters, 10(3):124– 126, April 18, 1980. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hertel:1983:SBP

[Her83] Stefan Hertel. Smoothsort’s behavior on presorted sequences. Information Processing Letters, 16(4): 165–170, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hershberger:1989:FUE

[Her89] John Hershberger. Finding the upper envelope of n line segments in O(n log n) time. Information Processing Letters, 33(4):169–174, December 21, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hanatani:1986:SCD

[HF86] Yoshito Hanatani and Ronald Fagin. A simple characterization of database dependency implication. Information Processing Letters, 22 (6):281–283, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoare:1987:WP

[HH87] C. A. R. Hoare and Jifeng He. The weakest prespecification. Information Processing Letters, 24(2):127– 132, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hartmanis:1988:SOS

[HH88] Juris Hartmanis and Lane Hemachandra. On sparse oracles separating feasible complexity classes. Information Processing Letters, 28(6): 291–295, August 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoare:1987:PDR

[HHS87] C. A. R. Hoare, Jifeng He, and J. W. Sanders. Prespecification in data refinement. Information Processing Letters, 25(2):71–76, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Highnam:1982:EPC

[Hig82] P. T. Highnam. The ears of a polygon. Information Processing Letters, 15(5):196–198, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Highnam:1986:OAF

[Hig86] P. T. Highnam. Optimal algorithms for finding the symmetries of a planar point set. Information Processing Letters, 22(5):219–222, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hikita:1983:LCS

[Hik83] Teruo Hikita. Listing and counting subtrees of equal size of a binary tree. Information Processing Letters, 17(4):225–229, November 8, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hatzopoulos:1981:OPD

[HK81a] M. Hatzopoulos and J. G. Kollias. Optimal policy for database backup and recovery. Information Processing Letters, 12(2):55–58, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hell:1981:GMP

[HK81b] P. Hell and D. G. Kirkpatrick. On generalized matching problems. Information Processing Letters, 12(1): 33–35, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoffmann:1988:ERG

[HK88] Frank Hoffmann and Klaus Kriegel.

Embedding rectilinear graphs in linear time. Information Processing Letters, 29(2):75–79, September 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Horspool:1987:CEO

[HL87a] R. Nigel Horspool and Michael R.

Levy. Correctness of an extended operator-precedence parsing algorithm. Information Processing Letters, 24(4):265–273, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hyslop:1987:PDP

[HL87b] Gary A. Hyslop and Edmund A. Lamagna. Performance of distributive partitioned sort in a demand paging environment. Information Processing Letters, 25(1):61–64, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho:1988:EPA

[HL88] Chin Wen Ho and Richard C. T.

Lee. Efficient parallel algorithms for finding maximal cliques, clique trees, and minimum coloring on chordal graphs. Information Processing Letters, 28(6):301–309, August 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ho:1989:CCT

[HL89] Chin-Wen W. Ho and R. C. T. Lee. Counting clique trees and computing perfect elimination schemes in parallel. Information Processing Letters, 31(2):61–68, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Honeyman:1980:TUI

[HLY80] Peter Honeyman, Richard E. Ladner, and Mihalis Yannakakis. Testing the universal instance assumption. Information Processing Letters, 10(1):14–19, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hood:1981:RQO

[HM81] R. Hood and R. Melville. Real-time queue operations in Pure LISP. Information Processing Letters, 13(2): 50–54, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsu:1988:STC

[HM88] Meichun Hsu and Stuart E. Madnick. Shifting timestamps for concurrency control in an information hierarchy. Information Processing Letters, 27(6):291–297, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hinrichs:1988:PSS

[HNS88] Klaus Hinrichs, Jürg Nievergelt, and Peter Schorn. Plane-sweep solves the closest pair problem elegantly. Information Processing Letters, 26 (5):255–261, January 11, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hochschild:1987:MCI

[Hoc87] Peter Hochschild. Multiple cuts, input repetition, and VLSI complexity. Information Processing Letters, 24(1):19–24, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hofri:1983:STH

[Hof83] Micha Hofri. Should the two-headed disk be greedy? — yes, it should. Information Processing Letters, 16(2): 83–85, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Holenderski:1984:NSV

[Hol84] Leszek Holenderski. A note on specifying and verifying concurrent processes. Information Processing Letters, 18(2):77–85, February 28, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Holenderski:1986:CNP

[Hol86] Leszek Holenderski. The correctness of nondeterministic programs revisited. Information Processing Letters, 23(6):299–303, December 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoogerwoord:1986:IMI

[Hoo86] Rob R. Hoogerwoord. An implementation of mutual inclusion. Information Processing Letters, 23(2):77–80, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hou:1987:TAA

[Hou87] Y. Hou. Trinity algebra and its application to machine decompositions. Information Processing Letters, 26(3):127–134, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harel:1985:MLV

[HP85] David Harel and David Peleg. More on looping vs. repeating in dynamic logic. Information Processing Letters, 20(2):87–90, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hagerup:1989:OMS

[HR89] Torben Hagerup and Christine Rüb. Optimal merging and sorting on the EREW PRAM. Information Processing Letters, 33(4):181–185, December 21, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hromkovic:1985:AMM

[Hro85a] Juraj Hromkovič. Alternating multicounter machines with constant number of reversals. Information Processing Letters, 21(1):7–9, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hromkovic:1985:LLB

[Hro85b] Juraj Hromkovič. Linear lower bounds on unbounded fan-in Boolean circuits. Information Processing Letters, 21(2):71–74, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Howell:1987:ADB

[HRY87] Rodney R. Howell, Louis E. Rosier, and Hsu-Chun Yen. An O(n^{1.5}) algorithm to decide boundedness for conflict-free vector replacement systems. Information Processing Letters, 25(1):27–33, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hehner:1981:IPV

[HS81] E. C. R. Hehner and R. K. Shyamasundar. An implementation of P and V. Information Processing Letters, 12(4):196–198, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hsiao:1985:KBS

[HS85] Ching C. Hsiao and Nien-Tsu Shen.

k-fold bitonic sort on a meshconnected parallel computer. Information Processing Letters, 21(4): 207–212, October 7, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Harary:1986:LAC

[HS86] Frank Harary and Peter J. Slater. A linear algorithm for the cutting center of a tree. Information Processing Letters, 23(6):317–319, December 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Haldar:1988:RBT

[HS88a] S. Haldar and D. K. Subramanian. Ring based termination detection algorithm for distributed computations. Information Processing Letters, 29(3):149–153, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [TM89].

Holenderski:1988:PDF

[HS88b] Leszek Holenderski and Andrzej Szałas. Propositional description of finite cause-effect structures. Information Processing Letters, 27(3): 111–117, March 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hochbaum:1989:AMW

[HS89] Dorit S. Hochbaum and Ron Shamir. An O(n log^{2} n) algorithm for the maximum weighted tardiness problem. Information Processing Letters, 31(4):215–219, May 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1981:AIP

[Hua81] Bing Chao Huang. An algorithm for inverting a permutation. Information Processing Letters, 12(5):237– 238, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1988:FDT

[Hua88a] Shing Tsaan Huang. A fully distributed termination detection scheme. Information Processing Letters, 29(1):13–18, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1988:LBE

[Hua88b] Xiaoqiu Huang. A lower bound for the edit-distance problem under an arbitrary cost function. Information Processing Letters, 27(6):

319–321, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huang:1989:TDU

[Hua89] Shing-Tsaan Huang. Termination detection by using distributed snapshots. Information Processing Letters, 32(3):113–119, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hughes:1986:NRL

[Hug86] R. John Muir Hughes. A novel representation of lists and its application to the function ‘reverse’. Information Processing Letters, 22(3): 141–144, March 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hull:1984:PVS

[Hul84] M. Elizabeth C. Hull. A parallel view of stable marriages. Information Processing Letters, 18(2): 63–66, February 28, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Huynh:1987:SHP

[Huỳ87] Dũng T. Huỳnh. On solving hard problems by polynomial-size circuits. Information Processing Letters, 24(3):171–176, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hirschberg:1987:IUQ

[HV87] D. S. Hirschberg and Dennis James Volper. Improved update/query algorithms for the interval valuation problem. Information Processing Letters, 24(5):307–310, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hoshi:1982:CEM

[HY82] Mamoru Hoshi and Toshitsugu Yuba. A counter example to a monotonicity property of k-d trees. Information Processing Letters, 15(4): 169–173, October 31, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Hazari:1987:DAD

[HZ87] Cyrus Hazari and Hussein Zedan. A distributed algorithm for distributed termination. Information Processing Letters, 24(5):293–297, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [TV87].

Ida:1983:SFA

[Ida83] T. Ida. Some FP algebra with Currying operation. Information Processing Letters, 17(5):259–261, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Israeli:1986:FSR

[II86] Amos Israeli and A. Itai. A fast and simple randomized parallel algorithm for maximal matching. Information Processing Letters, 22(2): 77–80, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:1989:OST

[IJ89] Oscar H. Ibarra and Tao Jiang. Optimal simulation of tree arrays by linear arrays. Information Processing Letters, 30(6):295–302, March 28, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:1988:SSC

[IJR88] Oscar H. Ibarra, Tao Jiang, and Bala Ravikumar. Some subclasses of context-free languages in NC^{1}. Information Processing Letters, 29(3): 111–117, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibaraki:1983:LCT

[IK83] T. Ibaraki and N. Katoh. On-line computation of transitive closures of graphs. Information Processing Letters, 16(2):95–97, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iliopoulos:1985:AAP

[Ili85] Costas S. Iliopoulos. Analysis of algorithms on problems in general Abelian groups. Information Processing Letters, 20(4):215–220, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:1981:DPA

[IM81] Oscar H. Ibarra and Shlomo Moran.

Deterministic and probabilistic algorithms for maximum bipartite matching via fast matrix multiplication. Information Processing Letters, 13(1):12–15, October 27, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Imase:1988:FTR

[IM88] Makoto Imase and Yoshifumi Manabe. Fault-tolerant routings in a κ-connected network. Information Processing Letters, 28(4):171–175, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Imai:1982:FCC

[Ima82] Hiroshi Imai. Finding connected components of an intersection graph of squares in the Euclidean plane. Information Processing Letters, 15 (3):125–128, October 11, 1982. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iri:1981:LTA

[IMM81] Masao Iri, Kazuo Murota, and Shouichi Matsui. Linear-time approximation algorithms for finding the minimum-weight perfect matching on a plane. Information Processing Letters, 12(4):206–209, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:1980:NPC

[IMR80] Oscar H. Ibarra, Shlomo Moran, and Louis E. Rosier. A note on the parallel complexity of computing the rank of order n matrices. Information Processing Letters, 11(4–5):162, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:1981:PAS

[IMR81] Oscar H. Ibarra, Shlomo Moran, and Louis E. Rosier. Probabilistic algorithms and straight-line programs for some rank decision problems. Information Processing Letters, 12(5): 227–232, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ihm:1984:LPP

[IN84] Heung-Soon S. Ihm and Simeon C.

Ntafos. On legal path problems in digraphs. Information Processing Letters, 18(2):93–98, February 28, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ibarra:1981:DED

[IR81] Oscar H. Ibarra and Louis E. Rosier.

On the decidability of equivalence for deterministic pushdown transducers. Information Processing Letters, 13(3):89–93, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iyer:1988:ONR

[IRV88] Ananth V. Iyer, H. Donald Ratliff, and G. Vijayan. Optimal node ranking of trees. Information Processing Letters, 28(5):225–229, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Israeli:1986:IPA

[IS86] Amos Israeli and Y. Shiloach.

An improved parallel algorithm for maximal matching. Information Processing Letters, 22(2):57– 60, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Incerpi:1987:PVS

[IS87] Janet Incerpi and Robert Sedgewick.

Practical variations of Shellsort. Information Processing Letters, 26(1): 37–43, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Itano:1988:IPM

[ISHY88] Kozo Itano, Yutaka Sato, Hidemi Hirai, and Tomoyoshi Yamagata. An incremental pattern matching algorithm for the pipelined lexical scanner. Information Processing Letters, 27(5):253–258, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1980:NDP

[IT80] Katsushi Inoue and Itsuo Takanami. A note on decision problems for three-way two-dimensional finite automata. Information Processing Letters, 10(4–5):245–248, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1988:SCA

[IT88] Katsushi Inoue and Itsuo Takanami. Some considerations about NPRIORITY(1) without ROM. Information Processing Letters, 28(4): 215–219, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Itoh:1989:EAD

[IT89] Toshiya Itoh and Shigeo Tsujii. An efficient algorithm for deciding quadratic residuosity in finite fields GF(p^{m}). Information Processing Letters, 30(3):111–114, February 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Italiano:1988:FPD

[Ita88] Giuseppe F. Italiano. Finding paths and deleting edges in directed acyclic graphs. Information Processing Letters, 28(1):5–11, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Inoue:1982:NAL

[ITT82] Katsushi Inoue, Itsuo Takanami, and Hiroshi Taniguchi. A note on alternating on-line Turing machines. Information Processing Letters, 15 (4):164–168, October 31, 1982. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ingemarsson:1981:UAS

[IW81] Ingemar Ingemarsson and C. K.

Wong. A user authentication scheme for shared data based on a trap-door one-way function. Information Processing Letters, 12(2):63–67, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Iwano:1989:IGP

[Iwa89] Kazuo Iwano. An improvement of Goldberg, Plotkin and Vaidya’s maximal node-disjoint paths algorithm. Information Processing Letters, 32(1):25–27, July 3, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jaffar:1981:PAA

[Jaf81] Joxan Jaffar. Presburger arithmetic with array segments. Information Processing Letters, 12(2):79–82, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jajodia:1985:ERN

[Jaj85] Sushil Jajodia. On equivalence of relational and network database models. Information Processing Letters, 20(1):51–54, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jakobsson:1982:EHB

[Jak82] M. Jakobsson. Evaluation of a hierarchical bit-vector compression technique. Information Processing Letters, 14(4):147–149, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jantzen:1985:NSO

[Jan85] M. Jantzen. A note on a special onerule semi-Thue system. Information Processing Letters, 21(3):135– 140, September 5, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Janiak:1989:MRC

[Jan89] Adam Janiak. Minimization of resource consumption under a given deadline in the two-processor flowshop scheduling problem. Information Processing Letters, 32(3):101– 112, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jaromczyk:1981:LDT

[Jar81] Jerzy W. Jaromczyk. Linear decision trees are too weak for convex hull problem. Information Processing Letters, 12(3):138–141, June 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jazayeri:1980:IID

[Jaz80] M. Jazayeri. An improvement in the iterative data flow analysis algorithm. Information Processing Letters, 10(2):108–110, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johnson:1984:FSM

[JB84] Dale Johnson and Barrett R.

Bryant. Formal syntax methods for natural language. Information Processing Letters, 19(3):135–143, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jedrzejowicz:1983:ECR

[Jȩd83] Joanna Jȩdrzejowicz. On the enlargement of the class of regular languages by the shuffle closure. Information Processing Letters, 16(2): 51–54, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jedrzejowicz:1987:NSC

[Jȩd87] Joanna Jȩdrzejowicz. Nesting of shuffle closure is important. Information Processing Letters, 25(6): 363–367, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jedrzejowicz:1988:IHE

[Jȩd88] Joanna Jȩdrzejowicz. Infinite hierarchy of expressions containing shuffle closure operator. Information Processing Letters, 28(1):33–37, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jou:1982:CRR

[JF82] Jiann H. Jou and Patrick C. Fischer. The complexity of recognizing 3NF relation schemes. Information Processing Letters, 14(4):187–190, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jin:1984:LPG

[Jin84] Xiao Long Jin. Large processors are good in VLSI chips. Information Processing Letters, 18(1):47– 49, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Janiga:1985:NFM

[JK85] Ladislav Janiga and Václav Koubek. A note on finding minimum cuts in directed planar networks by parallel computations. Information Processing Letters, 21(2):75–78, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jouannaud:1982:MO

[JL82] Jean Pierre Jouannaud and Pierre Lescanne. On multiset orderings. Information Processing Letters, 15 (2):57–63, September 6, 1982. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jaumard:1986:EAT

[JM86] Brigitte Jaumard and Michel Minoux. An efficient algorithm for the transitive closure and a linear worstcase complexity result for a class of sparse graphs. Information Processing Letters, 22(4):163–169, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jung:1988:PAC

[JM88] Hermann Jung and Kurt Mehlhorn.

Parallel algorithms for computing maximal independent sets in trees and for updating minimum spanning trees. Information Processing Letters, 27(5):227–236, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johnson:1980:CVA

[Joh80] Robert B. Johnson, Jr. The complexity of a VLSI adder. Information Processing Letters, 11(2):92–93, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jokinen:1989:CGC

[Jok89] Matti O. Jokinen. Customizable garbage collectors. Information Processing Letters, 30(3):115–118, February 13, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Joseph:1985:PRC

[Jos85] Mathai Joseph. On a problem in real-time computing. Information Processing Letters, 20(4):173–177, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Josephs:1988:DRC

[Jos88] Mark B. Josephs. The data refinement calculator for Z specifications. Information Processing Letters, 27 (1):29–33, February 15, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jammel:1980:ECD

[JS80] Alfons J. Jammel and Helmut G. Stiegler. On expected costs of deadlock detection. Information Processing Letters, 11(4–5):229–231, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jaumard:1987:CMS

[JS87] Brigitte Jaumard and Bruno Simeone. On the complexity of the maximum satisfiability problem for Horn formulas. Information Processing Letters, 26(1):1–4, September 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

JaJa:1985:ILB

[JT85] Joseph JáJá and Jean Takche. Improved lower bounds for some matrix multiplication problems. Information Processing Letters, 21(3): 123–127, September 5, 1985. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Jacobs:1986:TRT

[JV86] Christiaan T. M. Jacobs and Peter Van Emde Boas. Two results on tables. Information Processing Letters, 22(1):43–48, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Johnson:1988:GAM

[JYP88] David S. Johnson, Mihalis Yannakakis, and Christos H. Papadimitriou. On generating all maximal independent sets. Information Processing Letters, 27(3):119–123, March 25, 1988. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Kak:1985:HDT

[Kak85] Subhash C. Kak. How to detect tampering of data. Information Processing Letters, 20(2):109– 110, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kallay:1984:CIC

[Kal84] Michael Kallay. The complexity of incremental convex hull algorithms in R^{d}. Information Processing Letters, 19(4):197, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kaldewaij:1985:DSA

[Kal85] Anne Kaldewaij. On the decomposition of sequences into ascending subsequences. Information Processing Letters, 21(2):69, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kallay:1986:CHM

[Kal86] Michael Kallay. Convex hull made easy. Information Processing Letters, 22(3):161, March 3, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kanellakis:1980:CCC

[Kan80] Paris C. Kanellakis. On the computational complexity of cardinality constraints in relational databases. Information Processing Letters, 11 (2):98–101, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kantabutra:1983:TSC

[Kan83] Vitit Kantabutra. Traveling salesman cycles are not always subgraphs of Voronoĭ duals. Information Processing Letters, 16(1):11–12, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kasyanov:1984:LC

[Kas84] V. N. Kas′yanov. Loop cleaning. Information Processing Letters, 18 (1):1–6, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katsuno:1984:WDN

[Kat84] Hirofumi Katsuno. When do non-conflict-free multivalued dependency sets appear? Information Processing Letters, 18(2):87– 92, February 28, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kawai:1980:FSP

[Kaw80] Hiroya Kawai. A formal system for parallel programs in discrete time and space. Information Processing Letters, 11(4–5):204–210, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kawai:1987:LAI

[Kaw87] Satoru Kawai. Local authentication in insecure environments. Information Processing Letters, 25(3): 171–174, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1987:PMM

[KC87] Taenam Kim and Kyung-Yong Y.

Chwa. An O(n log n log log n) parallel maximum matching algorithm for bipartite graphs. Information Processing Letters, 24(1):15– 17, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keil:1985:FHC

[Kei85] J. Mark Keil. Finding Hamiltonian circuits in interval graphs. Information Processing Letters, 20(4): 201–206, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keil:1986:TDI

[Kei86] J. Mark Keil. Total domination in interval graphs. Information Processing Letters, 22(4):171–174, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keller:1986:STP

[Kel86] Arthur M. Keller. Set-theoretic problems of null completion in relational databases. Information Processing Letters, 22(5):261–265, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kelter:1987:CSS

[Kel87] Udo Kelter. The complexity of strict serializability revisited. Information Processing Letters, 25(6): 407–411, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kelter:1989:PPS

[Kel89] Udo Kelter. The pitfall paradox and its solution with virtual objects. Information Processing Letters, 30(3): 139–143, February 13, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kemp:1987:NNL

[Kem87] Rainer Kemp. A note on the number of leftist trees. Information Processing Letters, 25(4):227–232, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kemp:1989:OOC

[Kem89] R. Kemp. A one-to-one correspondence between two classes of ordered trees. Information Processing Letters, 32(5):229–234, September 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kessels:1980:RWP

[Kes80] J. L. W. I. Kessels. The readers and writers problem avoided. Information Processing Letters, 10(3): 159–162, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kessels:1988:EPS

[Kes88] J. L. W. Kessels. An exercise in proving self-stabilization with a variant function. Information Processing Letters, 29(1):39– 42, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Keedy:1985:EUS

[KF85] J. L. Keedy and B. Freisleben.

On the efficient use of semaphore primitives. Information Processing Letters, 21(4):199–205, October 7, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kfoury:1985:UPP

[Kfo85] A. J. Kfoury. The unwind property for programs with bounded memory. Information Processing Letters, 21(5):233–238, November 18, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Khuller:1989:CGC

[Khu89] Samir Khuller. On computing graph closures. Information Processing Letters, 31(5):249–255, June 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kieburtz:1986:WCY

[Kie86] Richard B. Kieburtz. When chasing your tail saves time graph theory. Information Processing Letters, 23(6):321–324, December 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kinber:1980:IPD

[Kin80] E. B. Kinber. On inclusion problem for deterministic multitape automata. Information Processing Letters, 11(3):144–146, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kirkpatrick:1980:NDO

[Kir80] David G. Kirkpatrick. A note on Delaunay and optimal triangulations.

Information Processing Letters, 10 (3):127–128, April ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kirchherr:1988:TMR

[Kir88] Walter W. Kirchherr. Transportation of an l × l matrix requires Ω(log l) reversals on conservative Turing machines. Information Processing Letters, 28(2):55–59, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kriegel:1983:ISB

[KK83] H. P. Kriegel and Y. S. Kwong. Insertion-safeness in balanced trees. Information Processing Letters, 16 (5):259–264, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kapur:1985:WCC

[KK85] Deepak Kapur and Mukkai S. Krishnamoorthy. Worst-case choice for the stable marriage problem. Information Processing Letters, 21(1): 27–30, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comment [PG87].

Kamada:1989:ADG

[KK89a] Tomihisa Kamada and Satoru Kawai. An algorithm for drawing general undirected graphs. Information Processing Letters, 31(1):7–15, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1989:CMM

[KK89b] Young Joo Kim and Gil Chang Kim.

Coordinator: a modification to the monitor concept. Information Processing Letters, 32(2):73–80, July 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kirsig:1987:SRS

[KL87] Bernd Kirsig and Klaus-Jörn J.

Lange. Separation with the Ruzzo, Simon, and Tompa relativization implies Dspace(log n) ≠ Nspace(log n). Information Processing Letters, 25(1):13–15, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korel:1988:DPS

[KL88] Bogdan Korel and Janusz Laski. Dynamic program slicing. Information Processing Letters, 29(3):155– 163, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Klapper:1987:LBC

[Kla87] Andrew Klapper. A lower bound on the complexity of the convex hull problem for simple polyhedra. Information Processing Letters, 25(3): 159–161, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Klupsz:1984:LAD

[Klu84] Ewa Klupsz. A linear algorithm of a deadlock avoidance for nonpreemptible resources. Information Processing Letters, 19(2):87–94, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krause:1987:PIT

[KLV87] Kadri Krause, Lawrence L. Larmore, and Dennis James Volper. Packing items from a triangular distribution. Information Processing Letters, 25(6):351–361, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kandzia:1980:CBC

[KM80] Peter Kandzia and Margret Mangelmann. On covering Boyce-Codd normal forms. Information Processing Letters, 11(4–5):218–223, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumar:1988:NCR

[KM88] Ashok Kumar and V. M. Malhotra. A new computation rule for Prolog. Information Processing Letters, 27(5):249–252, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kakuta:1986:PRC

[KNI86] K. Kakuta, H. Nakamura, and S. Iida. A parallel reference counting algorithm. Information Processing Letters, 23(1):33–37, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Katajainen:1987:LET

[KNT87] Jyrki Katajainen, Olli Nevalainen, and Jukka Teuhola. A linear expected-time algorithm for computing planar relative neighbourhood graphs. Information Processing Letters, 25(2):77–86, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ko:1982:SOP

[Ko82] Ker-I Ko. Some observations on the probabilistic algorithms and NPhard problems. Information Processing Letters, 14(1):39–43, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karlsson:1988:NDC

[KO88] Rolf G. Karlsson and Mark H.

Overmars. Normalized divide-and-conquer: a scaling technique for solving multi-dimensional problems. Information Processing Letters, 26 (6):307–312, January 25, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kobuchi:1983:SDS

[Kob83] Youichi Kobuchi. Stability of desynchronized 0L systems. Information Processing Letters, 17(2):85–90, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kobuchi:1987:NSC

[Kob87] Youichi Kobuchi. A note on symmetrical cellular spaces. Information Processing Letters, 25(6):413– 415, July 26, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koch:1987:ASC

[Koc87] Gregers Koch. Automating the semantic component. Information Processing Letters, 24(5):299–305, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kochol:1989:EMC

[Koc89] Martin Kochol. Efficient monotone circuits for threshold functions. Information Processing Letters, 32(3): 121–122, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comment [Dun90].

Kolla:1983:WOS

[Kol83] Reiner Kolla. Where-oblivious is not sufficient. Information Processing Letters, 17(5):263–268, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korsh:1981:GBS

[Kor81] James F. Korsh. Greedy binary search trees are nearly optimal. Information Processing Letters, 13(1): 16–19, October 27, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korsh:1982:GNO

[Kor82] J. F. Korsh. Growing nearly optimal binary search trees. Information Processing Letters, 14(3):139– 143, May 16, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korel:1987:PDG

[Kor87a] Bogdan Korel. The program dependence graph in static program testing. Information Processing Letters, 24(2):103–108, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kurtz:1987:HPR

[KOR87b] Stuart A. Kurtz, Michael J.

O’Donnell, and James S. Royer. How to prove representationindependent independence results. Information Processing Letters, 24 (1):5–10, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kotov:1989:APE

[Kot89] V. Kotov. Andrei P. Ershov (1931– 1988). Information Processing Letters, 31(1):1–2, April ??, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kowaltowski:1984:ASM

[KP84] Tomasz Kowaltowski and Antonio Palma. Another solution of the mutual exclusion problem. Information Processing Letters, 19(3):145– 146, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kim:1989:MDC

[KPK89] June Hyoung Kim, Kyu Ho Park, and Myunghwan Kim. A model of distributed control: dependency and uncertainty. Information Processing Letters, 30(2):73–77, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karloff:1988:UTS

[KPS88] Howard J. Karloff, Ramamohan Paturi, and Janos Simon. Universal traversal sequences of length n^{O(log n)} for cliques. Information Processing Letters, 28(5):241–243, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kreowski:1984:NNR

[KR84] Hans-Jörg Kreowski and Grzegorz Rozenberg. Note on node-rewriting graph grammars. Information Processing Letters, 18(1):21–24, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kothari:1985:GAA

[KR85] Suresh C. Kothari and K. V. S. Ramarao. General algorithms for the address calculation of lexicographically ordered tuples. Information Processing Letters, 21(3):113– 116, September 5, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krawczyk:1980:ECM

[Kra80] Tomasz Krawczyk. Error correction by mutational grammars. Information Processing Letters, 11(1):9–15, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kral:1984:SE

[Kra84] J. Kral. On software equations. Information Processing Letters, 19(4): 191–196, November 12, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krentel:1987:NTB

[Kre87] Mark W. Krentel. A note on the transaction backout problem. Information Processing Letters, 24(3): 149–152, February 13, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krichevsky:1982:LEF

[Kri82] R. E. Krichevsky. Letter to the editor: “A family of codes commutatively equivalent to prefix codes” [Inform. Process. Lett. 12 (1981), no. 1, 1–4; MR 82j:94021] by S. Mauceri and A. Restivo. Information Processing Letters, 14(5):238, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [MR81].

Kriauciukas:1986:TPP

[Kri86a] V. Kriaučiukas. Tree-like parse and polynomial subclasses of search problems. Information Processing Letters, 22(1):49–54, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krivanek:1986:HUN

[Kři86b] Mirko Křivánek. Hexagonal unit network—a tool for proving the NP-completeness results of geometric problems. Information Processing Letters, 22(1):37–41, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krishnaprasad:1988:CC

[Kri88a] T. Krishnaprasad. On the computability of circumscription. Information Processing Letters, 27(5): 237–243, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Krivanek:1988:CUP

[Kři88b] Mirko Křivánek. The complexity of ultrametric partitions on graphs. Information Processing Letters, 27(5): 265–270, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

KrusemanAretz:1988:RAP

[Kru88] F. E. J. Kruseman Aretz. On a recursive ascent parser. Information Processing Letters, 29(4):201– 206, November 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kennaway:1987:VAS

[KS87a] J. R. Kennaway and M. R. Sleep. Variable abstraction in O(n log n) space. Information Processing Letters, 24(5):343–349, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kunde:1987:WRC

[KS87b] Manfred Kunde and Horst Steppat. On the worst-case ratio of a compound multiprocessor scheduling algorithm. Information Processing Letters, 25(6):389–396, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

King:1982:OAS

[KST82] K. N. King and Barbara Smith-Thomas. An optimal algorithm for sink-finding. Information Processing Letters, 14(3):109–111, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Koo:1988:EML

[KT88] Richard Koo and Sam Toueg. Effects of message loss on the termination of distributed protocols. Information Processing Letters, 27(4): 181–188, April 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kubale:1987:CSI

[Kub87] Marek Kubale. The complexity of scheduling independent twoprocessor tasks on dedicated processors. Information Processing Letters, 24(3):141–147, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kucera:1982:PCC

[Kuč82] Luděk Kučera. Parallel computation and conflicts in memory access. Information Processing Letters, 14 (2):93–96, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum and addendum [Kuč83].

Kucera:1983:EAP

[Kuč83] Luděk Kučera. Erratum and addendum to: “Parallel computation and conflicts in memory access” [Inform. Process. Lett. 14 (1982), no. 2, 93–96; MR 83g:68038]. Information Processing Letters, 17(2):107, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Kuč82].

Kucera:1989:GSC

[Kuč89] Luděk Kučera. Graphs with small chromatic numbers are easy to color. Information Processing Letters, 30 (5):233–236, March 13, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumar:1989:CCE

[Kum89] Vijay Kumar. Concurrency control on extendible hashing. Information Processing Letters, 31(1):35–41, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kutten:1988:OFT

[Kut88] Shay Kutten. Optimal fault-tolerant distributed construction of a spanning forest. Information Processing Letters, 27(6):299–307, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Konneker:1981:NHD

[KV81] Lloyd K. Konneker and Yaakov L. Varol. A note on heuristics for dynamic organization of data structures. Information Processing Letters, 12(5):213–216, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Korczynski:1982:CCD

[KW82] Waldemar Korczyński and Józef Winkowski. A communication concept for distributed systems. Information Processing Letters, 15(3): 111–114, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Karhumaki:1984:IME

[KW84] Juhani Karhumäki and Derick Wood. Inverse morphic equivalence on languages. Information Processing Letters, 19(5):213–218, November 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lamb:1988:BSE

[Lam88] David Alex Lamb. Benign side effects. Information Processing Letters, 29(6):301–305, December 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laski:1986:ADC

[Las86] Janusz Laski. An algorithm for the derivation of codefinitions in computer programs. Information Processing Letters, 23(2):85–90, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Latteux:1983:LS

[Lat83] Michel Latteux. On a language without star. Information Processing Letters, 16(1):27–30, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laut:1980:SPI

[Lau80] A. Laut. Safe procedural implementations of algebraic types. Information Processing Letters, 11(4–5): 147–151, December 12, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lautemann:1983:BPH

[Lau83] Clemens Lautemann. BPP and the polynomial hierarchy. Information Processing Letters, 17(4):215– 217, November 8, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laumond:1985:EAP

[Lau85] Jean-Paul Laumond. Enumeration of articulation pairs of a planar graph. Information Processing Letters, 21(4):173–179, October 7, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laumond:1987:OGN

[Lau87] Jean-Paul P. Laumond. Obstacle growing in a nonpolygonal world. Information Processing Letters, 25 (1):41–50, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lavault:1989:ANM

[Lav89] Christian Lavault. Average number of messages for distributed leaderfinding in rings of processors. Information Processing Letters, 30(4): 167–176, February 27, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ligozat:1989:RBI

[LB89] Gérard Ligozat and Hélène Bestougeff. On relations between intervals. Information Processing Letters, 32(4):177–182, September 1, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lewis:1981:ECF

[LBB81] Gilbert N. Lewis, Nancy J. Boynton, and F. Warren Burton. Expected complexity of fast search with uniformly distributed data. Information Processing Letters, 13(1):4–7, October 27, 1981. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leszczylowski:1988:LPE

[LBM88] Jacek Leszczylowski, Staffan Bonnier, and Jan Maluszyński. Logic programming with external procedures: introducing S-unification. Information Processing Letters, 27 (3):159–165, March 25, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1985:PGD

[LC85] D. T. Lee and Y. T. Ching. The power of geometric duality revisited. Information Processing Letters, 21(3):117–122, September 5, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liu:1980:ATL

[LD80] Lishing Liu and Alan Demers. An algorithm for testing lossless join property in relational databases. Information Processing Letters, 11(2): 73–76, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lamb:1988:TCM

[LD88] David A. Lamb and Robin Dawes.

Testing for class membership in multi-parent hierarchies. Information Processing Letters, 28(1):21–25, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lea:1988:DHT

[Lea88] Douglas Lea. Digital and Hilbert K-D trees. Information Processing Letters, 27(1):35–41, February 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1989:AAD

[Lee89] De-Lei Lee. On access and alignment of data in a parallel processor. Information Processing Letters, 33(1):11–14, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leiss:1980:NSS

[Lei80] E. Leiss. A note on a signature system based on probabilistic logic. Information Processing Letters, 11(2): 110–113, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leipala:1982:OMI

[Lei82] Timo Leipälä. On optimal multilevel indexed sequential files. Information Processing Letters, 15(5):191– 195, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leiss:1986:ISC

[Lei86] Ernst L. Leiss. The inaccessible set: a classification by query type of security risks in statistical databases. Information Processing Letters, 23 (6):275–279, December 3, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leiss:1988:DDT

[Lei88] Ernst L. Leiss. On the degree of dominator trees. Information Processing Letters, 29(4):199–200, November 12, 1988. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lester:1983:AAQ

[Les83] Lewis Neale Lester. Accuracy of approximating queueing network departure processes with independent renewal processes. Information Processing Letters, 16(1):43–48, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1982:ODR

[Leu82] C. H. C. Leung. Optimal database reorganisation: some practical difficulties. Information Processing Letters, 15(1):23–27, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1984:ASU

[Leu84] Clement H. C. Leung. Approximate storage utilisation of B-trees: a simple derivation and generalisations. Information Processing Letters, 19(4):199–201, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1989:PRP

[Leu89] Joseph Y.-T. Leung. Bin packing with restricted piece sizes. Information Processing Letters, 31(3):145– 149, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levy:1982:ILS

[Lev82] Leon S. Levy. An improved listsearching algorithm. Information Processing Letters, 15(1):43–45, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levcopoulos:1987:LBN

[Lev87] Christos Levcopoulos. An Ω(\sqrt{n}) lower bound for the nonoptimality of the greedy triangulation. Information Processing Letters, 25(4): 247–251, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leland:1981:HDG

[LFQ^{+}81] W. Leland, R. Finkel, Li Qiao, M. Solomon, and L. Uhr. High density graphs for processor interconnection. Information Processing Letters, 12(3):117–120, June 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:1988:STO

[Li88a] Ming Li. A separator theorem for one-dimensional graphs under linear mapping. Information Processing Letters, 27(1):9–11, February 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:1988:RPP

[Li88b] Shuo-Yen Robert Li. Reconstruction of polygons from projections. Information Processing Letters, 28(5): 235–240, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liang:1980:LBL

[Lia80] Frank M. Liang. A lower bound for on-line bin packing. Information Processing Letters, 10(2):76–79, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lickteig:1984:NBR

[Lic84] Thomas Lickteig. A note on border rank. Information Processing Letters, 18(3):173–178, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lickteig:1986:GEO

[Lic86] Thomas Lickteig. Gaussian elimination is optimal for solving linear equations in dimension two. Information Processing Letters, 22(6): 277–279, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lifschitz:1984:VPG

[Lif84] Vladimir Lifschitz. On verification of programs with GOTO statements. Information Processing Letters, 18(4):221–225, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lingas:1986:GDT

[Lin86] Andrzej Lingas. The Greedy and Delauney triangulations are not bad in the average case. Information Processing Letters, 22(1):25–31, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lingas:1989:VDB

[Lin89] Andrzej Lingas. Voronoĭ diagrams with barriers and the shortest diagonal problem. Information Processing Letters, 32(4):191–198, September 1, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lipski:1984:MPA

[Lip84] Witold Lipski, Jr. An O(n log n) Manhattan path algorithm. Information Processing Letters, 19(2):99– 102, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lischke:1981:TTP

[Lis81] Gerhard Lischke. Two types of properties for complexity measures. Information Processing Letters, 12(3): 123–126, June 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leiss:1986:HVB

[LJ86] Ernst L. Leiss and Chamaiporn Jitmedha. Horizontally and vertically bounded propagation of privileges. Information Processing Letters, 22(6):319–327, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lingas:1989:SIN

[LK89] Andrzej Lingas and Marek Karpiński. Subtree isomorphism is NC reducible to bipartite perfect matching. Information Processing Letters, 30(1):27–32, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1984:CCM

[LL84] C. C. Lee and D. T. Lee. On a circlecover minimization problem. Information Processing Letters, 18(2): 109–115, February 28, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lodi:1985:SSH

[LL85] E. Lodi and F. Luccio. Split sequence hash search. Information Processing Letters, 20(3):131–136, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Liskiewicz:1988:ART

[LL88] Maciej Liśkiewicz and Krzysztof Loryś. Alternating real-time computations. Information Processing Letters, 28(6):311–316, August 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Laih:1989:NTS

[LLH89] Chi Sung Laih, Jau Yien Lee, and Lein Harn. A new threshold scheme and its application in designing the conference key distribution cryptosystem. Information Processing Letters, 32(3):95–99, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lloyd:1980:ELS

[Llo80a] E. L. Lloyd. Errata: “List scheduling bounds for UET systems with resources” [Inform. Process. Lett. 10 (1980), no. 1, 28–31; MR 81a:68042]. Information Processing Letters, 11 (1):57, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Llo80b].

Lloyd:1980:LSB

[Llo80b] Errol L. Lloyd. List scheduling bounds for UET systems with resources. Information Processing Letters, 10(1):28–31, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also errata [Llo80a, vW80a].

Lloyd:1981:CGS

[Llo81] Errol L. Lloyd. Coffman–Graham scheduling of UET task systems with 0-1 resources. Information Processing Letters, 12(1):40–45, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lloyd:1983:SDR

[Llo83] Errol L. Lloyd. On a simple deadlock recovery problem. Information Processing Letters, 16(4):175– 178, May 13, 1983. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lloyd:1989:FAF

[Llo89] Errol L. Lloyd. A fast algorithm for finding interlocking sets. Information Processing Letters, 32(1):47– 50, July 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Leung:1980:NPS

[LM80a] Joseph Y.-T. Leung and M. L. Merrill. A note on preemptive scheduling of periodic, real-time tasks. Information Processing Letters, 11(3): 115–118, November 18, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1980:CMC

[LM80b] F. Luccio and S. Mazzone. A cryptosystem for multiple communication. Information Processing Letters, 10(4–5):180–183, July 5, 1980.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See notes [Mei81, Hel81].

Lawler:1981:SPO

[LM81] Eugene L. Lawler and Charles U.

Martel. Scheduling periodically occurring tasks on multiple processors. Information Processing Letters, 12(1):9–12, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lautemann:1985:LTB

[LM85] Clemens Lautemann and Friedhelm Meyer auf der Heide. Lower time bounds for integer programming with two variables. Information Processing Letters, 21(2):101–105, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lakshmanan:1987:TMD

[LMT87] K. B. Lakshmanan, N. Meenakshi, and K. Thulasiraman. A time-optimal message-efficient distributed algorithm for depth-first-search. Information Processing Letters, 25(2):103–109, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loui:1986:ECN

[LMW86] Michael C. Loui, Teresa A. Matsushita, and Douglas B. West. Election in a complete network with a sense of direction. Information Processing Letters, 22(4):185–187, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [LMW88].

Loui:1988:CEC

[LMW88] Michael C. Loui, Teresa A. Matsushita, and Douglas B. West. Corrigendum: “Election in a Complete Network with a Sense of Direction” [Inform. Process. Lett. 22(4), 17 April 1986, pp. 185–187]. Information Processing Letters, 28(6):327, August 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LMW86].

Liu:1988:DPU

[LN88] Robin Liu and Simeon Ntafos. On decomposing polygons into uniformly monotone parts. Information Processing Letters, 27(2):85– 89, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lassez:1982:FPT

[LNS82] J.-L. Lassez, V. L. Nguyen, and E. A. Sonenberg. Fixed point theorems and semantics: a folk tale. Information Processing Letters, 14(3): 112–116, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lentfert:1989:DSR

[LO89] Patrick Lentfert and Mark H. Overmars. Data structures in a real-time environment. Information Processing Letters, 31(3):151–155, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loizou:1980:CFA

[Loi80] G. Loizou. On a cycle finding algorithm. Information Processing Letters, 11(1):33–36, August 29, 1980.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lopriore:1989:SCC

[Lop89] Lanfranco Lopriore. Softwarecontrolled cache coherence protocol for multicache systems. Information Processing Letters, 33(3):125– 130, November 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lotti:1984:ATT

[Lot84] Grazia Lotti. Area-time tradeoff for rectangular matrix multiplication in VLSI models. Information Processing Letters, 19(2):95–98, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Loui:1980:NPG

[Lou80] Michael C. Loui. A note on the pebble game. Information Processing Letters, 11(1):24–26, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1982:LAD

[LP82] F. Luccio and L. Pagli. A linear algorithm to determine minimal spanning forests in chain graphs. Information Processing Letters, 15(1): 1–4, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lifschitz:1983:NCP

[LP83] Vladimir Lifschitz and Leon Pesotchinsky. A note on the complexity of a partition algorithm. Information Processing Letters, 17(3): 117–120, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lee:1989:PBP

[LP89a] D. T. Lee and F. P. Preparata. Parallel batched planar point location on the CCC. Information Processing Letters, 33(4):175–179, December 21, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Levcopoulos:1989:NAP

[LP89b] Christos Levcopoulos and Ola Petersson. A note on adaptive parallel sorting. Information Processing Letters, 33(4):187–191, December 21, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1989:UBR

[LP89c] F. Luccio and L. Pagli. On the upper bound on the rotation distance of binary trees. Information Processing Letters, 31(2):57–60, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luccio:1988:PSP

[LPP88] Fabrizio Luccio, A. Pietracaprina, and G. Pucci. A probabilistic simulation of PRAMs on a bounded degree network. Information Processing Letters, 28(3):141–147, July 4, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luque:1980:TAM

[LR80] E. Luque and A. Ripoll. Tuning architecture via microprogramming. Information Processing Letters, 11 (2):102–109, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Luque:1984:ILP

[LR84] E. Luque and A. Ripoll. Integer linear programming for microprograms register allocation. Information Processing Letters, 19(2):81–85, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lavallee:1986:FDM

[LR86] Ivan Lavallée and Gérard Roucairol.

A fully distributed (minimal) spanning tree algorithm. Information Processing Letters, 23(2):55–62, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lagarias:1982:MSR

[LS82a] J. C. Lagarias and D. E. Swartwout.

Minimal storage representations for binary relations. Information Processing Letters, 14(2):63–66, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lewis:1982:UCC

[LS82b] Harry R. Lewis and Richard Statman. Unifiability is complete for co-NLogSpace. Information Processing Letters, 15(5):220–222, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:1986:NAP

[LS86a] Ten-Hwang H. Lai and Alan Sprague. A note on anomalies in parallel branch-and-bound algorithms with one-to-one bounding functions. Information Processing Letters, 23(3):119–122, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lin:1986:LEL

[LS86b] Ferng-Ching Lin and Wei Kuan Shih. Long edges in the layouts of shuffle-exchange and cubeconnected cycles graphs. Information Processing Letters, 23(1):5–9, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lindhorst:1989:RSC

[LS89] Greg Lindhorst and Farhad Shahrokhi. On renaming a set of clauses as a Horn set. Information Processing Letters, 30(6):289–293, March 28, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lengauer:1980:SCP

[LT80] Thomas Lengauer and Robert E. Tarjan. The space complexity of pebble games on trees. Information Processing Letters, 10(4–5):184–188, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Latteux:1986:FGL

[LT86] M. Latteux and E. Timmerman. Finitely generated ω-languages. Information Processing Letters, 23(4): 171–175, November 8, 1986. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Latteux:1988:BST

[LT88a] M. Latteux and E. Timmerman. Bifaithful starry transductions. Information Processing Letters, 28(1):1– 4, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Louka:1988:DPT

[LT88b] Basile Louka and Maurice Tchuente.

Dynamic programming on twodimensional systolic arrays. Information Processing Letters, 29(2):97– 104, September 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lukaszewicz:1982:UG

[Łuk82] Leon Łukaszewicz. Universal Grammars. Information Processing Letters, 15(2):76–80, September 6, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lumelsky:1985:FCD

[Lum85] Vladimir J. Lumelsky. On fast computation of distance between line segments. Information Processing Letters, 21(2):55–61, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lueker:1982:DSD

[LW82] George S. Lueker and Dan E.

Willard. A data structure for dynamic range queries. Information Processing Letters, 15(5):209–213, December 10, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Li:1986:SMC

[LY86] Ming Li and Yaacov Yesha. Stringmatching cannot be done by a twohead one-way deterministic finite automaton. Information Processing Letters, 22(5):231–235, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lai:1987:DS

[LY87] Ten H. Lai and Tao H. Yang. On distributed snapshots. Information Processing Letters, 25(3):153–158, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Lyon:1988:TML

[Lyo88] Gordon Lyon. A tagless marking that is linear over subtrees. Information Processing Letters, 27(1):23– 28, February 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meijer:1980:DAN

[MA80] Henk Meijer and Selim G. Akl. The design and analysis of a new hybrid sorting algorithm. Information Processing Letters, 10(4–5):213–218, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mirzaian:1985:SMS

[MA85] A. Mirzaian and E. Arjomandi. Selection in x + y and matrices with sorted rows and columns. Information Processing Letters, 20(1):13–17, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Magazine:1983:OIC

[Mag83] Michael J. Magazine. Optimality of intuitive checkpointing policies. Information Processing Letters, 17(2): 63–66, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Magott:1984:PEC

[Mag84] Jan Magott. Performance evaluation of concurrent systems using Petri nets. Information Processing Letters, 18(1):7–13, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Magott:1985:PES

[Mag85] Jan Magott. Performance evaluation of systems of cyclic sequential processes with mutual exclusion using Petri nets. Information Processing Letters, 21(5):229–232, November 18, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Magott:1987:PEC

[Mag87] Jan Magott. Performance evaluation of concurrent systems using conflict-free and persistent Petri nets. Information Processing Letters, 26(2):77–80, October 19, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Main:1982:PCF

[Mai82] Michael G. Main. Permutations are not context-free: an application of the Interchange Lemma. Information Processing Letters, 15(2):68– 71, September 6, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Main:1985:ISC

[Mai85] Michael G. Main. An infinite squarefree co-CFL. Information Processing Letters, 20(2):105–107, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makarov:1981:UDS

[Mak81] O. M. Makarov. Using duality for the synthesis of an optimal algorithm involving matrix multiplication. Information Processing Letters, 13(2):48–49, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1983:BTU

[Mäk83] Erkki Mäkinen. Boundedness testing for unambiguous context-free grammars. Information Processing Letters, 17(4):181–183, November 8, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1984:DPC

[Mäk84] Erkki Mäkinen. On derivation preservation. Information Processing Letters, 19(5):225–228, November 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1985:NUR

[Mäk85a] Erkki Mäkinen. A note on undercover relation. Information Processing Letters, 20(1):19–21, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1985:UPC

[Mäk85b] Erkki Mäkinen. An undecidable problem for context-free grammars.

Information Processing Letters, 20 (3):141–142, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makarov:1986:NAM

[Mak86a] O. M. Makarov. A noncommutative algorithm for multiplying 5 × 5 matrices using 102 multiplications. Information Processing Letters, 23(3): 115–117, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1986:NPG

[Mäk86b] Erkki Mäkinen. A note on pure grammars. Information Processing Letters, 23(5):271–274, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1988:LSH

[Mak88a] Erkki Makinen. On linear search heuristics. Information Processing Letters, 29(1):35–36, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1988:RDB

[Mäk88b] Erkki Mäkinen. On the rotation distance of binary trees. Information Processing Letters, 26(5):271– 272, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Makinen:1989:SIP

[Mäk89] Erkki Mäkinen. On the subtree isomorphism problem for ordered trees. Information Processing Letters, 32 (5):271–273, September 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [Gro91].

Maley:1987:OCC

[Mal87] F. Miller Maley. An observation concerning constraint-based compaction. Information Processing Letters, 25(2):119–122, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manacher:1982:SFA

[Man82] G. K. Manacher. Steady-paced-output and fractional-on-line algorithms on a RAM. Information Processing Letters, 15(2):47– 52, September 6, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Manber:1984:PLB

[Man84] Udi Manber. A probabilistic lower bound for checking disjointness of sets. Information Processing Letters, 19(1):51–53, July 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/0020019084901297.

Manolopoulos:1986:BSI

[Man86] Yannis Manolopoulos. Batched search of index sequential files. Information Processing Letters, 22(5): 267–272, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:1983:PTF

[Mar83] J. J. Martin. Precise typing and filters. Information Processing Letters, 17(3):109–112, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:1985:EPA

[Mar85a] Alain J. Martin. Erratum: “The probe: an addition to communication primitives”. Information Processing Letters, 21(2):107, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Mar85b].

Martin:1985:PAC

[Mar85b] Alain J. Martin. The probe: an addition to communication primitives. Information Processing Letters, 20 (3):125–130, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [Mar85a].

Martel:1986:LBP

[Mar86a] Charles U. Martel. Lower bounds on parallel algorithms for finding the first maximal independent set. Information Processing Letters, 22(2): 81–85, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:1986:NGD

[Mar86b] Alain J. Martin. A new generalization of Dekker’s algorithm for mutual exclusion. Information Processing Letters, 23(6):295–297, December 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:1987:EFM

[Mar87] Ursula Martin. Extension functions for multiset orderings. Information Processing Letters, 26(4):181– 186, December 4, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matousek:1988:LAR

[Mat88] Jiří Matoušek. Line arrangements and range search. Information Processing Letters, 27(6):275–280, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matousek:1989:OCC

[Mat89a] Jiří Matoušek. On-line computation of convolutions. Information Processing Letters, 32(1):15–16, July 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mattern:1989:EDT

[Mat89b] Friedemann Mattern. An efficient distributed termination test. Information Processing Letters, 31(4): 203–208, May 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mattern:1989:GQD

[Mat89c] Friedemann Mattern. Global quiescence detection based on credit distribution and recovery. Information Processing Letters, 30(4):195– 200, February 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mayoh:1983:MPP

[May83] Brian H. Mayoh. Models of programs and processes. Information Processing Letters, 17(4):211–214, November 8, 1983. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mazurkiewicz:1988:SAR

[Maz88] Antoni Mazurkiewicz. Solvability of the asynchronous ranking problem. Information Processing Letters, 28 (5):221–224, August 12, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martin:1985:FME

[MB85] Alain J. Martin and Jerry R. Burch.

Fair mutual exclusion with unfair P and V operations. Information Processing Letters, 21(2):97–100, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Majster-Cederbaum:1980:SRB

[MC80] Mila E. Majster-Cederbaum. A simple relation between relational and predicate transformer semantics for nondeterministic programs. Information Processing Letters, 11(4–5): 190–192, December 12, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Majster-Cederbaum:1988:UFP

[MC88] Mila E. Majster-Cederbaum. On the uniqueness of fixed points of endofunctors in a category of complete metric spaces. Information Processing Letters, 29(6):277–281, December 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Majster-Cederbaum:1989:CPS

[MC89a] Mila E. Majster-Cederbaum. The contraction property is sufficient to guarantee the uniqueness of fixed points of endofunctors in a category of complete metric spaces. Information Processing Letters, 33(1): 15–19, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

MixBarrington:1989:RCS

[MC89b] David A. Mix Barrington and James Corbett. On the relative complexity of some languages in NC^{1}. Information Processing Letters, 32(5):251– 256, September 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McCune:1988:SCS

[McC88] William W. McCune. Un-Skolemizing clause sets. Information Processing Letters, 29(5):257– 263, November 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McCurley:1989:AVP

[McC89] E. Robert McCurley. Auxiliary variables in partial correctness programming logics. Information Processing Letters, 33(3):131–133, November 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McGlinn:1983:SBT

[McG83] Robert J. McGlinn. Is SSS^{∗} better than alpha-beta. Information Processing Letters, 16(3):113–120, April 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McKenzie:1984:PBD

[McK84] Pierre McKenzie. Permutations of bounded degree generate groups of polynomial diameter. Information Processing Letters, 19(5):253–254, November 26, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McKeown:1985:SPM

[McK85] G. P. McKeown. A special purpose MIMD parallel processor. Information Processing Letters, 20(1):23–27, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McLean:1985:CBS

[McL85] John McLean. A comment on the ‘basic security theorem’ of Bell and LaPadula. Information Processing Letters, 20(2):67–70, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mehlhorn:1988:FAA

[Meh88] Kurt Mehlhorn. A faster approximation algorithm for the Steiner problem in graphs. Information Processing Letters, 27(3):125–128, March 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [Flo91].

Meijer:1981:NCM

[Mei81] H. Meijer. A note on “A cryptosystem for multiple communication” [Inform. Process. Lett. 10(4– 5), 5 July 1980, pp. 180–183]. Information Processing Letters, 12(4): 179–181, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [LM80b, Hel81].

Meira:1985:LAS

[Mei85] Silvio Lemos Meira. A linear applicative solution for the set union problem. Information Processing Letters, 20(1):43–45, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Melkman:1987:LCC

[Mel87] Avraham A. Melkman. On-line construction of the convex hull of a simple polyline. Information Processing Letters, 25(1):11–12, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Metivier:1983:ARS

[Mét83] Yves Métivier. About the rewriting systems produced by the Knuth Bendix completion algorithm. Information Processing Letters, 16(1): 31–34, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

MeyeraufderHeide:1983:ICC

[Mey83] Friedhelm Meyer Auf Der Heide. Infinite cube-connected cycles. Information Processing Letters, 16(1): 1–2, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Meyer:1985:ISM

[Mey85] Bertrand Meyer. Incremental string matching. Information Processing Letters, 21(5):219–227, November 18, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mauney:1981:IIE

[MF81] J. Mauney and C. N. Fischer. An improvement to immediate error detection in Strong LL(1) parsers. Information Processing Letters, 12(5): 211–212, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Madduri:1984:EBA

[MF84] Hari Madduri and Raphael Finkel.

Extension of the Banker’s algorithm for resource allocation in a distributed operating system. Information Processing Letters, 19 (1):1–8, July 26, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/0020019084901194.

Melville:1980:CDS

[MG80] Robert Melville and David Gries.

Controlled density sorting. Information Processing Letters, 10(4–5): 169–172, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Martel:1989:FPQ

[MG89] Charles U. Martel and Dan Gusfield. A fast parallel quicksort algorithm. Information Processing Letters, 30(2):97–102, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miller:1988:CIJ

[MGKL88] Leslie L. Miller, S. K. Gadia, S. Kothari, and K. C. Liu. Completeness issues for join dependencies derived from the universal relation join dependency. Information Processing Letters, 28(5):269– 274, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Micali:1981:TWD

[Mic81] Silvio Micali. Two-way deterministic finite automata are exponentially more succinct than sweeping automata. Information Processing Letters, 12(2):103–105, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mihail:1989:CAP

[Mih89] Milena Mihail. On coupling and the approximation of the permanent. Information Processing Letters, 30(2):91–95, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mili:1983:VPI

[Mil83] Ali Mili. Verifying programs by induction on their data structure: General format and applications. Information Processing Letters, 17 (3):155–160, October 5, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minoux:1988:LSL

[Min88] Michel Minoux. LTUR: a simplified linear-time unit resolution algorithm for Horn formulae and computer implementation. Information Processing Letters, 29(1):1–12, September 15, 1988. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miola:1984:AAA

[Mio84] Alfonso Miola. Algebraic approach to p-adic conversion of rational numbers. Information Processing Letters, 18(3):167–171, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mirzaian:1987:HTL

[Mir87] Andranik Mirzaian. A halving technique for the longest stuttering subsequence problem. Information Processing Letters, 26(2):71–75, October 19, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mishra:1988:SAU

[Mis88] Prateek Mishra. Strictness analysis of the untyped λ-calculus. Information Processing Letters, 28(3):121– 125, July 4, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Misra:1989:SPS

[Mis89] Jayadev Misra. A simple proof of a simple consensus algorithm. Information Processing Letters, 33(1): 21–24, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mittal:1988:FBA

[Mit88] Hari Ballabh Mittal. A fast backtrack algorithm for graph isomorphism. Information Processing Letters, 29(2):105–110, September 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miyano:1984:RTW

[Miy84] Satoru Miyano. Remarks on twoway automata with weak-counters. Information Processing Letters, 18 (2):105–107, February 28, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Miyano:1988:PLF

[Miy88] Satoru Miyano. A parallelizable lexicographically first maximal edgeinduced subgraph problem. Information Processing Letters, 27(2):75– 78, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Matsuoka:1989:AEW

[MKK89] Satoshi Matsuoka, Tomihisa Kamada, and Satoru Kawai. Asymptotic evaluation of window visibility. Information Processing Letters, 31 (3):119–126, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mannila:1985:FAR

[MM85] Heikki Mannila and Kurt Mehlhorn. A fast algorithm for renaming a set of clauses as a Horn set. Information Processing Letters, 21(5):269– 272, November 18, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maluszynski:1982:GU

[MN82] Jan Małuszyński and Jörgen Fischer Nilsson. Grammatical unification. Information Processing Letters, 15 (4):150–158, October 31, 1982. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moller-Nielsen:1984:EFS

[MNS84] Peter Moller-Nielsen and Jorgen Staunstrup. Experiments with a fast string searching algorithm. Information Processing Letters, 18(3): 129–135, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mehlhorn:1981:ODD

[MO81] Kurt Mehlhorn and Mark H. Overmars. Optimal dynamization of decomposable searching problems. Information Processing Letters, 12(2): 93–98, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Madlener:1987:USR

[MO87a] Klaus Madlener and Friedrich Otto.

Using string-rewriting for solving the word problem for finitely presented groups. Information Processing Letters, 24(5):281–284, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Metivier:1987:LSC

[MO87b] Y. Métivier and E. Ochmanski. On lexicographic semi-commutations. Information Processing Letters, 26 (2):55–59, October 19, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Monien:1984:DTW

[Mon84] B. Monien. Deterministic two-way one-head pushdown automata are very powerful. Information Processing Letters, 18(5):239–242, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morris:1979:SSM

[Mor79] J. M. Morris. A starvation-free solution to the mutual exclusion problem. Information Processing Letters, 8(2):76–80, February 15, 1979. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [AJV86].

Moran:1981:NSP

[Mor81] Shlomo Moran. A note on: “Shortest-path problem is not harder than matrix multiplication” [Inform. Process. Lett. 11 (1980), no. 3, 134–136; MR 81k:68036] by F. Romani. With a reply by Romani. Information Processing Letters, 13 (2):85–86, November 13, 1981. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Rom80].

Morris:1982:ACG

[Mor82] F. L. Morris. Another compacting garbage collector. Information Processing Letters, 15(4):139–142, October 31, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morgan:1985:GLT

[Mor85] Carroll Morgan. Global and logical time in distributed algorithms. Information Processing Letters, 20(4): 189–194, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moran:1987:GLB

[Mor87a] Shlomo Moran. Generalized lower bounds derived from Håstad’s main lemma (small depth circuits). Information Processing Letters, 25(6): 383–388, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morris:1987:VWL

[Mor87b] Joseph M. Morris. Varieties of weakest liberal preconditions. Information Processing Letters, 25(3):207– 210, May 29, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morgan:1988:AVD

[Mor88a] Carroll Morgan. Auxiliary variables in data refinement. Information Processing Letters, 29(6):293– 296, December 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morgan:1988:DRM

[Mor88b] Carroll C. Morgan. Data refinement by miracles. Information Processing Letters, 26(5):243–246, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moriya:1988:CPP

[Mor88c] Tetsuo Moriya. Closure property of principal cones under substitution. Information Processing Letters, 29 (6):315–317, December 8, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morrison:1988:PAC

[Mor88d] John F. Morrison. Parallel p-adic computation. Information Processing Letters, 28(3):137–140, July 4, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Morris:1989:WFI

[Mor89] Joseph M. Morris. Well-founded induction and the invariance theorem for loops. Information Processing Letters, 32(3):155–158, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mostowski:1982:DSA

[Mos82] A. Włodzimierz Mostowski. Determinancy of sinking automata on infinite trees and inequalities between various Rabin’s pair indices. Information Processing Letters, 15(4): 159–163, October 31, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Motoki:1982:NUB

[Mot82] Tatsuya Motoki. A note on upper bounds for the selection problem. Information Processing Letters, 15(5):214–219, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moura:1980:NGC

[Mou80] Arnaldo Moura. A note on grammatical covers. Information Processing Letters, 11(3):127–129, November 18, 1980. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McColl:1987:PRB

[MP87] W. F. McColl and M. S. Paterson. The planar realization of Boolean functions. Information Processing Letters, 24(3):165–170, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maggs:1988:MCS

[MP88] Bruce M. Maggs and Serge A.

Plotkin. Minimum-cost spanning tree as a path-finding problem. Information Processing Letters, 26(6): 291–293, January 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mathieu:1987:AED

[MPY87] C. Mathieu, Claude Puech, and Hossein Yahia. Average efficiency of data structures for binary image processing. Information Processing Letters, 26(2):89–93, October 19, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mauceri:1981:FCC

[MR81] S. Mauceri and A. Restivo. A family of codes commutatively equivalent to prefix codes. Information Processing Letters, 12(1):1–4, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See letter to the editor [Kri82].

Martin:1984:PFA

[MR84] Alain J. Martin and Martin Rem. A presentation of the Fibonacci algorithm. Information Processing Letters, 19(2):67–68, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

McKeown:1984:CPM

[MRS84] G. P. McKeown and V. J. Rayward-Smith. Communication problems on MIMD parallel computers. Information Processing Letters, 19(2):69–73, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maier:1981:HB

[MS81] D. Maier and S. C. Salveter. Hysterical B-trees. Information Processing Letters, 12(4):199–202, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Murphy:1986:EUT

[MS86] O. J. Murphy and S. M. Selkow. The efficiency of using k-d trees for finding nearest neighbors in discrete space. Information Processing Letters, 23(4):215–218, November 8, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Marchetti-Spaccamela:1985:DAC

[MSR85] A. Marchetti-Spaccamela and G. Romano. On different approximation criteria for subset product problems. Information Processing Letters, 21 (4):213–218, October 7, 1985. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Maio:1984:EAC

[MST84] D. Maio, M. R. Scalas, and P. Tiberio. On estimating access costs in relational databases. Information Processing Letters, 19(3): 157–161, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moffat:1984:PQA

[MT84] Alistair Moffat and Tadao Takaoka. A priority queue for the all pairs shortest path problem. Information Processing Letters, 18(4):189– 193, May 14, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mints:1986:SDL

[MT86] G. Mints and E. Tyugu. Semantics of a declarative language. Information Processing Letters, 23(3): 147–151, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [MT87b].

Marie:1987:NEP

[MT87a] Raymond Marie and Kishor S.

Trivedi. A note on the effect of preemptive policies on the stability of a priority queue. Information Processing Letters, 24(6):397–401, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mints:1987:CSD

[MT87b] G. Mints and E. Tyugu. Corrigendum: “Semantics of a declarative language” [Inform. Process. Lett. 23(3), 22 October 1997, pp. 147– 151]. Information Processing Letters, 25(2):139, May 6, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [MT86].

Malhotra:1989:IDB

[MTK89] Vishv Mohan Malhotra, Tang Van To, and Kanchana Kanchanasut. An improved data-dependency-based backtracking scheme for Prolog. Information Processing Letters, 31(4): 185–189, May 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mannila:1984:SLA

[MU84] Heikki Mannila and Esko Ukkonen. A simple linear-time algorithm for in situ merging. Information Processing Letters, 18(4):203–208, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Mukkamala:1989:SPV

[Muk89] Ravi Mukkamala. Some properties of view-based replication control algorithms for distributed systems. Information Processing Letters, 31(6):295–298, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Muller:1987:SNU

[Mül87] Heinrich Müller. Sorting numbers using limited systolic coprocessors. Information Processing Letters, 24 (6):351–354, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Murtagh:1983:ETC

[Mur83] F. Murtagh. Expected-time complexity results for hierarchic clustering algorithms which use cluster centres. Information Processing Letters, 16(5):237–241, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Moran:1987:EIR

[MW87] Shlomo Moran and Yaron Wolfstahl. Extended impossibility results for asynchronous complete networks. Information Processing Letters, 26(3):145–151, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Myers:1983:ARA

[Mye83] Eugene W. Myers. An applicative random-access stack. Information Processing Letters, 17(5):241– 248, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Minker:1982:ELR

[MZ82] Jack Minker and Guy Zanon. An extension to linear resolution with selection function. Information Processing Letters, 14(4):191–194, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakamura:1981:AIP

[NA81] Akira Nakamura and Kunio Aizawa.

Acceptors for isometric parallel context-free array languages. Information Processing Letters, 13(4–5): 182–186, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nazareth:1989:ANM

[NA89] J. L. Nazareth and K. A. Ariyawansa.

On accelerating Newton’s method based on a conic model. Information Processing Letters, 30(6):277–281, March 28, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakamura:1981:SRO

[Nak81] Akira Nakamura. Some remarks on one-pebble rectangular array acceptors. Information Processing Letters, 13(2):80–84, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Naor:1987:FPC

[Nao87] Joseph Naor. A fast parallel coloring of planar graphs with five colors. Information Processing Letters, 25 (1):51–53, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Narasimhan:1989:NHC

[Nar89] Giri N. Narasimhan. A note on the Hamiltonian circuit problem on directed path graphs. Information Processing Letters, 32(4):167– 170, September 1, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nawrocki:1983:CSL

[Naw83] Jerzy R. Nawrocki. Contiguous segmentation with limited compacting. Information Processing Letters, 17 (2):57–62, August 24, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nguyen:1985:IMC

[Ngu85] Van Nguyen. The incompleteness of Misra and Chandy’s proof systems. Information Processing Letters, 21(2):93–96, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakamura:1980:RTD

[NI80] Akira Nakamura and Katsushi Inoue. A remark on two-dimensional finite automata: “Some properties of two-dimensional on-line tessellation acceptors” [Inform. Sci. 13 (1977), no. 2, 95–121; MR 80e:68143]. Information Processing Letters, 10(4–5):219–222, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nishihara:1983:ACA

[NI83] Seiichi Nishihara and Katsuo Ikeda.

Address calculation algorithms for ordered sets of combinations. Information Processing Letters, 17(5): 251–253, December 15, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nagamochi:1989:MFM

[NI89] Hiroshi Nagamochi and Toshihide Ibaraki. On max-flow min-cut and integral flow properties for multicommodity flows in directed networks. Information Processing Letters, 31(6):279–285, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nijholt:1982:NSS

[Nij82a] Anton Nijholt. A note on the sufficiency of Sokołowski’s criterion for context-free languages. Information Processing Letters, 14(5):207, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nijholt:1982:RBG

[Nij82b] Anton Nijholt. On the relationship between the LL(k) and LR(k) grammars. Information Processing Letters, 15(3):97–101, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Njau:1985:DDC

[Nja85] Ernest C. Njau. Details of distortions in the computed Fourier transforms of signals. Part I. Short periodic signals. Information Processing Letters, 20(3):111–113, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nath:1982:PAC

[NM82] Dhruva Nath and S. N. Maheshwari. Parallel algorithms for the connected components and minimal spanning tree problems. Information Processing Letters, 14(1):7–11, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nawrocki:1987:SAM

[NM87] Jerzy R. Nawrocki and J. Martinek. A storage allocation method with invalidating dangling references. Information Processing Letters, 25(5): 305–310, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Narendran:1988:PUT

[NO88] Paliath Narendran and Friedrich Otto. Preperfectness is undecidable for Thue systems containing only length-reducing rules and a single commutation rule. Information Processing Letters, 29(3):125–130, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Noshita:1985:TTC

[Nos85] Kohei Noshita. Translation of Turner combinators in O(n log n) space. Information Processing Letters, 20(2):71–74, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Negri:1985:JDM

[NP85] M. Negri and G. Pelagatti. Join during merge: an improved sort based algorithm. Information Processing Letters, 21(1):11–16, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nakata:1986:LLL

[NS86] Ikuo Nakata and Masataka Sassa. Lattributed LL(1)-grammars are LRattributed. Information Processing Letters, 23(6):325–328, December 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ntafos:1986:GWG

[Nta86] Simeon Ntafos. On gallery watchmen in grids. Information Processing Letters, 23(2):99–102, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Nawrocki:1988:FBO

[NU88] Jerzy R. Nawrocki and Andrzej Urbanski. Fixed-sized blocks optimization. Information Processing Letters, 29(4):165–169, November 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ozguner:1988:RAF

[ÖA88] Füsun Özgüner and C. Aykanat. A reconfiguration algorithm for fault tolerance in a hypercube multiprocessor. Information Processing Letters, 29(5):247–254, November 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ochmanski:1987:ICS

[Och87] Edward Ochmanski. Inevitability in concurrent systems. Information Processing Letters, 25(4):221– 225, June 17, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

ODunlaing:1988:TLB

[Ó’D88] Colm Ó’Dúnlaing. A tight lower bound for the complexity of pathplanning for a disc. Information Processing Letters, 28(4):165–170, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ohya:1984:FVA

[OIM84] T. Ohya, M. Iri, and K. Murota. A fast Voronoi-diagram algorithm with quaternary tree bucketing. Information Processing Letters, 18(4): 227–231, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olariu:1988:UCP

[Ola88a] Stephan Olariu. On the unimodality of convex polygons. Information Processing Letters, 29(6):289– 292, December 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olariu:1988:PFG

[Ola88b] Stephan Olariu. Paw-free graphs. Information Processing Letters, 28 (1):53–54, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olariu:1989:SLT

[Ola89] Stephan Olariu. A simple lineartime algorithm for computing the RNG and MST of unimodal polygons. Information Processing Letters, 31(5):243–247, June 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olivie:1980:RBS

[Oli80] H. Olivié. On the relationship between son-trees and symmetric binary B-trees. Information Processing Letters, 10(1):4–8, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oh:1985:NRL

[OP85] Se Man Oh and J. C. H. Park. A note on removing loops from tabledriven code generators. Information Processing Letters, 21(1):31–34, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

OHallaron:1986:GDP

[OR86] David R. O’Hallaron and Paul F.

Reynolds, Jr. A generalized deadlock predicate. Information Processing Letters, 23(4):181–188, November 8, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Olariu:1989:WPO

[OR89] S. Olariu and J. Randall. Welsh-Powell opposition graphs. Information Processing Letters, 31(1):43–46, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Orenstein:1982:MTU

[Ore82] Jack A. Orenstein. Multidimensional tries used for associative searching. Information Processing Letters, 14(4):150–157, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Orlowska:1988:PSW

[Orł88] Ewa Orłowska. Proof system for weakest prespecification. Information Processing Letters, 27(6):309– 313, May 13, 1988. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ouksel:1988:IDS

[OS88] Mohamed Ouksel and Peter Scheuermann. Implicit data structures for linear hashing schemes. Information Processing Letters, 29(4):183– 189, November 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ottmann:1981:SGF

[OSW81] T. Ottmann, A. Salomaa, and D. Wood. Sub-regular grammar forms. Information Processing Letters, 12(4):184–187, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Omiecinski:1989:HBJ

[OT89] Edward Omiecinski and Eileen Tien. A hash-based join algorithm for a cube-connected parallel computer. Information Processing Letters, 30 (5):269–275, March 13, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Overmars:1980:FCB

[Ov80] Mark H. Overmars and Jan van Leeuwen. Further comments on A. Bykat’s convex hull algorithm: “Convex hull of a finite set of points in two dimensions” [Inform. Process. Lett. 7 (1978), no. 6, 296–298; MR 80b:68041]. Information Processing Letters, 10(4–5):209–212, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Byk78, Fou79, DS79].

Overmars:1981:WOI

[Ov81a] M. H. Overmars and J. van Leeuwen. Worst-case optimal insertion and deletion methods for decomposable searching problems. Information Processing Letters, 12(4): 168–173, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Overmars:1981:SPD

[Ov81b] Mark H. Overmars and Jan van Leeuwen. Some principles for dynamizing decomposable searching problems. Information Processing Letters, 12(1):49–53 (or 49–54??), February 13, 1981. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Overmars:1981:GMA

[Ove81] Mark H. Overmars. General methods for “all elements” and “all pairs” problems. Information Processing Letters, 12(2):99–102, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Oyamaguchi:1984:SRS

[Oya84] Michio Oyamaguchi. Some remarks on subclass containment problems for several classes of DPDA’s. Information Processing Letters, 19 (1):9–12, July 26, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/0020019084901200.

Ozawa:1983:CSM

[Oza83] K. Ozawa. Considerations on the similarity measures between index terms. Information Processing Letters, 16(5):243–246, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pagan:1980:GCL

[Pag80] Frank G. Pagan. On the generation of compilers from language definitions. Information Processing Letters, 10(2):104–107, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pagli:1985:SAH

[Pag85] Linda Pagli. Self-adjusting hash tables. Information Processing Letters, 21(1):23–25, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pajunen:1980:TTL

[Paj80] Seppo Pajunen. On two theorems of Lenstra. Information Processing Letters, 11(4–5):224–228, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pallo:1987:RDL

[Pal87] Jean Pallo. On the rotation distance in the lattice of binary trees. Information Processing Letters, 25(6): 369–373, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pan:1981:LBA

[Pan81a] V. Ya. Pan. The lower bounds on the additive complexity of bilinear problems in terms of some algebraic quantities. Information Processing Letters, 13(2):71–72, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pansiot:1981:MSI

[Pan81b] J. J. Pansiot. The Morse sequence and iterated morphisms. Information Processing Letters, 12(2):68–70, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pansiot:1981:NPC

[Pan81c] J.-J. Pansiot. A note on Post’s correspondence problem. Information Processing Letters, 12(5):233, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pan:1986:TBA

[Pan86a] Victor Ya. Pan. The trade-off between the additive complexity and the asynchronicity of linear and bilinear algorithms. Information Processing Letters, 22(1):11–14, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Panny:1986:NHM

[Pan86b] Wolfgang Panny. A note on the higher moments of the expected behavior of straight insertion sort. Information Processing Letters, 22(4): 175–177, April 17, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pan:1988:CDC

[Pan88] Victor Pan. Computing the determinant and the characteristic polynomial of a matrix via solving linear systems of equations. Information Processing Letters, 28(2):71–75, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Papadimitriou:1985:ASP

[Pap85] Christos H. Papadimitriou. An algorithm for shortest-path motion in three dimensions. Information Processing Letters, 20(5):259–263, June 12, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parker:1984:CCS

[Par84] J. R. Parker. On converting character strings to integers. Information Processing Letters, 19 (1):17–19, July 26, 1984. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/0020019084901224.

Parberry:1986:RRI

[Par86] Ian Parberry. On recurrent and recursive interconnection patterns. Information Processing Letters, 22(6):

285–289, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parberry:1987:ISS

[Par87] Ian Parberry. An improved simulation of space and reversal bounded deterministic Turing machines by width and depth bounded uniform circuits. Information Processing Letters, 24(6):363–367, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Passy:1980:SPT

[Pas80] Solomon Passy. Structured programs for Turing machines. Information Processing Letters, 10(2): 63–67, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peper:1987:DCC

[Pep87] F. Peper. Determining connected components in linear time by a linear number of processors. Information Processing Letters, 25(6):401– 406, July 26, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Perko:1983:RDS

[Per83] A. Perko. A representation of disjoint sets with fast initialization. Information Processing Letters, 16(1): 21, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peterson:1980:NCP

[Pet80a] J. L. Peterson. A note on colored Petri nets. Information Processing Letters, 11(1):40–43, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pettorossi:1980:DAC

[Pet80b] Alberto Pettorossi. Derivation of an O(k^{2} log n) algorithm for computing order-k Fibonacci numbers from the O(k^{3} log n) matrix multiplication method. Information Processing Letters, 11(4–5):172–179, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Peterson:1981:MAM

[Pet81] G. L. Peterson. Myths about the mutual exclusion problem. Information Processing Letters, 12(3):115– 116, June 13, 1981. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pettersson:1991:LEC

[Pet91] Jimmi S. Pettersson. Letter to the editor: Comments on “Alwaystrue is not invariant”: Assertional reasoning about invariance [Inform. Process. Lett. 35(6), 15 September 1990, pp. 277–279]. Information Processing Letters, 40(5):231– 233, December 13, 1991. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ver87, VT90].

Pujari:1987:CWC

[PG87] A. K. Pujari and S. Gupta. Comment on: “Worst-case choice for the stable marriage problem” [Inform. Process. Lett. 21 (1985), no. 1, 27– 30; MR 87b:68081] by D. Kapur and M. S. Krishnamoorthy. Information Processing Letters, 24(2):139, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [KK85].

Pawagi:1987:CDP

[PGR87a] Shaunak R. Pawagi, P. S. Gopalakrishnan, and I. V. Ramakrishnan. Computing dominators in parallel. Information Processing Letters, 24 (4):217–221, March 2, 1987. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigendum [PGR87b].

Pawagi:1987:CCD

[PGR87b] Shaunak R. Pawagi, P. S. Gopalakrishnan, and I. V. Ramakrishnan. Corrigendum: “Computing dominators in parallel”. Information Processing Letters, 25(6):417, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [PGR87a].

Pitt:1987:NEK

[Pit87] Leonard Pitt. A note on extending Knuth’s tree estimator to directed acyclic graphs. Information Processing Letters, 24(3):203–206, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pramanik:1988:HTH

[PK88] Sakti Pramanik and Myoung Ho Kim. HCB tree a height compressed B tree for parallel processing. Information Processing Letters, 29(4): 213–220, November 12, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Plaisted:1985:USE

[Pla85] David A. Plaisted. The undecidability of self-embedding for term rewriting systems. Information Processing Letters, 20(2):61–64, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Palvia:1984:ABA

[PM84] Prashant Palvia and Salvatore T. March. Approximating block accesses in database organizations. Information Processing Letters, 19(2): 75–79, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Platt:1988:ACU

[PM88] Don Platt and Moneeb A. Magdy. Adaptive control using switched capacitor filters. Information Processing Letters, 28(5):231–234, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Poblete:1986:AFT

[Pob86] Patricio V. Poblete. Approximating functions by their Poisson transform. Information Processing Letters, 23(3):127–130, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Paredaens:1980:GLA

[PP80] J. Paredaens and F. Ponsaert. Grant levels in an authorization mechanism. Information Processing Letters, 11(4–5):152–155, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Parker:1989:CSN

[PP89] Bruce Parker and Ian Parberry.

Constructing sorting networks from k-sorters. Information Processing Letters, 33(3):157–162, November 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Park:1988:EBE

[PPK88] Chan-Ik Park, Kyu Ho Park, and Myunghwan Kim. Efficient backward execution in AND/OR process model. Information Processing Letters, 29(4):191–198, November 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pawagi:1986:APU

[PR86] Shaunak Pawagi and I. V. Ramakrishnan. An O(log n) algorithm for parallel update of minimum spanning trees. Information Processing Letters, 22(5):223–229, April ??, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Provan:1989:SEW

[Pro89] J. Scott Provan. Shortest enclosing walks and cycles in embedded graphs. Information Processing Letters, 30(3):119–125, February 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Petreschi:1980:SAS

[PS80] Rossella Petreschi and Bruno Simeone. A switching algorithm for the solution of quadratic Boolean equations. Information Processing Letters, 11(4–5):193–198, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [PS81a].

Petreschi:1981:ESA

[PS81a] R. Petreschi and B. Simeone. Erratum: “A switching algorithm for the solution of quadratic Boolean equations”. Information Processing Letters, 12(2):109, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [PS80].

Preparata:1981:TSP

[PS81b] Franco P. Preparata and Kenneth J. Supowit. Testing a simple polygon for monotonicity. Information Processing Letters, 12(4):161–164, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Patterson:1988:FVN

[PS88] Nicholas J. Patterson and Kenneth J. Supowit. Finding the vertices nearest to a point in a hypercube. Information Processing Letters, 27(2):99–102, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Padberg:1989:CMB

[PS89] Manfred Padberg and Antonio Sassano. The complexity of matching with bonds. Information Processing Letters, 32(6):297–300, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Passy:1985:PDC

[PT85] Solomon Passy and Tinko Tinchev.

PDL with data constants. Information Processing Letters, 20(1):35–41, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Preparata:1980:AOV

[PV80] Franco P. Preparata and Jean E.

Vuillemin. Area-time optimal VLSI networks for multiplying matrices. Information Processing Letters, 11 (2):77–80, October ??, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1982:NHT

[pW82] Patrick Shen pei Wang. A new hierarchy of two-dimensional array languages. Information Processing Letters, 15(5):223–226, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Pinter:1987:ETT

[PW87] Shlomit S. Pinter and Yaron Wolfstahl. Embedding ternary trees in VLSI arrays. Information Processing Letters, 26(4):187–191, December 4, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Papadimitriou:1981:CPP

[PY81a] Christos H. Papadimitriou and Mihalis Yannakakis. The clique problem for planar graphs. Information Processing Letters, 13(4–5):131–133, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Papadimitriou:1981:MEG

[PY81b] Christos H. Papadimitriou and Mihalis Yannakakis. On minimal Eulerian graphs. Information Processing Letters, 12(4):203–205, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rangarajan:1985:FDE

[RAK85] K. Rangarajan and S. Arun-Kumar. Fair derivations in E0L systems. Information Processing Letters, 20(4): 183–188, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramanan:1987:OLB

[Ram87] Prakash Ramanan. Obtaining lower bounds using artificial components (fixed order algebraic decision tree model). Information Processing Letters, 24(4):243–246, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramakrishna:1989:ARP

[Ram89a] M. V. Ramakrishna. Analysis of random probing hashing. Information Processing Letters, 31(2):83–90, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramanan:1989:ACA

[Ram89b] Prakash Ramanan. Average-case analysis of the smart next fit algorithm. Information Processing Letters, 31(5):221–225, June 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rana:1983:DSD

[Ran83] S. P. Rana. A distributed solution of the distributed termination problem. Information Processing Letters, 17(1):43–46, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raoult:1988:POP

[Rao88] Jean-Claude Raoult. Proving open properties by induction. Information Processing Letters, 29(1):19– 23, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raynal:1987:DAP

[Ray87] Michel Raynal. A distributed algorithm to prevent mutual drift between n logical clocks. Information Processing Letters, 24(3):199– 202, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raymond:1989:DAM

[Ray89a] Kerry Raymond. A distributed algorithm for multiple entries to a critical section. Information Processing Letters, 30(4):189–193, February 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See note [SM92].

Raynal:1989:PNT

[Ray89b] Michel Raynal. Prime numbers as a tool to design distributed algorithms. Information Processing Letters, 33(1):53–58, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ridjanovic:1982:DDD

[RB82] Dzenan Ridjanovic and Michael L. Brodie. Defining database dynamics with attribute grammars. Information Processing Letters, 14(3): 132–138, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reddy:1982:RCC

[RBP82a] P. G. Reddy, S. Bhalla, and B. E. Prasad. Robust, centralized certifier based concurrency control for distributed databases. Information Processing Letters, 15(3):105– 110, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reddy:1982:MCC

[RBP82b] P. G. Reddy, Subhash Bhalla, and B. E. Prasad. A model of concurrency control in distributed database systems. Information Processing Letters, 14(5):208–213, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roussille:1987:GCP

[RD87] M. Roussille and P. Dufour. Generation of convex polygons with individual angular constraints. Information Processing Letters, 24(3):159– 164, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramarao:1988:MCS

[RDM88] K. V. S. Ramarao, Robert Daley, and Rami Melhem. Message complexity of the set intersection problem. Information Processing Letters, 27(4):169–174, April 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Regnier:1981:AHT

[Reg81] Mireille Regnier. On the average height of trees in digital search and dynamic hashing. Information Processing Letters, 13(2):64– 66, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reif:1985:DFS

[Rei85] John H. Reif. Depth-first search is inherently sequential. Information Processing Letters, 20(5):229– 234, June 12, 1985. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reichling:1987:SSQ

[Rei87a] Matthias Reichling. A simplified solution of the N queens’ problem. Information Processing Letters, 25(4): 253–255, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reif:1987:TAD

[Rei87b] John H. Reif. A topological approach to dynamic graph connectivity. Information Processing Letters, 25(1):65–70, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Reichling:1988:DCI

[Rei88] Matthias Reichling. On the detection of a common intersection of k convex objects in the plane. Information Processing Letters, 29(1):25– 29, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Remmers:1984:TDL

[Rem84] J. H. Remmers. A technique for developing loop invariants. Information Processing Letters, 18(3):137– 139, March 30, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rolim:1988:ICA

[RG88] José D. P. Rolim and Sheila A. Greibach. On the IO-complexity and approximation languages. Information Processing Letters, 28(1): 27–31, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Richardson:1989:IAV

[RG89a] Stephen Richardson and Mahadevan Ganapathi. Interprocedural analysis vs. procedure integration. Information Processing Letters, 32(3): 137–142, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rolim:1989:NBC

[RG89b] José D. P. Rolim and Sheila A. Greibach. A note on the best-case complexity. Information Processing Letters, 30(3):133–138, February 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rajan:1989:EPA

[RGG89] Vijayan Rajan, R. K. Ghosh, and P. Gupta. An efficient parallel algorithm for random sampling. Information Processing Letters, 30(5):

265–268, March 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ravi:1987:APS

[RH87] S. S. Ravi and H. B. Hunt, III. An application of the Planar Separator Theorem to counting problems. Information Processing Letters, 25(5): 317–321, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rosenkrantz:1988:MMF

[RH88] Daniel J. Rosenkrantz and Harry B.

Hunt, III. Matrix multiplication for finite algebraic systems. Information Processing Letters, 28(4): 189–192, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Richards:1983:SDH

[Ric83] Ronald C. Richards. Shape distribution of height-balanced trees. Information Processing Letters, 17(1):17– 20, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Richards:1986:DCG

[Ric86] Dana Richards. Data compression and Gray-code sorting. Information Processing Letters, 22(4):201– 205, April 17, 1986. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramesh:1983:LPF

[RM83] S. Ramesh and S. L. Mehndiratta.

The liveness property of on-the-fly garbage collector—a proof. Information Processing Letters, 17(4):

189–195, November 8, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:1985:ALC

[RM85] S. Upendra Rao and A. K. Majumdar. An algorithm for local compaction of horizontal microprograms. Information Processing Letters, 20(1):29–33, January 2, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robson:1980:SAN

[Rob80] J. M. Robson. Storage allocation is NP-hard. Information Processing Letters, 11(3):119–125, November 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roberts:1989:ANR

[Rob89a] George H. Roberts. Another note on recursive ascent. Information Processing Letters, 32(5):263–266, September 22, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robson:1989:SSS

[Rob89b] J. M. Robson. Separating strings with small automata. Information Processing Letters, 30(4):209– 214, February 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rohrich:1982:HQW

[Röh82] Johannes Röhrich. A hybrid of Quicksort with O(n log n) worst case complexity. Information Processing Letters, 14(3):119–123, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rohl:1983:FLQ

[Roh83] J. S. Rohl. A faster lexicographical N queens algorithm. Information Processing Letters, 17(5):231– 233, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rohnert:1986:SPP

[Roh86] Hans Rohnert. Shortest paths in the plane with convex polygonal obstacles. Information Processing Letters, 23(2):71–76, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rohnert:1988:TSE

[Roh88] Hans Rohnert. Time and space efficient algorithms for shortest paths between convex polygons. Information Processing Letters, 27(4):175– 179, April 8, 1988. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rolim:1989:PIC

[Rol89] José D. P. Rolim. On the polynomial IO-complexity. Information Processing Letters, 33(4):199–204, December 21, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Romani:1980:SPP

[Rom80] Francesco Romani. Shortest-path problem is not harder than matrix multiplication. Information Processing Letters, 11(3):134–136, November 18, 1980. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic). See also note [Mor81].

Romeuf:1988:SPU

[Rom88] Jean-François Romeuf. Shortest path under rational constraint. Information Processing Letters, 28(5): 245–248, August 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rosier:1986:NPA

[Ros86] Louis E. Rosier. A note on Presburger arithmetic with array segments, permutation and equality. Information Processing Letters, 22 (1):33–35, January 2, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Roth:1989:NWC

[Rot89] Peter Roth. A note on word chains and regular languages. Information Processing Letters, 30(1):15– 18, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rowe:1984:DIS

[Row84] Neil C. Rowe. Diophantine inference on a statistical database. Information Processing Letters, 18(1): 25–31, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramalingam:1988:TDI

[RR88a] G. Ramalingam and C. Pandu Rangan. Total domination in interval graphs revisited. Information Processing Letters, 27(1):17–21, February 15, 1988. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramalingam:1988:UAD

[RR88b] G. Ramalingam and C. Pandu Rangan. A unified approach to domination problems on interval graphs. Information Processing Letters, 27(5): 271–274, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:1989:LAD

[RR89a] A. Srinivasa Rao and C. Pandu Rangan. Linear algorithm for domatic number problem on interval graphs. Information Processing Letters, 33 (1):29–33, October 27, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rao:1989:OPA

[RR89b] A. Srinivasa Rao and C. Pandu Rangan. Optimal parallel algorithms on circular-arc graphs. Information Processing Letters, 33(3):147– 156, November 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rayward-Smith:1987:CPS

[RS87] V. J. Rayward-Smith. The complexity of preemptive scheduling given interprocessor communication delays. Information Processing Letters, 25(2):123–125, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1989:PAC

[RS89] Wojciech Rytter and Tomasz Szymacha. Parallel algorithms for a class of graphs generated recursively. Information Processing Letters, 30(5):225–231, March 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ramamohanarao:1981:HAT

[RSD81] K. Ramamohanarao and R. Sacks-Davis. Hardware address translation for machines with a large virtual memory. Information Processing Letters, 13(1):23–29, October 27, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rayward-Smith:1980:FLC

[RSR80] V. J. Rayward-Smith and R. N. Rolph. Finding linear and circular sequences of minimal and maximal total adjacency. Information Processing Letters, 10(1):9–13, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rauchle:1985:EDC

[RT85a] Thomas Rauchle and Sam Toueg. Exposure to deadlock for communicating processes is hard to detect. Information Processing Letters, 21(2):63–68, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robert:1985:SAL

[RT85b] Yves Robert and Maurice Tchuente. A systolic array for the longest common subsequence problem. Information Processing Letters, 21(4):191– 198, October 7, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Robert:1989:DAS

[RTV89] Yves Robert, Bernard Tourancheau, and Gilles Villard. Data allocation strategies for the Gauss and Jordan algorithms on a ring of processors. Information Processing Letters, 31 (1):21–29, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ruby:1987:LPP

[Rub87] Jonathan Ruby. A liveness property of a parallel algorithm. Information Processing Letters, 24(4): 275–277, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rusinowitch:1987:TDS

[Rus87] Michael Rusinowitch. On termination of the direct sum of termrewriting systems. Information Processing Letters, 26(2):65–70, October 19, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Raghavan:1987:BBC

[RV87] Vijay Raghavan and Shankar M.

Venkatesan. On bounds for a board covering problem. Information Processing Letters, 25(5):281–284, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rey:1987:DLM

[RW87] Claudio Rey and Rabab Ward. On determining the on-line minimax linear fit to a discrete point set in the plane. Information Processing Letters, 24(2):97–101, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rystsov:1983:PCP

[Rys83] I. K. Rystsov. Polynomial complete problems in automata theory. Information Processing Letters, 16(3): 147–151, April 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1981:DSR

[Ryt81a] Wojciech Rytter. The dynamic simulation of recursive and stack manipulating programs. Information Processing Letters, 13(2):58– 63, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1981:HLR

[Ryt81b] Wojciech Rytter. A hardest language recognized by two-way nondeterministic pushdown automata. Information Processing Letters, 13(4– 5):145–146, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1981:TCL

[Ryt81c] Wojciech Rytter. Time complexity of languages recognized by one-way multihead pushdown automata. Information Processing Letters, 13(4– 5):142–144, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1982:NTW

[Ryt82a] Wojciech Rytter. A note on twoway nondeterministic pushdown automata. Information Processing Letters, 15(1):5–9, August 19, 1982.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1982:TCU

[Ryt82b] Wojciech Rytter. Time complexity of unambiguous path systems. Information Processing Letters, 15(3): 102–104, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1983:SRT

[Ryt83a] Wojciech Rytter. A simulation result for two-way pushdown automata. Information Processing Letters, 16(4):199–202, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1983:TCL

[Ryt83b] Wojciech Rytter. Time complexity of loop-free two-way pushdown automata. Information Processing Letters, 16(3):127–129, April 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1984:LCF

[Ryt84] Wojciech Rytter. On linear contextfree languages and one-way multihead automata. Information Processing Letters, 19(4):163–166, November 12, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1986:AMA

[Ryt86a] Wojciech Rytter. An application of Mehlhorn’s algorithm for bracket languages to log(N) space recognition of input-driven languages. Information Processing Letters, 23(2): 81–84, August 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1986:SCU

[Ryt86b] Wojciech Rytter. The space complexity of the unique decipherability problem. Information Processing Letters, 23(1):1–3, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1988:EPC

[Ryt88] Wojciech Rytter. On efficient parallel computations of costs of paths on a grid graph. Information Processing Letters, 29(2):71–74, September 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Rytter:1989:NOP

[Ryt89] Wojciech Rytter. A note on optimal parallel transformations of regular expressions to nondeterministic finite automata. Information Processing Letters, 31(2):103–109, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sabelfeld:1980:LTE

[Sab80] V. K. Sabel′fel′d. The logic-termal equivalence is polynomial-time decidable. Information Processing Letters, 10(2):57–62, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sabelfeld:1981:TEL

[Sab81] V. K. Sabel′fel′d. Tree equivalence of linear recursive schemata is polynomial-time decidable. Information Processing Letters, 13(4–5):

147–153, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sakkinen:1988:CDS

[Sak88] M. Sakkinen. Comments on “Manipulation of data structures without pointers” [Inform. Process. Lett. 26(3), 23 November 1997, pp. 135– 143]. Information Processing Letters, 27(6):327–328, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [SO87].

Salowe:1989:IIS

[Sal89] Jeffrey S. Salowe. L-infinity interdistance selection by parametric search. Information Processing Letters, 30(1):9–14, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Samet:1985:BC

[Sam85] Hanan Samet. Bidirectional coroutines. Information Processing Letters, 21(1):1–6, July 10, 1985. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Santoro:1980:EFR

[San80] Nicola Santoro. Extending the four Russians’ bound to general matrix multiplication. Information Processing Letters, 10(2):87–88, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sandhu:1988:CIT

[San88] Ravinderpal S. Sandhu. Cryptographic implementation of a tree hierarchy for access control. Information Processing Letters, 27(2):95– 98, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sandhu:1989:DOS

[San89a] Ravinderpal Singh Sandhu. The demand operation in the schematic protection model. Information Processing Letters, 32(4):213–219, September 1, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sandhu:1989:RTH

[San89b] Ravinderpal Singh Sandhu. The reflected tree hierarchy for protection and sharing. Information Processing Letters, 30(1):21–26, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Saoudi:1984:ITL

[Sao84] A. Saoudi. Infinitary tree languages recognized by ω-automata. Information Processing Letters, 18(1):15– 19, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sarwate:1980:NUC

[Sar80] Dilip V. Sarwate. A note on: “Universal classes of hash functions” [J. Comput. System Sci. 18 (1979), no. 2, 143–154; MR 80f:68110a] by J. L. Carter and M. N. Wegman. Information Processing Letters, 10(1): 41–45, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sattolo:1986:AGR

[Sat86] Sandra Sattolo. An algorithm to generate a random cyclic permutation. Information Processing Letters, 22(6):315–317, May 30, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Savage:1980:MMT

[Sav80] Carla Savage. Maximum matchings and trees. Information Processing Letters, 10(4–5):202–205, July 5, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Savage:1982:DFS

[Sav82] Carla Savage. Depth-first search and the vertex cover problem. Information Processing Letters, 14(5): 233–235, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Savage:1988:RMO

[Sav88] Carla Savage. Recognizing majority on a one-way mesh. Information Processing Letters, 27(5):221– 225, April 28, 1988. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Skillicorn:1989:PPC

[SB89] D. B. Skillicorn and D. T. Barnard.

Parallel parsing on the Connection Machine. Information Processing Letters, 31(3):111–117, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schnorr:1981:HMP

[Sch81a] C. P. Schnorr. How many polynomials can be approximated faster than they can be evaluated? Information Processing Letters, 12(2):76–78, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwartz:1981:FMD

[Sch81b] Jacob T. Schwartz. Finding the minimum distance between two convex polygons. Information Processing Letters, 13(4–5):168–170, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schmitt:1982:CPF

[Sch82a] Alfred Schmitt. On the computational power of the floor function. Information Processing Letters, 14 (1):1–3, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schoor:1982:FAS

[Sch82b] Amir Schoor. Fast algorithm for sparse matrix multiplication. Information Processing Letters, 15(2):87– 89, September 6, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schoning:1983:S

[Sch83a] Uwe Schöning. On the structure of Δ_{2}^{p}. Information Processing Letters, 16(4):209–211, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schorr:1983:PPD

[Sch83b] Amir Schorr. Physical parallel devices are not much faster than sequential ones. Information Processing Letters, 17(2):103–106, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schlorer:1984:ISC

[Sch84] Jan Schlörer. Insecurity of set controls for statistical databases. Information Processing Letters, 18(2): 67–71, February 28, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schmeck:1986:MEL

[Sch86a] Hartmut Schmeck. On the maximum edge length in VLSI layouts of complete binary trees. Information Processing Letters, 23(1):19–23, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schwer:1986:RPN

[Sch86b] Sylvianne R. Schwer. On the rationality of Petri net languages. Information Processing Letters, 22(3): 145–146, March 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schorn:1988:CST

[Sch88] Peter Schorn. A canonical simplifier for trigonometric expressions in the kinematic equation. Information Processing Letters, 29(5):241– 246, November 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schaffer:1989:ONR

[Sch89] Alejandro A. Schäffer. Optimal node ranking of trees in linear time. Information Processing Letters, 33(2): 91–96, November 10, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shen:1987:TLB

[SE87] Xiaojun Shen and Herbert Edelsbrunner. A tight lower bound on the size of visibility graphs. Information Processing Letters, 26(2):61–64, October 19, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seiferas:1988:VBO

[Sei88] Joel Seiferas. A variant of Ben-Or’s lower bound for algebraic decision trees. Information Processing Letters, 26(5):273–276, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Semba:1981:GAB

[Sem81] Ichirō Semba. Generation of all the balanced parenthesis strings in lexicographical order. Information Processing Letters, 12(4):188–192, August 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Senizergues:1981:NCC

[Sen81] G. Senizergues. A new class of C.F.L. for which the equivalence is decidable. Information Processing Letters, 13(1):30–34, October 27, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shih:1989:MWC

[SH89] Wei Kuan Shih and Wen-Lian Hsu. An O(n log n + m log log n) maximum weight clique algorithm for circular-arc graphs. Information Processing Letters, 31(3):129–134, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shanthikumar:1982:RAG

[Sha82a] J. G. Shanthikumar. A recursive algorithm to generate joint probability distribution of arrivals from exponential sources during a random time interval. Information Processing Letters, 14(5):214–217, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sharir:1982:FCS

[Sha82b] Micha Sharir. Fast composition of sparse maps. Information Processing Letters, 15(4):183–185, October 31, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shamir:1983:ECT

[Sha83] Adi Shamir. Embedding cryptographic trapdoors in arbitrary knapsack systems. Information Processing Letters, 17(2):77–79, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shannon:1988:LPA

[Sha88a] Gregory E. Shannon. A linearprocessor algorithm for depth-first search in planar graphs. Information Processing Letters, 29(3):119– 123, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sharir:1988:SWR

[Sha88b] Micha Sharir. The shortest watchtower and related problems for polyhedral terrains. Information Processing Letters, 29(5):265–270, November 24, 1988. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sharir:1989:NPS

[Sha89] Micha Sharir. A note on the Papadimitriou-Silverberg algorithm for planning optimal piecewiselinear motion of a ladder. Information Processing Letters, 32(4):187– 190, September 1, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shiloach:1981:ALD

[Shi81] Yossi Shiloach. Another look at the degree constrained subgraph problem. Information Processing Letters, 12(2):89–92, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shing:1983:OOB

[Shi83] Man-Tak Shing. Optimum ordered bi-weighted binary trees. Information Processing Letters, 17(2):67–70, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shiraishi:1989:PAM

[Shi89] Shuji Shiraishi. A parallel algorithm for the maximum 2-chain edge packing problem. Information Processing Letters, 32(6):277–279, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shmueli:1983:DCD

[Shm83] Oded Shmueli. Dynamic cycle detection. Information Processing Letters, 17(4):185–188, November 8, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shock:1986:CMC

[Sho86] Robert C. Shock. Computing the minimum cover of functional dependencies. Information Processing Letters, 22(3):157–159, March 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shonkwiler:1989:IAC

[Sho89a] R. Shonkwiler. An image algorithm for computing the Hausdorff distance efficiently in linear time. Information Processing Letters, 30(2): 87–89, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shoudai:1989:LFT

[Sho89b] Takayoshi Shoudai. The lexicographically first topological order problem is NLOG-complete. Information Processing Letters, 33(3): 121–124, November 30, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sieminski:1988:FDH

[Sie88] Andrzej Sieminski. Fast decoding of the Huffman codes. Information Processing Letters, 26(5):237– 241, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Siklossy:1981:EQE

[Sik81] L. Siklóssy. Efficient query evaluation in relational data bases with missing values. Information Processing Letters, 13(4–5):160–163, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

SilvaFilho:1981:OCD

[Sil81] Ysmar V. Silva Filho. Optimal choice of discriminators in a balanced k-d binary search tree. Information Processing Letters, 13(2): 67–70, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sharma:1989:EDD

[SIM89] Mohan B. Sharma, Sitharama S. Iyengar, and Narasimha K. Mandyam. An efficient distributed depth-first-search algorithm. Information Processing Letters, 32(4):183–186, September 1, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sassa:1987:EGP

[SIN87] Masataka Sassa, Harushi Ishizuka, and Ikuo Nakata. ECLR-attributed grammars: a practical class of LRattributed grammars. Information Processing Letters, 24(1):31– 41, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sistla:1989:VCP

[Sis89] A. P. Sistla. On verifying that a concurrent program satisfies a nondeterministic specification. Information Processing Letters, 32(1):17–23, July 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Skiena:1989:RGC

[Ski89] Steven S. Skiena. Reconstructing graphs from cut-set sizes. Information Processing Letters, 32(3):123– 127, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Skyum:1983:MWB

[Sky83] Sven Skyum. A measure in which Boolean negation is exponentially powerful. Information Processing Letters, 17(3):125–128, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slater:1982:LAN

[Sla82] Peter J. Slater. A linear algorithm for the number of degree constrained subforests of a tree. Information Processing Letters, 15(4): 186–188, October 31, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sleator:1980:TOA

[Sle80] Daniel D. K. D. B. Sleator. A 2.5 times optimal algorithm for packing in two dimensions. Information Processing Letters, 10(1):37–40, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slisenko:1982:CFG

[Sli82] A. O. Slisenko. Context-free grammars as a tool for describing polynomial-time subclasses of hard problems. Information Processing Letters, 14(2):52–56, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slusarek:1985:NDS

[Slu85] Maciej Slusarek. A note on the dynamic storage allocation problem. Information Processing Letters, 20 (5):223–227, June 12, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Slusarek:1987:OSA

[Slu87] Maciej Slusarek. An off-line storage allocation algorithm. Information Processing Letters, 24(2):71– 75, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Samet:1980:PED

[SM80] Hanan Samet and Leo Marcus. Purging in an equality data base. Information Processing Letters, 10(2): 89–95, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Seroussi:1983:ACM

[SM83] Gadiel Seroussi and Fai Ma. On the arithmetic complexity of matrix Kronecker powers. Information Processing Letters, 17(3):145–148, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Kumar:1986:PFQ

[SM86] P. Srinivas Kumar and M. Manohar. On probability of forest of quadtrees reducing to quadtrees. Information Processing Letters, 22(3): 109–111, March 3, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Suros:1987:FDR

[SM87] Rina Suros and E. Montagne. Fitted diagonals for reducing I/O bandwidth in systolic systems. Information Processing Letters, 25(5):335– 341, July 10, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Satyanarayanan:1992:NRT

[SM92] R. Satyanarayanan and D. R.

Muthukrishnan. A note on Raymond’s tree based algorithm for distributed mutual exclusion. Information Processing Letters, 43(5): 249–255, October 5, 1992. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ray89a].

Sassa:1987:SRL

[SN87] Masataka Sassa and Ikuo Nakata. A simple realization of LR-parsers for regular right part grammars. Information Processing Letters, 24(2): 113–120, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sassa:1988:TOS

[SN88] Masataka Sassa and Ikuo Nakata.

Time-optimal short-circuit evaluation of Boolean expressions. Information Processing Letters, 29(1):43– 51, September 15, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Staunstrup:1989:BSO

[SN89] Jorgen Staunstrup and Jurg Nievergelt. The behavior of shared objects: concepts, pitfalls, and a new model. Information Processing Letters, 30(3):145–151, February 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Snir:1981:CSQ

[Sni81] M. Snir. On the complexity of simplifying quadratic forms. Information Processing Letters, 12(5):217– 220, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Snir:1986:EBA

[Sni86] Marc Snir. Exact balancing is not always good. Information Processing Letters, 22(2):97–102, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sajeev:1987:MDS

[SO87] A. S. M. Sajeev and J. Olszewski. Manipulation of data structures without pointers. Information Processing Letters, 26(3):135–143, November 23, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [Sak88].

Sontag:1985:RAP

[Son85] Eduardo D. Sontag. Real addition and the polynomial hierarchy. Information Processing Letters, 20(3): 115–120, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Song:1989:DDD

[SPL^{+}89] Gyuyoun Song, Donghyeon Park, Dongmyun Lee, Kyu Ho Park, and Myunghwan Kim. A distributed deadlock detection algorithm: distributed graph reconstruction algorithm. Information Processing Letters, 30(5):245–252, March 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Spyratos:1982:HTD

[Spy82] Michael Spyratos. A homomorphism theorem for data base mappings. Information Processing Letters, 15(3): 91–96, October 11, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Steinberg:1981:LSE

[SR81] David Steinberg and Michael Rodeh. A layout for the shuffleexchange network with Θ(N^{2} log N) area. Information Processing Letters, 12(2):83–88, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schreiber:1988:ADS

[SR88a] Fabio A. Schreiber and G. Rosolini.

An algebraic description of some state-dependent failure mechanisms. Information Processing Letters, 29 (4):207–211, November 12, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shankar:1988:EPC

[SR88b] Narayan Shankar and Vijaya Ramachandran. Efficient parallel circuits and algorithms for division. Information Processing Letters, 29(6): 307–313, December 8, 1988. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sridhar:1987:NAP

[Sri87] M. K. Sridhar. A new algorithm for parallel solution of linear equations. Information Processing Letters, 24 (6):407–412, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sridhar:1988:CBG

[Sri88] M. A. Sridhar. On the connectivity of the De Bruijn graph. Information Processing Letters, 27(6): 315–318, May 13, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stoyan:1981:APR

[SS81] Yu. G. Stoyan and S. V. Smelyakov. An approach to the problems of routing optimization in the regions of intricate shape. Information Processing Letters, 13(1):39–43, October 27, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Soisalon-Soininen:1983:CAC

[SS83] Eljas Soisalon-Soininen. On computing approximate convex hulls. Information Processing Letters, 16 (3):121–126, April 15, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Scarioni:1984:PAE

[SS84a] F. Scarioni and H. G. Speranza. A probabilistic analysis of an errorcorrecting algorithm for the Towers of Hanoi puzzle. Information Processing Letters, 18(2):99– 103, February 28, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shamir:1984:CCV

[SS84b] A. Shamir and C. P. Schnorr. Cryptanalysis of certain variants of Rabin’s signature scheme. Information Processing Letters, 19(3):113– 115, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Santoro:1985:IBS

[SS85a] Nicola Santoro and Jeffrey B. Sidney. Interpolation-binary search. Information Processing Letters, 20(4): 179–181, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Shallit:1985:NTF

[SS85b] Jeffrey Shallit and Adi Shamir.

Number-theoretic functions which are equivalent to number of divisors. Information Processing Letters, 20 (3):151–153, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Salowe:1987:SUL

[SS87] Jeffrey S. Salowe and W. L. Steiger.

Stable unmerging in linear time and constant space. Information Processing Letters, 25(5):285–294, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sarkar:1989:OPC

[SS89] Dilip Sarkar and Ivan Stojmenović.

An optimal parallel circle-cover algorithm. Information Processing Letters, 32(1):3–6, July 3, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Subramanian:1987:DTP

[SSA87] K. G. Subramanian, Rani Siromoney, and P. Jeyanthi Abisha. A D0L-T0L public key cryptosystem. Information Processing Letters, 26(2): 95–97, October 19, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sagar:1989:MAT

[SSA89] G. Sagar, Anil K. Sarje, and Kamal U. Ahmed. On module assignment in two-processor distributed systems: A modified algorithm. Information Processing Letters, 32(3): 151–153, August 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stojmenovic:1986:NAC

[SSS86] Ivan Stojmenović and Eljas Soisalon-Soininen. A note on approximate convex hulls. Information Processing Letters, 22(2):55–56, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Soisalon-Soininen:1988:LLP

[SST88] Eljas Soisalon-Soininen and Jorma Tarhio. Looping LR parsers. Information Processing Letters, 26(5): 251–253, January 11, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Samet:1987:DPC

[SSW87] Hanan Samet, Clifford A. Shaffer, and Robert E. Webber. Digitizing the plane with cells of nonuniform size. Information Processing Letters, 24(6):369–375, April 6, 1987.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Staalmarck:1989:NCC

[Stå89] Gunnar Stålmarck. A note on the computational complexity of the pure classical implication calculus. Information Processing Letters, 31 (6):277–278, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stewart:1989:ACP

[Ste89] Iain A. Stewart. An algorithm for colouring perfect planar graphs. Information Processing Letters, 31(2): 97–101, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Storer:1981:CFS

[Sto81] James A. Storer. Constructing full spanning trees for cubic graphs. Information Processing Letters, 13(1): 8–11, October 27, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stout:1982:DSL

[Sto82] Quentin F. Stout. Drawing straight lines with a pyramid cellular automaton. Information Processing Letters, 15(5):233–237, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stojanovski:1986:NIP

[Sto86] Jordan Stojanovski. A note on implementing Prolog in Lisp. Information Processing Letters, 23(5):261– 264, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sundar:1989:WDS

[Sun89] Rajamani Sundar. Worst-case data structures for the priority queue with attrition. Information Processing Letters, 31(2):69–75, April 26, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Supowit:1985:DSP

[Sup85] Kenneth J. Supowit. Decomposing a set of points into chains, with applications to permutation and circle graphs. Information Processing Letters, 21(5):249–252, November 18, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Suzuki:1988:PPR

[Suz88] Ichiro Suzuki. Proving properties of a ring of finite-state machines. Information Processing Letters, 28(4): 213–214, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Stockmeyer:1982:NCS

[SV82] Larry J. Stockmeyer and Vijay V. Vazirani. NP-completeness of some generalizations of the maximum matching problem. Information Processing Letters, 15(1):14–19, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Savitch:1984:PRT

[SV84] Walter J. Savitch and Paul M. B. Vitányi. On the power of real-time two-way multihead finite automata with jumps. Information Processing Letters, 19(1):31–35, July 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/002001908490125X.

Sommerhalder:1986:PAD

[SV86] R. Sommerhalder and S. C. Van Westrhenen. A parallel O(log n) algorithm for the drawing of algebraic curves in an n × n square. Information Processing Letters, 23(5):221– 226, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Sykora:1987:TCA

[SV87] O. Sýkora and I. Vř’to. Tight chip area lower bounds for string matching. Information Processing Letters, 26(3):117–119, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Schevon:1988:PAR

[SV88] Catherine A. Schevon and Jeffrey Scott Vitter. A parallel algorithm for recognizing unordered depth-first search. Information Processing Letters, 28(2):105–110, June 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Spirakis:1984:SNH

[SY84a] Paul Spirakis and Chee K. Yap.

Strong NP-hardness of moving many discs. Information Processing Letters, 19(1):55–59, July 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/0020019084901303.

Subrahmanyam:1984:EFL

[SY84b] P. A. Subrahmanyam and J.-H. You. On embedding functions in logic. Information Processing Letters, 19(1):41–46, July 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

                                        http://www.sciencedirect.com/
                                        science/article/pii/0020019084901273.

Singhal:1988:PAC

[SY88] Mukesh Singhal and Yelena Yesha. A polynomial algorithm for computation of the probability of conflicts in a database under arbitrary data access distribution. Information Processing Letters, 27(2):69– 74, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Steiner:1993:NSU

[SY93] George Steiner and Scott Yeomans. A note on: “Scheduling unit-time tasks with integer release times and deadlines” [Inform. Process. Lett. 16 (1983), no. 4, 171–173; MR 84m:68030] by G. Frederickson. Information Processing Letters, 47(3): 165–166, September 14, 1993. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Fre83a].

Syslo:1982:LAR

[Sys82] Maciej M. Sysło. A labeling algorithm to recognize a line digraph and output its root graph. Information Processing Letters, 15(1):28–30, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szalas:1987:AAF

[Sza87] Andrzej Szałas. Arithmetical axiomatization of first-order temporal logic. Information Processing Letters, 26(3):111–116, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szalas:1988:IRP

[Sza88] Andrzej Szałas. An incompleteness result in Process Algebra. Information Processing Letters, 29(2):67– 70, September 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:1982:FPA

[Sze82] Andrzej Szepietowski. A finite 5pebble-automaton can search every maze. Information Processing Letters, 15(5):199–204, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:1987:TNF

[Sze87] Andrzej Szepietowski. There are no fully space constructible functions between log log n and log n. Information Processing Letters, 24(6): 361–362, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:1988:RLA

[Sze88] Andrzej Szepietowski. Remarks on languages acceptable in log log n space. Information Processing Letters, 27(4):201–203, April 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:1989:SNS

[Sze89a] Andrzej Szepietowski. Some notes on strong and weak log log n space complexity. Information Processing Letters, 33(2):109–112, November 10, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szepietowski:1989:SRA

[Sze89b] Andrzej Szepietowski. Some remarks on the alternating hierarchy and closure under complement for sublogarithmic space. Information Processing Letters, 33(2):73– 78, November 10, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szpankowski:1986:AAT

[Szp86] Wojciech Szpankowski. On an asymptotic analysis of a tree-type algorithm for broadcast communications. Information Processing Letters, 23(3):135–142, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szpankowski:1988:EAS

[Szp88] Wojciech Szpankowski. The evaluation of an alternative sum with applications to the analysis of some data structures. Information Processing Letters, 28(1):13–19, May 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Szwarcfiter:1987:NCC

[Szw87] Jayme Luiz Szwarcfiter. A note on the computation of the k-closure of a graph. Information Processing Letters, 24(4):279–280, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Taghva:1986:SCF

[Tag86] Kazem Taghva. Some characterizations of finitely specifiable implicational dependency families. Information Processing Letters, 23(3): 153–158, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takaoka:1986:OPM

[Tak86] Tadao Takaoka. An on-line pattern matching algorithm. Information Processing Letters, 22(6):329– 330, May 30, 1986. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takaoka:1987:DRH

[Tak87] Tadao Takaoka. A decomposition rule for the Hoare logic. Information Processing Letters, 26(4):205– 208, December 4, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Takada:1988:GIE

[Tak88] Yuji Takada. Grammatical inference for even linear languages based on control sets. Information Processing Letters, 28(4):193–199, July 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tamminen:1982:EHO

[Tam82] Markku Tamminen. Extendible hashing with overflow. Information Processing Letters, 15(5):227– 232, December 10, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tamminen:1983:AT

[Tam83] Markku Tamminen. Analysis of Ntrees. Information Processing Letters, 16(3):131–137, April 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarhio:1982:LPS

[Tar82a] Jorma Tarhio. LR parsing of some ambiguous grammars. Information Processing Letters, 14(3):101–103, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1982:HCA

[Tar82b] R. Endre Tarjan. A hierarchical clustering algorithm using strong components. Information Processing Letters, 14(1):26–29, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1982:SAM

[Tar82c] Robert Endre Tarjan. Sensitivity analysis of minimum spanning trees and shortest path trees. Information Processing Letters, 14(1): 30–33, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See corrigendum [Tar86].

Tarjan:1983:IAH

[Tar83a] Robert E. Tarjan. An improved algorithm for hierarchical clustering using strong components. Information Processing Letters, 17(1):37–41, July 19, 1983. CODEN IFPLAT.

ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1983:UBS

[Tar83b] Robert Endre Tarjan. Updating a balanced search tree in O(1) rotations. Information Processing Letters, 16(5):253–257, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tarjan:1986:SAM

[Tar86] R. E. Tarjan. Corrigendum: “Sensitivity analysis of minimum spanning trees and shortest path trees” [Inform. Process. Lett. 14(1), 27 March 1982, pp. 30–33]. Information Processing Letters, 23(4):219, November 8, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Tar82c].

Taubenfeld:1989:LEP

[Tau89] Gadi Taubenfeld. Leader election in the presence of n − 1 initial failures. Information Processing Letters, 33(1):25–28, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tazawa:1984:CRP

[Taz84] Shinsei Tazawa. On the consecutive retrieval property for generalized binary queries. Information Processing Letters, 18(5):291–293, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Treat:1984:EGS

[TB84] Joseph M. Treat and Timothy A.

Budd. Extensions to grid selector composition and compilation in APL. Information Processing Letters, 19(3):117–123, October 19, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsin:1983:GPS

[TC83] Yung H. Tsin and Francis Y. Chin. A general program scheme for finding bridges. Information Processing Letters, 17(5):269–272, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turpin:1984:EBB

[TC84] Russell Turpin and Brian A. Coan. Extending binary Byzantine Agreement to multivalued Byzantine Agreement. Information Processing Letters, 18(2):73–76, February 28, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

tenHoopen:1980:CRR

[ten80] J. ten Hoopen. Consecutive retrieval with redundancy: an optimal linear and an optimal cyclic arrangement and their storage space requirements. Information Processing Letters, 11(4–5):211–217, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tennent:1987:NUE

[Ten87a] R. D. Tennent. A note on undefined expression values in programming logics. Information Processing Letters, 24(5):331–333, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tennent:1987:QAL

[Ten87b] R. D. Tennent. Quantification in Algol-like languages. Information Processing Letters, 25(2):133–137, May 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

terBekke:1980:CD

[ter80] J. H. ter Bekke. Convertibility in databases. Information Processing Letters, 11(4–5):168–171, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thatte:1985:CBT

[Tha85] Satish R. Thatte. On the correspondence between two classes of reduction systems. Information Processing Letters, 20(2):83–85, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thompson:1985:VDM

[Tho85] Clark D. Thompson. VLSI design with multiple active layers. Information Processing Letters, 21(3):109– 111, September 5, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tiekenheinrich:1984:LBM

[Tie84] Jürgen Tiekenheinrich. A 4n-lower bound on the monotone network complexity of a one-output Boolean function. Information Processing Letters, 18(4):201–202, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tirri:1983:SRP

[Tir83] Henry R. Tirri. Simulation, reduction and preservation of correctness properties of parallel systems. Information Processing Letters, 17(1):21– 27, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Thazhuthaveetil:1987:SLR

[TJP87] Matthew Thazhuthaveetil, J., and Andrew R. Pleszkun. On the structural locality of reference in LISP list access streams. Information Processing Letters, 26(2):105–110, October 19, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tong:1983:FAF

[TL83] Po Tong and E. L. Lawler. A faster algorithm for finding edge-disjoint branchings. Information Processing Letters, 17(2):73–76, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tel:1989:CRB

[TM89] G. Tel and F. Mattern. Comments on “Ring based termination detection algorithm for distributed computations” [Inform. Process. Lett. 29(3), 26 October 1988, pp. 149– 153]. Information Processing Letters, 31(3):127–128, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [HS88a].

Tompa:1981:EST

[Tom81] Martin Tompa. An extension of Savitch’s theorem to small space bounds. Information Processing Letters, 12(2):106–108, April 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Topor:1984:TDD

[Top84] Rodney W. Topor. Termination detection for distributed computations. Information Processing Letters, 18(1):33–36, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tourlakis:1984:INT

[Tou84] George Tourlakis. An inductive number-theoretic characterization of NP. Information Processing Letters, 19(5):245–247, November 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Toyama:1987:CTD

[Toy87] Yoshihito Toyama. Counterexamples to termination for the direct sum of term rewriting systems. Information Processing Letters, 25(3): 141–143, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Toyama:1989:FKC

[Toy89] Yoshihito Toyama. Fast Knuth– Bendix completion with a term rewriting system compiler. Information Processing Letters, 32(6):325– 328, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Trakhtenbrot:1984:SET

[Tra84] M. B. Trakhtenbrot. Some equivalent transformations of recursive programs based on their schematic properties. Information Processing Letters, 18(5):275–283, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Truszczynski:1980:OMS

[Tru80] Mirosław Truszczyński. Once more on storage for consecutive retrieval. Information Processing Letters, 10 (1):21–24, February 12, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tolle:1981:CVC

[TS81] Donald MacDavid Tolle and William E. Siddall. On the complexity of vector computations in binary tree machines. Information Processing Letters, 13(3):120–124, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tinhofer:1986:BSS

[TS86] G. Tinhofer and H. Schreck. The bounded subset sum problem is almost everywhere randomly decidable in O(N). Information Processing Letters, 23(1):11–17, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsao:1981:NIB

[Tsa81] Nai Kuan Tsao. The numerical instability of Bini’s algorithm. Information Processing Letters, 12(1): 17–19, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See reply [Bin82].

Tse:1987:DUF

[Tse87] T. H. Tse. On the detection of unstructuredness in flowgraphs. Information Processing Letters, 25(3): 189–193, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsin:1985:OPP

[Tsi85] Yung Hyang Tsin. An optimal parallel processor bound in strong orientation of an undirected graph. Information Processing Letters, 20(3): 143–146, April 8, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsin:1988:HVD

[Tsi88] Yung Hyang Tsin. On handling vertex deletion in updating minimum spanning trees. Information Processing Letters, 27(4):167–168, April 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsitsiklis:1989:URN

[Tsi89] John N. Tsitsiklis. On the use of random numbers in asynchronous simulation via rollback. Information Processing Letters, 31(3):139–144, May 8, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tsuda:1983:GFM

[TST83] Takao Tsuda, Takashi Sato, and Takaaki Tatsumi. Generalization of Floyd’s model on permuting information in idealized two-level storage. Information Processing Letters, 16(4):183–188, May 13, 1983.

CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tan:1986:CDT

[TTv86] R. B. Tan, G. Tel, and J. van Leeuwen. Comments on: “Distributed termination detection algorithm for distributed computations” [Inform. Process. Lett. 22 (1986), no. 6, 311–314; MR 87j:68034a] by R. K. Arora, S. P. Rana and M. N. Gupta. Information Processing Letters, 23(3):163, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [ARG86, AG88b].

Turan:1984:CCG

[Tur84] György Turán. The critical complexity of graph properties. Information Processing Letters, 18(3): 151–153, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Turan:1989:LBS

[Tur89] György Turán. Lower bounds for synchronous circuits and planar circuits. Information Processing Letters, 30(1):37–40, January 16, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Troya:1981:AAR

[TV81] J. M. Troya and A. Vaquero. An approximation algorithm for reducing expected head movement in linear storage devices. Information Processing Letters, 13(4–5):218–220, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Tel:1987:CDA

[TV87] G. Tel and J. Van Leeuwen. Comments on “A distributed algorithm for distributed termination” [Inform. Process. Lett. 24(5), 16 March 1987, pp. 293–297]. Information Processing Letters, 25(5):349, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [HZ87].

Thome:1989:SCA

[TW89] Winfried Thome and Reinhard Wilhelm. Simulating circular attribute grammars through attribute reevaluation. Information Processing Letters, 33(2):79–81, November 10, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Udding:1986:AIS

[Udd86] Jan Tijmen Udding. Absence of individual starvation using weak semaphores. Information Processing Letters, 23(3):159–162, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ukkonen:1985:UBS

[Ukk85] Esko Ukkonen. Upper bounds on the size of LR(k) parsers. Information Processing Letters, 20(2):99– 103, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Umeo:1982:DOW

[UMS82] Hiroshi Umeo, Kenichi Morita, and Kazuhiro Sugata. Deterministic one-way simulation of two-way realtime cellular automata and its related problems. Information Processing Letters, 14(4):158–161, June 13, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Upfal:1982:FCP

[Upf82] Eliezer Upfal. Formal correctness proofs of a nondeterministic program. Information Processing Letters, 14(2):86–92, April 20, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Urbanek:1980:ACE

[Urb80] Friedrich J. Urbanek. An O(log n) algorithm for computing the nth element of the solution of a difference equation. Information Processing Letters, 11(2):66–67, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ural:1988:STS

[UY88] Hasan Ural and Bo Yang. A structural test selection criterion. Information Processing Letters, 28(3): 157–163, July 4, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Valiant:1980:CMP

[Val80] L. G. Valiant. Computing multivariate polynomials in parallel. Information Processing Letters, 11(1): 44–45, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanderNat:1980:FSA

[van80a] M. van der Nat. A fast sorting algorithm, a hybrid of distributive and merge sorting. Information Processing Letters, 10(3):163–167, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanEmdeBoas:1980:LBC

[van80b] Peter van Emde Boas. On the Ω(n log n) lower bound for convex hull and maximal vector determination. Information Processing Letters, 10(3):132–136, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [Avi80].

vandeSnepscheut:1981:SCB

[Van81] J. L. A. Van De Snepscheut. Synchronous communication between asynchronous components. Information Processing Letters, 13(3):127– 130, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VandeSnepscheut:1985:EEQ

[Van85a] Jan L. A. Van de Snepscheut. Evaluating expressions with a queue. Information Processing Letters, 20(2): 65–66, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VanGils:1985:HCF

[Van85b] W. J. Van Gils. How to cope with faulty processors in a completely connected network of communicating processors. Information Processing Letters, 20(4):207–213, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VanDeSnepscheut:1987:AFG

[Van87a] Jan L. A. Van De Snepscheut.

“Algorithms for on-the-fly garbage collection” revisited. Information Processing Letters, 24(4):211–216, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanLeijenhorst:1987:NFS

[Van87b] D. C. Van Leijenhorst. A note on the formula size of the ‘ mod k’ functions. Information Processing Letters, 24(4):223–224, March 2, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vandenBroek:1988:CIR

[Van88] P. M. Van Den Broek. Confluence of indirection reductions in graph rewrite systems. Information Processing Letters, 29(3):143–148, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vandenBos:1989:PPC

[van89a] Jan van den Bos. PROCOL: A protocol-constrained concurrent object-oriented language. Information Processing Letters, 32(5):221– 227, September 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanEmdeBoas:1989:SMS

[Van89b] Peter Van Emde Boas. Space measures for storage modification machines. Information Processing Letters, 30(2):103–110, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vardi:1981:DPD

[Var81] M. Y. Vardi. The decision problem for database dependencies. Information Processing Letters, 12(5):

251–254, October 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vardi:1984:NLD

[Var84] Moshe Y. Vardi. A note on lossless database decompositions. Information Processing Letters, 18(5): 257–260, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vardi:1989:NRT

[Var89a] Moshe Y. Vardi. A note on the reduction of two-way automata to one-way automata. Information Processing Letters, 30(5):261–264, March 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Varma:1989:FRU

[Var89b] Anujan Varma. Fault-tolerant routing in unique-path multistage interconnection networks. Information Processing Letters, 31(4):197– 201, May 22, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vikas:1986:DAA

[VB86] O. M. Vikas and Suresh Kumar Basandra. Data algebra and its application in database design. Information Processing Letters, 23(1):47– 54, July 20, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Valentine:1987:ASL

[VD87] Mark Valentine and Robert H.

Davis. The automated solution of logic puzzles. Information Processing Letters, 24(5):317–324, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Verjus:1987:PDA

[Ver87] J. Pierre Verjus. On the proof of a distributed algorithm. Information Processing Letters, 25(3):145– 147, May 29, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See comments [VT90, Pet91].

Veroy:1988:ACD

[Ver88a] Boris S. Veroy. Average complexity of divide-and-conquer algorithms. Information Processing Letters, 29(6):319–326, December 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigenda [Ver89a].

Veroy:1988:OSA

[Ver88b] Boris S. Veroy. Optimal search algorithm for a minimum of a discrete periodic bimodal function. Information Processing Letters, 29(5):233– 239, November 24, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also corrigenda [Ver89b].

Veroy:1989:CAC

[Ver89a] Boris S. Veroy. Corrigenda: “Average complexity of divide-and-conquer algorithms” [Inform. Process. Lett. 29 (1988), no. 6, 319–326; MR 89m:68068]. Information Processing Letters, 32(6):329, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ver88a].

Veroy:1989:COS

[Ver89b] Boris S. Veroy. Corrigenda: “Optimal search algorithm for a minimum of a discrete periodic bimodal function” [Inform. Process. Lett. 29 (1988), no. 5, 233–239; MR 90b:90099]. Information Processing Letters, 32(6):329, October 3, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ver88b].

Vidyasankar:1988:CLR

[Vid88] K. Vidyasankar. Converting Lamport’s regular register to atomic register. Information Processing Letters, 28(6):287–290, August 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vidyasankar:1989:EWM

[Vid89] K. Vidyasankar. An elegant 1-writer multireader multivalued atomic register. Information Processing Letters, 30(5):221–223, March 13, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vishkin:1985:EPS

[Vis85] Uzi Vishkin. On efficient parallel strong orientation. Information Processing Letters, 20(5):235–240, June 12, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vitanyi:1980:AHS

[Vit80] Paul M. B. Vitányi. Achievable high scores of ε-moves and running times in DPDA computations. Information Processing Letters, 10(2):83–86, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vitter:1981:SMS

[Vit81] J. S. Vitter. A shared-memory scheme for coalesced hashing. Information Processing Letters, 13(2): 77–79, November 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vitanyi:1985:LBT

[Vit85a] Paul M. B. Vitányi. An n^{1.618} lower bound on the time to simulate one queue or two pushdown stores by one tape. Information Processing Letters, 21(3):147–152, September 5, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vitanyi:1985:STO

[Vit85b] Paul M. B. Vitányi. Square time is optimal for simulation of one pushdown store or one queue by an oblivious one-head tape unit. Information Processing Letters, 21(2):87–91, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanLeeuwen:1982:ERR

[vN82] Jan van Leeuwen and Maurice Nivat. Efficient recognition of rational relations. Information Processing Letters, 14(1):34–38, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vogt:1989:NAO

[Vog89] Carsten Vogt. A new approach to optimal cache scheduling. Information Processing Letters, 30(6): 303–310, March 28, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Volger:1985:SRA

[Vol85] Hugo Volger. Some results on addition/subtraction chains. Information Processing Letters, 20(3):155– 160, April 8, 1985. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vossen:1988:NCF

[Vos88] Gottfried Vossen. A new characterization of FD implication with an application to update anomalies. Information Processing Letters, 29(3): 131–135, October 26, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanBaronaigien:1988:GAT

[vR88] D. Roelants van Baronaigien and Frank Ruskey. Generating t-ary trees in A-order. Information Processing Letters, 27(4):205–213, April 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Venkatesh:1987:OCL

[VRL87] K. Venkatesh, T. Radhakrishnan, and H. F. Li. Optimal checkpointing and local recording for domino-free rollback recovery. Information Processing Letters, 25(5):295–303, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vrto:1987:ATC

[Vř’t87] Imrich Vř’to. The area-time complexity of the VLSI counter. Information Processing Letters, 25(6): 397–400, July 26, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

VanGasteren:1990:CPD

[VT90] A. J. M. Van Gasteren and G. Tel. Comments on “On the proof of a distributed algorithm”: always-true is not invariant [Inform. Process. Lett. 25(3), 29 May 1987, pp. 145– 147]. Information Processing Letters, 35(6):277–279, September 15, 1990. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Ver87, Pet91].

VandeSnepscheut:1986:AIC

[VU86] Jan L. A. Van de Snepscheut and Jan Tijmen Udding. An alternative implementation of communication primitives. Information Processing Letters, 23(5):231–238, November 24, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanLeijenhorst:1987:RCT

[VV87] D. C. Van Leijenhorst and Th. P. Van der Weide. On a recursion connected with tree balancing algorithms. Information Processing Letters, 24(3):189–192, February 13, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

vanLeeuwen:1980:ELS

[vW80a] J. van Leeuwen and D. Wood. Errata: “List scheduling bounds for UET systems with resources” [Inform. Process. Lett. 10 (1980), no. 1, 28–31; MR 81a:68042]. Information Processing Letters, 11(1):57, August 29, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Llo80b].

vanLeeuwen:1980:DDS

[vW80b] Jan van Leeuwen and Derick Wood. Dynamization of decomposable searching problems. Information Processing Letters, 10(2):51–56, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vyskoc:1983:NPI

[Vys83] Jozef Vyskoč. A note on the power of integer division. Information Processing Letters, 17(2):71–72, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Vyskoc:1984:NBM

[Vys84] Jozef Vyskoč. A note on Boolean matrix multiplication. Information Processing Letters, 19(5):249–251, November 26, 1984. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wagner:1986:ICL

[Wag86] Klaus W. Wagner. On the intersection of the class of linear context-free languages and the class of singlereset languages. Information Processing Letters, 23(3):143–146, October 22, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Walsh:1982:THR

[Wal82] T. R. Walsh. The Towers of Hanoi revisited: moving the rings by counting the moves. Information Processing Letters, 15(2):64– 67, September 6, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Walsh:1983:ISB

[Wal83] T. R. Walsh. Iteration strikes back—at the cyclic Towers of Hanoi. Information Processing Letters, 16 (2):91–93, February 26, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Walsh:1984:HES

[Wal84] T. R. Walsh. How evenly should one divide to conquer quickly? Information Processing Letters, 19(4):203– 208, November 12, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1980:SNR

[Wan80] Patrick Shen-Pei Wang. Some new results on isotonic array grammars. Information Processing Letters, 10 (3):129–131, April 18, 1980. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Watanabe:1980:AAR

[Wat80] O. Watanabe. Another application of recursion introduction. Information Processing Letters, 10(3): 116–119, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Watanabe:1981:FAF

[Wat81] Osamu Watanabe. A fast algorithm for finding all shortest paths. Information Processing Letters, 13(1): 1–3, October ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Watjen:1985:FAT

[Wät85] Dietmar Wätjen. Feedback automata and their languages. Information Processing Letters, 21(2): 83–86, August 16, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Watanabe:1988:HOW

[Wat88] Osamu Watanabe. On hardness of one-way functions. Information Processing Letters, 27(3):151–157, March 25, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Webster:1986:CDE

[WB86] Max B. Webster and Paul W. Baker.

A class of differential equations for testing variable step-size integration. Information Processing Letters, 22(2):103–107, January 18, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weems:1988:SPA

[Wee88] Bob P. Weems. A study of page arrangements for extendible hashing. Information Processing Letters, 27(5):245–248, April 28, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wegener:1982:BPA

[Weg82a] Ingo Wegener. Best possible asymptotic bounds on the depth of monotone functions in multivalued logic. Information Processing Letters, 15 (2):81–83, September 6, 1982. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wegner:1982:SLL

[Weg82b] Lutz M. Wegner. Sorting a linked list with equal keys. Information Processing Letters, 15(5):205–208, December 10, 1982. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wegener:1983:RMF

[Weg83] Ingo Wegener. Relating monotone formula size and monotone depth of Boolean functions. Information Processing Letters, 16(1):41–42, January 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wegener:1985:OSP

[Weg85] Ingo Wegener. Optimal search with positive switch cost is NP-hard. Information Processing Letters, 21(1): 49–52, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weiser:1983:RSB

[Wei83] Mark Weiser. Reconstructing sequential behavior from parallel behavior projections. Information Processing Letters, 17(3):129–135, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Welzl:1985:CVG

[Wel85] Emo Welzl. Constructing the visibility graph for n-line segments in O(n^{2}) time. Information Processing Letters, 20(4):167–171, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weyuker:1984:CDF

[Wey84] E. J. Weyuker. The complexity of data flow criteria for test data selection. Information Processing Letters, 19(2):103–109, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wood:1983:MDP

[WFL83] Christopher R. Wood, Eduardo B.

Fernandez, and Tomas Lang. Minimization of demand paging for the LRU stack model of program behavior. Information Processing Letters, 16(2):99–104, February 26, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Whitesides:1981:AFC

[Whi81] S. H. Whitesides. An algorithm for finding clique cut-sets. Information Processing Letters, 12(1):31– 32, February 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Waxman:1988:WCP

[WI88] Bernard M. Waxman and Makoto Imase. Worst-case performance of Rayward-Smith’s Steiner tree heuristic. Information Processing Letters, 29(6):283–287, December 8, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1980:BSB

[Wil80a] M. H. Williams. Batch sizes for the batching method of colouring planar maps. Information Processing Letters, 11(4–5):186–189, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1980:CMC

[Wil80b] M. H. Williams. Cubic map configurations. Information Processing Letters, 11(4–5):180–185, December 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wilhelm:1981:MTT

[Wil81a] Reinhard Wilhelm. A modified treeto-tree correction problem. Information Processing Letters, 12(3):127– 132, June 13, 1981. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1981:STE

[Wil81b] M. H. Williams. A systematic test for extended operator precedence. Information Processing Letters, 13 (4–5):187–190, End ??, 1981. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Willard:1983:LLW

[Wil83a] Dan E. Willard. Log-logarithmic worst-case range queries are possible in space Θ(N). Information Processing Letters, 17(2):81–84, August 24, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Willett:1983:TKS

[Wil83b] Michael Willett. Trapdoor knapsacks without superincreasing structure. Information Processing Letters, 17(1):7–11, July 19, 1983. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1983:ESS

[Wil83c] M. H. Williams. Is an exit statement sufficient? Information Processing Letters, 17(1):47–51, July 19, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Williams:1983:PAP

[Wil83d] M. H. Williams. The problem of absolute privacy. Information Processing Letters, 17(3):169–171, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wilf:1984:BET

[Wil84] Herbert S. Wilf. Backtrack: an O(1) expected time algorithm for the graph coloring problem. Information Processing Letters, 18(3): 119–121, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wiltink:1987:DND

[Wil87] J. G. Wiltink. A deficiency of natural deduction. Information Processing Letters, 25(4):233–234, June 17, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also comments [Bal87].

Winkowski:1981:PAO

[Win81] Jozef Winkowski. Protocols of accessing overlapping sets of resources. Information Processing Letters, 12 (5):239–243, October 13, 1981. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Winkler:1982:CMD

[Win82] Peter M. Winkler. On computability of the mean deviation. Information Processing Letters, 15(1):36–38, August 19, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wise:1985:RMQ

[Wis85] David S. Wise. Representing matrices as quadtrees for parallel processors. Information Processing Letters, 20(4):195–199, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Witaszek:1988:PMF

[Wit88] Jacek Witaszek. A practical method for finding the optimum postponement transformation for LR(k) parsers. Information Processing Letters, 27(2):63–67, February 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wang:1988:STG

[WK88] D. W. Wang and Yue-Sun Kuo. A study on two geometric location problems. Information Processing Letters, 28(6):281–286, August 29, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wong:1989:NAC

[WM89] Wing Shing Wong and Robert J. T. Morris. A new approach to choosing initial points in local search. Information Processing Letters, 30(2): 67–72, January 30, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wogulis:1989:SSS

[Wog89] James Wogulis. Self-adjusting and split sequence hash tables. Information Processing Letters, 30(4):185– 188, February 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wolfson:1987:CET

[Wol87] Ouri Wolfson. Concurrent execution of transaction copies. Information Processing Letters, 24(2):87– 93, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wolfram:1989:FCI

[Wol89] D. A. Wolfram. Forward checking and intelligent backtracking. Information Processing Letters, 32(2):85– 87, July 24, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wong:1989:CCC

[Won89] Kam-Fai Wong. Comments on “A comparison of concatenated and superimposed code word surrogate files for very large data/knowledge bases”. Information Processing Letters, 33(1):45–52, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wood:1984:CPR

[Woo84] Derick Wood. The contour problem for rectilinear polygons. Information Processing Letters, 19(5):229– 236, November 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wright:1987:NES

[Wri87] William E. Wright. A note on external sorting using almost single input buffering. Information Processing Letters, 24(6):403–405, April 6, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wallis:1980:EIA

[WS80a] Peter J. L. Wallis and Bernard W. Silverman. Efficient implementation of the Ada overloading rules. Information Processing Letters, 10(3): 120–123, April 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wilson:1980:ACG

[WS80b] Thomas C. Wilson and Joseph Shortt. An O(log n) algorithm for computing general order-k Fibonacci numbers. Information Processing Letters, 10(2):68–75, March 18, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Watjen:1982:AVE

[WS82] Dietmar Wätjen and Werner Struckmann. An algorithm for verifying equations of morphisms in a category. Information Processing Letters, 14(3):104–108, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Weiss:1988:BCS

[WS88] Mark Allen Weiss and Robert Sedgewick. Bad cases for Shakersort. Information Processing Letters, 28(3):133–136, July 4, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Woo:1984:PDP

[WSA84] Nam Sung Woo, Carl H. Smith, and Ashok Agrawala. A proof of the determinacy property of the data flow schema. Information Processing Letters, 19(1):13–16, July 26, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). URL

http://www.sciencedirect.com/
science/article/pii/0020019084901212.

Wang:1987:TPA

[WT87] Cao An Wang and Yung H. Tsin.

An O(log n) time parallel algorithm for triangulating a set of points in the plane. Information Processing Letters, 25(1):55–60, April 20, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Watjen:1988:DSL

[WU88] Dietmar Wätjen and Erwin Unruh.

On the degree of synchronization of k lT0L and k lET0L systems. Information Processing Letters, 29(2):87– 89, September 30, 1988. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wu:1989:FAD

[Wu89] Xiaolin Wu. Fast approximations to discrete optimal quantization. Information Processing Letters, 31(4): 175–179, May 22, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Widmayer:1985:OAM

[WW85] P. Widmayer and C. K. Wong. An optimal algorithm for the maximum alignment of terminals. Information Processing Letters, 20(2):75– 82, February 15, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Widmayer:1987:TSO

[WW87] Peter Widmayer and Derick Wood. Time- and space-optimal contour computation for a set of rectangles. Information Processing Letters, 24(5):335–338, March 16, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wan:1989:AAF

[WW89a] S. J. Wan and S. K. M. Wong. An adaptive algorithm for finding a covering hypersphere. Information Processing Letters, 33(1):7–10, October 27, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wanke:1989:UBP

[WW89b] Egon Wanke and Manfred Wiegers. Undecidability of the bandwidth problem on linear graph languages. Information Processing Letters, 33 (4):193–197, December 21, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Wyrostek:1983:CPP

[Wyr83] Piotr Wyrostek. On the “correct prefix property” in precedence parsers. Information Processing Letters, 17(3):161–165, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [Wyr84].

Wyrostek:1984:ECP

[Wyr84] Piotr Wyrostek. Erratum: “On the ‘correct prefix property’ in precedence parsers” [Inform. Process. Lett. 17 (1983), no. 3, 161–165, MR 85a:68112]. Information Processing Letters, 19(2):111, August 31, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [Wyr83].

Yamasaki:1981:WPP

[Yam81] Hideki Yamasaki. On weak persistency of Petri nets. Information Processing Letters, 13(3):94–97, December 13, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yao:1980:NAE

[Yao80] Andrew C. Yao. A note on the analysis of extendible hashing. Information Processing Letters, 11(2):84–86, October ??, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yap:1983:HAS

[Yap83] Chee K. Yap. A hybrid algorithm for the shortest path between two nodes in the presence of few negative arcs. Information Processing Letters, 16 (4):181–182, May 13, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yasuura:1981:WDC

[Yas81] Hiroto Yasuura. Width and depth of combinational logic circuits. Information Processing Letters, 13(4– 5):191–194, End ??, 1981. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yamasaki:1986:RDD

[YD86] Susumu Yamasaki and Shuji Doshita. Resolution deduction to detect satisfiability for another class including non-Horn sentences in propositional logic. Information Processing Letters, 23(4):201–207, November 8, 1986. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yehudai:1985:NCD

[Yeh85] Amiram Yehudai. A note on chains of deterministic pushdown transducers. Information Processing Letters, 20(4):221–222, May 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yu:1983:UDC

[YG83] Yao Tin Yu and Mohamed G. Gouda. Unboundedness detection for a class of communicating finitestate machines. Information Processing Letters, 17(5):235–240, December 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yannakakis:1987:MCS

[YG87] Mihalis Yannakakis and Fănică Gavril. The maximum k-colorable subgraph problem for chordal graphs. Information Processing Letters, 24(2):133–137, January 30, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuba:1987:BSN

[YH87] Toshitsugu Yuba and Mamoru Hoshi. Binary search networks: a new method for key searching. Information Processing Letters, 24(1): 59–65, January 15, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yokomori:1983:SLP

[YJ83] Takashi Yokomori and Aravind K.

Joshi. Semi-linearity, Parikhboundedness and tree adjunct languages. Information Processing Letters, 17(3):137–143, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yoccoz:1989:RRP

[Yoc89] Serge Yoccoz. Recursive ω-rule for proof systems. Information Processing Letters, 31(6):291–294, June 19, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yodogawa:1984:NAG

[Yod84] Eiji Yodogawa. A note on array grammars. Information Processing Letters, 18(1):51–54, January 20, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yoo:1985:ATC

[Yoo85] Chae Woo Yoo. An approach to the transportation of computer software. Information Processing Letters, 21(3):153–157, September 5, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yu:1989:PLD

[Yu89] Sheng Yu. A pumping lemma for deterministic context-free languages. Information Processing Letters, 31 (1):47–51, April 12, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yung:1985:SUK

[Yun85] Mordechai M. Yung. A secure and useful ‘keyless cryptosystem’. Information Processing Letters, 21(1):35– 38, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yokomori:1982:TRN

[YWL82] Takashi Yokomori, Derick Wood, and Klaus-Jörn Lange. A threerestricted normal form theorem for ETOL languages. Information Processing Letters, 14(3):97–100, May 16, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See also erratum [YWL85].

Yokomori:1985:ETR

[YWL85] Takashi Yokomori, Derick Wood, and Klaus-Jörn Lange. Erratum: “A three-restricted normal form theorem for ETOL languages” [Inform. Process. Lett. 14 (1982), no. 3, 97– 100; MR 83g:68115]. Information Processing Letters, 21(1):53, July 10, 1985. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic). See [YWL82].

Yamasaki:1984:NCI

[YYDH84] Susumu Yamasaki, Mikio Yoshida, Shuji Doshita, and Mikito Hirata. A new combination of input and unit deductions for Horn sentences. Information Processing Letters, 18(4): 209–213, May 14, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Yuba:1983:CML

[YYS83] Toshitsugu Yuba, Yoshinori Yamaguchi, and Toshio Shimada. A control mechanism of a Lisp-based data-driven machine. Information Processing Letters, 16(3):139–143, April 15, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zaks:1982:GRA

[Zak82] Shmuel Zaks. Generation and ranking of K-ary trees. Information Processing Letters, 14(1):44–48, March 27, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zak:1984:LEF

[Žák84] Stanislav Žák. Letter to the editor:

“Finitary and infinitary interpretations of languages” [Math. Systems Theory 15 (1981/82), no. 3, 251– 265; MR 83h:68119] by H. A. Maurer, A. Salomaa and D. Wood. Information Processing Letters, 18(5): 297–298, June 18, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zaniolo:1980:MTF

[Zan80] Carlo Zaniolo. Mixed transitivity for functional and multivalued dependencies in database relations. Information Processing Letters, 10(1): 32–34, February 12, 1980. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zahorjan:1983:EBT

[ZBS83] John Zahorjan, Barbara J. Bell, and Kenneth C. Sevcik. Estimating block transfers when record access probabilities are non-uniform. Information Processing Letters, 16(5): 249–252, June 10, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhu:1987:NDB

[ZC87] Yun Zhou Zhu and To Yat Cheung. A new distributed breadthfirst-search algorithm. Information Processing Letters, 25(5):329–333, July 10, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zemanek:1982:OVM

[Zem82] Heinz Zemanek. Obituary: Victor Mikhaĭlovich Glushkov (1923–1982). Information Processing Letters, 14 (5):236–237, July 23, 1982. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zemel:1984:ALM

[Zem84] Eitan Zemel. An O(n) algorithm for the linear multiple choice Knapsack problem and related problems. Information Processing Letters, 18(3): 123–128, March 30, 1984. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zerovnik:1989:RHA

[Zer89] J. Zerovnik. A randomised heuristical algorithm for estimating the chromatic number of a graph. Information Processing Letters, 33(4): 213–219, December 21, 1989. CO-DEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zalcstein:1989:ATS

[ZG89] Yechezkel Zalcstein and Max Garzon. An NC^{2} algorithm for testing similarity of matrices. Information Processing Letters, 30(5):253– 254, March 13, 1989. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zheng:1983:DMA

[Zhe83] Zhi Jie Zheng. The duodirun merging algorithm: a new fast algorithm for parallel merging. Information Processing Letters, 17(3):167– 168, October 5, 1983. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zippel:1989:ESR

[Zip89] Richard Zippel. An explicit separation of relativised random polynomial time and relativised deterministic polynomial time. Information Processing Letters, 33(4):207– 212, December 21, 1989. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zhu:1987:TMS

[ZLS87] Mingfa Zhu, Nan K. Loh, and Pepe Siy. Towards the minimum set of primitive relations in temporal logic. Information Processing Letters, 26(3):121–126, November 23, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zubair:1987:TES

[ZM87] M. Zubair and B. B. Madan. Time efficient systolic architecture for matrix*vector multiplication. Information Processing Letters, 24(4):225– 231, March 2, 1987. CODEN IF-PLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Zobel:1987:TCF

[Zöb87] Dieter Zöbel. Transformations for communication fairness in CSP. Information Processing Letters, 25(3): 195–198, May 29, 1987. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).